### v1

# Characterizations of certain matroids by maximizing valuative invariants

Abstract.
Luis Ferroni and Alex Fink recently introduced a polytope of all unlabeled matroids of rank $r$ on $n$ elements, and they showed that the vertices of this polytope come from matroids that can be characterized by maximizing a sequence of valuative invariants. We prove that a number of the matroids that they conjectured to yield vertices indeed do (these include cycle matroids of complete graphs, projective geometries, and Dowling geometries), and we give additional examples (including truncations of cycle matroids of complete graphs, Bose-Burton geometries, and binary and free spikes with tips). We prove a special case of a conjecture of Ferroni and Fink by showing that direct sums of uniform matroids yield vertices of their polytope, and we prove a similar result for direct sums whose components are in certain restricted classes of extremal matroids.
Key words and phrases:
Matroid, extremal matroid theory, valuative invariant, polytope of matroids
1991 Mathematics Subject Classification:
Primary: 05B35
1. Introduction
By a result of Derksen and Fink [15], when the base polytope of a matroid is written as an indicator function, the base polytope of any matroid of rank $r$ on $n$ elements is a unique linear combination of base polytopes of nested matroids of rank $r$ on $n$ elements and the coefficients are integers. Matroid invariants assign the same value to isomorphic matroids. Valuative invariants are matroid invariants $f$ that take values in an abelian group and for which, for any matroid $M$, the image $f(M)$ can be found by taking the integer linear combination of the base polytopes of nested matroids that give the base polytope of $M$ and replacing each nested matroid by its image under $f$. (See Section 2.3 for more precise statements.) Valuative invariants include many well-known invariants, such as the Tutte polynomial, but they encompass far more than deletion-contraction invariants. Ferroni and Fink [19] introduced two polytopes of matroids of a given rank $r$ and size $n$; one is the convex hull of the coordinate vectors of the base polytopes of matroids of rank $r$ and size $n$ relative to the base polytopes of nested matroids of rank $r$ and size $n$; the other, $\Omega_{r,n}$, is a counterpart for unlabeled matroids (that is, isomorphism types of matroids). They showed that the vertices of $\Omega_{r,n}$ come from matroids, or sets of matroids that give the same point in $\Omega_{r,n}$, that can be characterized by sequences of maximization (or, equivalently, minimization) problems of valuative invariants. We illustrate such characterizations, which are the focus of this paper, with Theorem 5.2 below, which shows that, as Ferroni and Fink conjectured, the cycle matroid $M(K_{r+1})$ of the complete graph $K_{r+1}$ yields a vertex.
[…]
To see how this conforms to the description of the vertices of $\Omega_{r,\binom{r+1}{2}}$, we use the fact that the number of flats of a given size in a matroid is a valuative invariant, so maximizing the number of flats of rank $1$ and size $1$ limits the scope to simple matroids. Then we can minimize the number of lines of sizes other than two or three to say that all lines in the matroid of interest have just two or three points, and likewise we can then impose the remaining conditions on the sizes of the flats of each rank. Relative to those conditions, we show that the greatest possible number of $3$-point lines is $\binom{r+1}{3}$, and, for the final link, we show that the maximum is achieved precisely by $M(K_{r+1})$. Thus, a sequence of optimization problems for valuative invariants hones in on $M(K_{r+1})$, so $M(K_{r+1})$ gives rise to a vertex of $\Omega_{r,\binom{r+1}{2}}$.
As the illustration above suggests, these problems have the flavor of many problems that have long been considered in extremal matroid theory, but with a new twist, namely, we are restricted to optimizing valuative invariants. Thus, above we could not assume at the outset that the number of $3$-point lines was $\binom{r+1}{3}$; we could do that only after showing that that value was the maximum among matroids that satisfy the prior constraints. The study of the vertices of Ferroni and Fink’s polytopes $\Omega_{r,n}$ thus opens up a new and interesting chapter in extremal matroid theory. Appropriately, they call the matroids that yield vertices in their polytope extremal matroids.
After reviewing the necessary background in Section 2, we show that several matroids that Ferroni and Fink conjectured to be extremal indeed are: $M(K_{r+1})$ (Theorems 5.2 and 5.3) and Dowling geometries $Q_{r}(G)$ based on groups (or quasigroups, if $r=3$; Theorems 6.2, 6.3, and 6.4) are extremal. They conjectured that projective geometries are extremal; in Section 3 we explain why, more broadly, all perfect matroid designs (i.e., matroids in which flats of the same rank have the same size) are extremal; this applies to projective geometries, affine geometries, their truncations, and much more. Among matroids that Ferroni and Fink did not mention, we show that truncations of $M(K_{r+1})$ to rank at least five (Theorem 5.4), Bose-Burton geometries (Theorems 4.4 and 4.5), and binary and free spikes with tips (Theorem 7.1) are extremal. Ferroni and Fink conjecture that direct sums of extremal matroids are extremal; we do not have a proof of that conjecture, but provide evidence for it by showing that direct sums of uniform matroids are extremal (Theorem 8.1) and giving a general result (Theorem 8.2) that shows that direct sums of cycle matroids of complete graphs are extremal, and likewise for projective geometries of the same order, affine geometries of the same order, Dowling geometries based on a group that is the only group of its order, and certain spikes (Corollary 8.3).
2. Background
For basic matroid theory, we recommend [23]; we follow the notation used there. All matroids considered in this paper are finite. We use $\mathbb{N}$ for the set of positive integers, $\mathbb{N}_{0}$ for the set of nonnegative integers, and $[n]$ for the set $\{1,2,\ldots,n\}$.
2.1. A few items about matroids
A set $X$ in a matroid $M$ is cyclic if it is a union of circuits, or, equivalently, $M|X$ has no coloops. A set $X$ is cyclic in $M$ if and only if $E(M)-X$ is a flat of the dual matroid, $M^{*}$, since $X$ is a union of circuits of $M$ if and only if $E(M)-X$ is an intersection of hyperplanes (cocircuit complements) of $M^{*}$. We are most interested in cyclic flats, that is, flats that are cyclic. By the second reformulation of cyclic sets above, $F$ is a cyclic flat of $M$ if and only if $E(M)-F$ is a cyclic flat of $M^{*}$. The set of cyclic flats of $M$ is denoted $\mathcal{Z}(M)$, which is a lattice under inclusion. The least element of $\mathcal{Z}(M)$ is $\operatorname{cl}_{M}(\emptyset)$, the set of loops of $M$, and the greatest element is $E(M)-\operatorname{cl}_{M^{*}}(\emptyset)$ since $\operatorname{cl}_{M^{*}}(\emptyset)$ is the set of coloops of $M$. By the remark on duality, $\mathcal{Z}(M)$ is isomorphic to the order dual of $\mathcal{Z}(M^{*})$.
We use the notion of cover in the lattice of flats of a matroid and in the lattice of cyclic flats $\mathcal{Z}(M)$. In general, $y$ is a cover of $x$, or $y$ covers $x$, in an ordered set if $x<y$ and there is no $z$ with $x<z<y$. If $Y$ covers $X$ in the lattice of flats, then $r(Y)=r(X)+1$, but if $Y$ covers $X$ in $\mathcal{Z}(M)$, then $r(Y)-r(X)$ can be any integer in $[r(M)-r(X)]$.
The single-element extensions of $M$ by $e$ are the matroids $N$ with $E(N)=E(M)\cup\{e\}$ and $N\backslash e=M$. Modular cuts of $M$ are sets $\mathcal{M}$ of flats of $M$ that have two properties: (i) if $F\in\mathcal{M}$ and $F^{\prime}$ is a flat of $M$ with $F\subseteq F^{\prime}$, then $F^{\prime}\in\mathcal{M}$, and (ii) if $F,F^{\prime}\in\mathcal{M}$ and $(F,F^{\prime})$ is a modular pair, that is, $r(F)+r(F^{\prime})=r(F\cup F^{\prime})+r(F\cap F^{\prime})$, then $F\cap F^{\prime}\in\mathcal{M}$. If $N$ is a single-element extension of $M$ by $e$, then the set of flats $F$ of $M$ for which $r_{N}(F)=r_{N}(F\cup\{e\})$ is a modular cut. Conversely, if $\mathcal{M}$ is a modular cut of $M$ and $e\not\in E(M)$, then the following sets give the flats of a single-element extension of $M$: the set $F\cup\{e\}$ with $F\in\mathcal{M}$; the flats $F$ of $M$ for which $F\not\in\mathcal{M}$; the sets $F\cup\{e\}$ for which $F$ is a flat of $M$ and is neither in $\mathcal{M}$ nor covered by a flat in $\mathcal{M}$. The free extension of $M$ by $e$, denoted $M+e$, is the single-element extension that is determined by the modular cut =$\mathcal{M}=\{E(M)\}$. Thus, $e$ is not a coloop of $M+e$, and no proper cyclic flat of $M+e$ contains $e$.
A matroid $M$ is nested if $\mathcal{Z}(M)$ is a chain, that is, for any cyclic flats $F$ and $G$ of $M$, either $F\subseteq G$ or $G\subseteq F$. Nested matroids are known by many names, including Schubert matroids. The number of nested matroids of rank $r$ on $n$ elements, up to isomorphism, is $\binom{n}{r}$, as we now justify. Consider sequences of matroids $M_{0},M_{1},\ldots,M_{n}$, where $M_{0}$ is empty, and for each $i\in[n]$, the element $i$ is added to $M_{i-1}$ either (a) as a coloop or (b) via free extension, and option (a) is applied $r$ times; then $M_{n}$ is a nested matroid of rank $r$ on $[n]$; there are $\binom{n}{r}$ options for which elements will be added as coloops; also, by that construction, each nonempty cyclic flat is $[s]$ for some $s\in[n]$, so the construction gives just one representative of each isomorphism class of nested matroids. It is useful to encode isomorphism types of nested matroids of rank $r$ on $n$ elements as lattice paths where adding $i\in[n]$ as a coloop is recorded by a north step (denoted $N$) in position $i$, otherwise an east step (denoted $E$) is in position $i$; for the rank to be $r$ and the number of elements to be $n$, the lattice path goes from $(0,0)$ to $(n-r,r)$. For instance, the lattice path $NENEEN$ has cyclic flats $\emptyset$ of rank 00, $[2]$ of rank $1$, and $[5]$ of rank $2$; the element $6$ is a coloop of the matroid.
Recall the Scum Theorem [23, Theorem 3.3.1]: if $N$ is a minor of $M$, then there are subsets $X$ and $Y$ of $E(M)$ for which $N=(M/Y)\backslash X$ and $r(N)=r(M/Y)$; also, if $N$ has no loops, then $Y$ can be taken to be a flat of $M$. In particular, a matroid $M$ has no minor isomorphic to $U_{2,q+2}$ if and only if, for each rank-$(r(M)-2)$ flat $F$, at most $q+1$ hyperplanes of $M$ contain $F$.
We will review the matroids that may be less familiar (Dowling geometries, spikes, and Bose-Burton geometries) at the beginnings of the sections in which they are treated.
2.2. The Tutte polynomial
The Tutte polynomial is the most well-known enumerative invariant in matroid theory (see, e.g., [12, 18]). For a matroid $M$ on the set $E(M)$, its Tutte polynomial is defined to be
$T(M;x,y)=\sum_{A\subseteq E(M)}(x-1)^{r(M)-r(A)}(y-1)^{|A|-r(A)},$
which is a generating function for the $2^{|E(M)|}$ pairs $(|A|,r(A))$ with $A\subseteq E(M)$. Among the most basic properties of the Tutte polynomial are (i) the deletion/contraction formula $T(M;x,y)=T(M\backslash e;x,y)+T(M/e;x,y)$ for any $e\in E(M)$, and (ii) the direct sum factorization $T(M;x,y)=T(M_{1};x,y)T(M_{2};x,y)$; both results are routine to prove. It is easy to see (e.g., by using induction and the deletion/contraction formula) that the coefficients of $T(M;x,y)$ are nonnegative integers and $T(M;x,y)$ has no constant term.
The following well-known result of Brylawski [10] is easy to prove by induction on $|E(M)|$ using properties (i) and (ii) cited above. Recall also a key result by Tutte [27]: for any connected matroid $M$ with $|E(M)|\geq 2$ and any $e\in E(M)$, at least one of $M\backslash e$ and $M/e$ is connected.
Lemma 2.1.
A matroid $M$ is connected if and only if the coefficient of at least one of $x$ or $y$ in its Tutte polynomial $T(M;x,y)$ is nonzero. If $M$ is connected and $|E(M)|>1$, then the coefficients of $x$ and $y$ are nonzero and equal.
Corollary 2.2.
A matroid $M$ has at least $k$ connected components if and only if all terms $x^{i}y^{j}$ in $T(M;x,y)$ with $i+j<k$ have zero coefficient.
The Tutte polynomial is the universal deletion/contraction invariant, as the following result by Oxley and Welsh [24] (extending a result of Brylawski [10]) makes precise.
Theorem 2.3.
Let $R$ be a commutative ring with unity. For any $u,v,\sigma,\tau\in R$ with $\sigma$ and $\tau$ nonzero, there is a unique function $t$ from the set of all matroids into $R$ that has the following properties.
  1. (1)
    If $E(M)=\emptyset$, then $t(M)=1$.
  2. (2)
    If $e$ is a loop of $M$, then $t(M)=v\cdot t(M\backslash e)$.
  3. (3)
    If $e$ is a coloop of $M$, then $t(M)=u\cdot t(M/e)$.
  4. (4)
    If $e$ is neither a loop nor a coloop, then $t(M)=\sigma\cdot t(M\backslash e)+\tau\cdot t(M/e)$.
Furthermore, $t$ is the following evaluation of the Tutte polynomial $T(M;x,y)$:
$t(M)=\sigma^{|E(M)|-r(M)}\tau^{r(M)}T(M;u/\tau,v/\sigma).$
A matroid $M$ is Tutte unique if $T(M;x,y)=T(N;x,y)$ implies that $N$ is isomorphic to $M$. Many earlier characterizations of matroids by numerical invariants that we will use, such as the characterization of $M(K_{r+1})$ in Theorem 5.1, were proven to show that these matroids are Tutte unique, but in most cases the characterizations do not fit the pattern of maximizing a sequence of valuative invariants that is required in order to show that the matroids yield vertices of the polytope of unlabeled matroids. Some Tutte-unique matroids do not yield vertices of the polytope of unlabeled matroids (e.g., the nested matroid that is encoded by the lattice path $NENE$; see [19, Examples 3.3 and 4.5]), and some matroids that yield vertices are not Tutte unique (e.g., Dowling geometries of the same rank based on nonisomorphic groups of the same order). However, there are strong connections between the two notions.
2.3. Matroid base polytopes, valuative invariants, and the $\mathcal{G}$-invariant
We discuss matroid base polytopes very lightly; readers wanting more on this topic are referred to [2, 15, 21]. To simplify notation when discussing matroid base polytopes, it is common to consider the ground set of a matroid to be $[n]$. Given a matroid $M$ on $[n]$, a basis $B$ of $M$ can be encoded by its characteristic vector $\mathbf{v}_{B}$ in $\mathbb{R}^{n}$, that is, the vector in which entry $i$ is $1$ if and only if $i\in B$, otherwise entry $i$ is 00. The convex hull in $\mathbb{R}^{n}$ of the vectors $\mathbf{v}_{B}$, as $B$ ranges overall bases of $M$, is the matroid base polytope $\mathcal{P}(M)$.
For a subset $X$ of $\mathbb{R}^{n}$, its indicator function $\bar{\mathbb{I}}(X):\mathbb{R}^{n}\to\mathbb{R}$ is given by
$\bar{\mathbb{I}}(X)(x)=\begin{cases}1,&\text{ if }x\in X,\\ 0,&\text{ if }x\not\in X.\end{cases}$
We focus on indicator functions $\bar{\mathbb{I}}(\mathcal{P}(M))$ of base polytopes of matroids. The next result is due to Derksen and Fink [15, Theorem 5.4].
Theorem 2.4.
Let $M$ be a rank-$r$ matroid on $[n]$ and let $N_{1},N_{2},\ldots,N_{t}$ be the nested matroids of rank $r$ on $[n]$. There are unique integers $a_{1},a_{2},\ldots,a_{t}$ for which
(2.1) $\bar{\mathbb{I}}(\mathcal{P}(M))=a_{1}\cdot\bar{\mathbb{I}}(\mathcal{P}(N_{1}))+a_{2}\cdot\bar{\mathbb{I}}(\mathcal{P}(N_{2}))+\cdots+a_{t}\cdot\bar{\mathbb{I}}(\mathcal{P}(N_{t})).$
While computing the coefficients $a_{i}$ in practice is difficult, there is an elegant and useful expression for them, due to Hampe [22] and Ferroni [20], that uses the Möbius function. (Readers who are not familiar with the Möbius function might consult Rota [25] or Stanley [26]. We use the Möbius function $\mu$ on lattices, and we need just two facts: $\mu(x,x)=1$ for all $x$, and if $a<b$, then the sum of $\mu(x,b)$ over all $x$ with $a\leq x\leq b$ is 00; with that, all values $\mu(a,b)$ can be computed recursively.) For any nested matroid that has a nonzero coefficient in Equation (2.1), its chain of cyclic flats is a chain of cyclic flats of $M$ that includes $\operatorname{cl}_{M}(\emptyset)$ and $E(M)-\operatorname{cl}_{M^{*}}(\emptyset)$, the least and greatest cyclic flats of $M$. Let $\mathcal{C}(\mathcal{Z}(M))$ consist of all chains in $\mathcal{Z}(M)$ that include $\operatorname{cl}_{M}(\emptyset)$ and $E(M)-\operatorname{cl}_{M^{*}}(\emptyset)$, ordered by inclusion, along with a greatest element $\hat{1}$. The intersection of two chains in $\mathcal{C}(\mathcal{Z}(M))$ is a chain in $\mathcal{C}(\mathcal{Z}(M))$, and $\hat{1}$ is the greatest element, so $\mathcal{C}(\mathcal{Z}(M))$ is a lattice. The coefficient of a nested matroid $N_{i}$ whose cyclic flats are those in the chain $C\in\mathcal{C}(\mathcal{Z}(M))$ is the Möbius value $-\mu(C,\hat{1})$ in $\mathcal{C}(\mathcal{Z}(M))$. Figures 1 and 2 give examples of computing these coefficients.
[…]
Figure 1. A matroid $M$, its lattice $\mathcal{Z}(M)$ of cyclic flats, and the lattice $\mathcal{C}(\mathcal{Z}(M))$ of chains of cyclic flats that include $\emptyset$ and $E(M)$, with $\mu(C,\hat{1})$ shown beneath $C$. For readability, we suppress $\emptyset$ and $E(M)$ in the chains and let $A=\{a,a^{\prime}\}$, $B=\{b,b^{\prime}\}$, $C=\{c,c^{\prime}\}$, and $D=\{d,d^{\prime}\}$.
$a^{\prime}$$a$$c$$c^{\prime}$$d$$d^{\prime}$$b^{\prime}$$b$$M^{\prime}$
[…]
Figure 2. A second example of computing the coefficients $-\mu(C,\hat{1})$ in Equation (2.1) following the same conventions as in Figure 1.
The form of Equation (2.1) motivates one of several equivalent formulations of the notion of a valuation. First, by a matroid invariant, we mean a function that is defined on the set of matroids and that assigns the same value to isomorphic matroids. A matroid invariant $f$ taking values in an abelian group $A$ is valuative, or is a valuation, if, for any matroids $M_{1},M_{2},\ldots,M_{t}$ of rank $r$ on $[n]$ and any integers $a_{1},a_{2},\ldots,a_{t}$, if
$a_{1}\cdot\bar{\mathbb{I}}(\mathcal{P}(M_{1}))(x)+a_{2}\cdot\bar{\mathbb{I}}(\mathcal{P}(M_{2}))(x)+\cdots+a_{t}\cdot\bar{\mathbb{I}}(\mathcal{P}(M_{t}))(x)=0$
for all $x\in\mathbb{R}^{n}$, then $a_{1}\cdot f(M_{1})+a_{2}\cdot f(M_{2})+\cdots+a_{t}\cdot f(M_{t})$ is the identity of $A$. Thus, by Equation (2.1), knowing the values of a valuative invariant $f$ on nested matroids completely determines $f$.
For example, the indicator functions of the matroid base polytopes of the examples in Figures 1 and 2, which must be different, are written in terms of different nested matroids, and for the matroid $M^{\prime}$ in Figure 2, we see a Möbius value, $-2$, that does not occur for the matroid $M$ in Figure 1. However, any valuative invariant $f$ gives the same values on these two matroids, namely,
$4\cdot f(NENENEEE)-2\cdot f(NENNEEEE)-f(NNEENEEE),$
where $f(P)$ is the value of $f$ at the nested matroid that is encoded by the lattice path $P$; the first term is from the four coatoms in $\mathcal{C}(\mathcal{Z}(M))$ (or $\mathcal{C}(\mathcal{Z}(M^{\prime}))$) and the other terms are from the atoms, some of which, besides $\emptyset$ and $E(M)$, include a cyclic flat of rank $1$ and size $2$ (giving $NENNEEEE$) and the others include a cyclic flat of rank $2$ and size $4$ (giving $NNEENEEE$).
Derksen and Fink [15] proved a counterpart of Theorem 2.3 for valuative invariants; the universal valuative invariant is the $\mathcal{G}$-invariant, to which we turn. For a rank-$r$ matroid $M$ and permutation $\pi=e_{1}e_{2}\ldots e_{n}$ of $E(M)$, the associated rank sequence is the sequence $\underline{r}(\pi)=r_{1}r_{2}\ldots r_{n}$ where $r_{1}=r(\{e_{1}\})$ and for $j\geq 2$,
$r_{j}=r(\{e_{1},e_{2},\ldots,e_{j}\})-r(\{e_{1},e_{2},\ldots,e_{j-1}\}).$
Thus, $r$ of the $n$ entries $r_{i}$ of $\underline{r}(\pi)$ are $1$ and the rest are 00; we call sequences of this type $(n,r)$-sequences. The $(n,r)$-sequence $\underline{r}(\pi)$ shows the rank increases as one adjoins the elements of $E(M)$, one at a time, to $\emptyset$ in the order that the permutation $\pi$ gives. Note that $\{e_{i}\,:\,r_{i}=1\}$ is a basis of $M$.
Fix integers $n$ and $r$ with $0\leq r\leq n$. For each $(n,r)$-sequence $\underline{r}$, let $[\underline{r}]$ be a formal symbol. Let $\mathcal{G}(n,r)$ be the vector space, over a field $F$ of characteristic zero, that consists of all formal linear combinations, with coefficients in $F$, of the symbols $[\underline{r}]$. For a matroid $M$ of rank $r$ on $n$ elements, its $\mathcal{G}$-invariant and its coefficients $g_{\underline{r}}(M)$ are defined by
(2.2) $\mathcal{G}(M)=\sum_{\underline{r}}g_{\underline{r}}(M)[\underline{r}]=\sum_{\pi}[\underline{r}(\pi)]$
where the second sum is over all $n!$ permutations $\pi$ of $E(M)$. Thus, $g_{\underline{r}}(M)$ is the number of permutations $\pi$ of $E(M)$ for which $\underline{r}(\pi)=\underline{r}$. One can think of $\mathcal{G}(M)$ as a generating function for rank sequences.
As mentioned above, Derksen and Fink [15] showed that the $\mathcal{G}$-invariant is the universal valuative invariant: one obtains any valuative invariant $f$ with values in an abelian group $A$ by assigning values in $A$ to each symbol $[\underline{r}]$ and extending linearly, that is, getting $f(M)$ by replacing each symbol $[\underline{r}(\pi)]$ in Equation (2.2) by its image $f([\underline{r}(\pi)])$ in $A$. For example, for $A=\mathbb{Z}$ (the case of interest for this paper) and a fixed $(n,r)$-sequence $\underline{r}$, by mapping $[\underline{r}]$ to $1$ and all other rank sequences $[\underline{r}^{\prime}]$ to 00, we see that the function mapping $M$ to $g_{\underline{r}}(M)$ is valuative. Also, integer linear combinations of valuations with values in $\mathbb{Z}$ are valuations. Map each $[\underline{r}]$ to $[\underline{r}^{*}]$ where we obtain $\underline{r}^{*}$ from $\underline{r}$ by replacing each $1$ by 00, and each 00 by $1$, and reversing the order of the entries; it is routine to check that $\mathcal{G}(M)$ then maps to $\mathcal{G}(M^{*})$. Thus, valuative invariants of $M^{*}$ are also valuative invariants of $M$.
Derksen [14] showed that the Tutte polynomial is a valuative invariant. From this and the definition of a valuation, it follows easily that, for fixed nonnegative integers $i$ and $j$, the map that sends a matroid $M$ to the coefficient of $x^{i}y^{j}$ in its Tutte polynomial $T(M;x,y)$ is a valuative invariant. The following theorem collects these results along with valuative invariants that were identified in [7, Theorem 5.6] and that we will use heavily.
Theorem 2.5.
Each of the following is a valuative invariant:
  1. (1)
    the map $M\mapsto\mathcal{G}(M)$,
  2. (2)
    the map $M\mapsto\mathcal{G}(M^{*})$,
  3. (3)
    for any rank sequence $\underline{r}$, the map $M\mapsto g_{\underline{r}}(M)$,
  4. (4)
    the map $M\mapsto T(M;x,y)$,
  5. (5)
    for $i,j\in\mathbb{N}_{0}$, the map $M\mapsto t_{ij}$ where $t_{ij}$ is the coefficient of $x^{i}y^{j}$ in $T(M;x,y)$,
  6. (6)
    the map $M\mapsto|\mathcal{I}(M)|$, where $\mathcal{I}(M)$ is the set of independent sets of $M$,
  7. (7)
    for any $h,k,s_{h},s_{h+1},\ldots,s_{k}\in\mathbb{N}_{0}$ for which $0\leq h\leq k\leq r(M)$, the map $M\mapsto F_{h,k}(M;s_{h},\ldots,s_{k})$, where $F_{h,k}(M;s_{h},\ldots,s_{k})$ is the number of flags $F_{h}\subsetneq F_{h+1}\subsetneq\cdots\subsetneq F_{k}$ of flats of $M$ with $r(F_{i})=i$ and $s_{i}=|F_{i}|$ for all $i$ with $h\leq i\leq k$,
  8. (8)
    the map $M\mapsto f_{k}(M;s)$, where $f_{k}(M;s)$ is the number of flats $X$ of $M$ with $r(X)=k$ and $|X|=s$, and
  9. (9)
    the map $M\mapsto f_{k}(M;s,c)$, where $f_{k}(M;s,c)$ is the number of flats $X$ of $M$ with $r(X)=k$ and $|X|=s$ for which $M|X$ has $c$ coloops.
Item (8) is a special case of item (7). By combining item (2) with other items, it follows that the number of circuits of $M$ of a given size is a valuative invariant, as is the number of cyclic sets of a given size and nullity. The number of cyclic flats of $M$ of a given rank and size is a valuative invariant; set $c=0$ in item (9). The number of circuit-hyperplanes is a valuative invariant; in a rank-$r$ matroid $M$, it is $f_{r-1}(M;r,0)$. Item (6) is a well-known evaluation of the Tutte polynomial, as is the number of bases.
2.4. Extremal matroids of the polytope of unlabeled matroids
To treat isomorphism classes of matroids, rather than labeled matroids, Ferroni and Fink use the symmetrized indicator function $\mathbb{I}(X):\mathbb{R}^{n}\to\mathbb{R}$ of a subset $X$ of $\mathbb{R}^{n}$, which is given by
$\mathbb{I}(X)(x_{1},x_{2},\ldots,x_{n})=\frac{1}{n!}\sum_{\sigma\in S_{n}}\bar{\mathbb{I}}(X)(x_{\sigma(1)},x_{\sigma(2)},\ldots,x_{\sigma(n)})$
where $S_{n}$ is the symmetric group on $[n]$. (Note that dropping the bar over $\mathbb{I}$ indicates the symmetrized version.) It follows that if $M$ and $N$ are isomorphic but unequal matroids on $[n]$, then $\bar{\mathbb{I}}(\mathcal{P}(M))\neq\bar{\mathbb{I}}(\mathcal{P}(N))$ but $\mathbb{I}(\mathcal{P}(M))=\mathbb{I}(\mathcal{P}(N))$. Theorem 2.4 carries over to the setting of unlabeled matroids: when both $M$ and the nested matroids are treated as unlabeled, the same equality holds with the same coefficients. Of course, by considering matroids up to isomorphism, typically some terms merge. For instance, for the matroid $M$ in Figure 1, when considering $\bar{\mathbb{I}}(\mathcal{P}(M))$, there are two different nested matroids whose cyclic flats are, besides $\emptyset$ and $E(M)$, a set of size two and rank $1$ (i.e., $A$ or $B$), and these give different terms in Equation (2.1), each with coefficient $-1$; however, in $\mathbb{I}(\mathcal{P}(M))$, they give the same unlabeled nested matroid, namely, the nested matroid that is encoded by the lattice path $NENNEEEE$, and its coefficient is $-2$. Taking this into account, Theorem 2.4 implies that the symmetrized indicator function of any unlabeled matroid of rank $r$ on $n$ elements can be written in exactly one way as an integer linear combination of the symmetrized indicator functions of unlabeled nested matroids of rank $r$ on $n$ elements. By imposing a linear order on the set of all $\binom{n}{r}$ unlabeled nested matroids of rank $r$ on $n$ elements and considering the set of coordinate vectors of unlabeled matroids of rank $r$ on $n$ elements relative to these nested matroids, one gets a set of points in $\mathbb{R}^{\binom{n}{r}}$. The convex hull of this set of points is the polytope $\Omega_{r,n}$ of unlabeled matroids of rank $r$ on $n$ elements. Matroids $M$ that give a vertex in $\Omega_{r,n}$ are called extremal. The starting point for our work is the following result from [19, Lemma 4.3].
Theorem 2.6.
Fix integers $r$ and $n$ with $0\leq r\leq n$. Let $f_{1},f_{2},\ldots,f_{t}$ be real-valued valuative invariants. Let the set $S_{0}$ consist of one matroid of each isomorphism type among rank-$r$ matroids on $n$ elements, and, for each $i\in[t]$, set
$S_{i}=\{M\in S_{i-1}\,:\,f_{i}(M)\geq f_{i}(N)\text{ for all }N\in S_{i-1}\}.$
If all matroids in $S_{t}$ yield the same point in the polytope $\Omega_{r,n}$, then that point is a vertex of $\Omega_{r,n}$, so all matroids in $S_{t}$ are then extremal.
Theorem 2.6 motivates characterizing certain matroids by maximizing a sequence of real-valued valuative invariants. Such characterizations are the focus of this paper. Note that $f$ is valuative if and only if $-f$ is, so we can also minimize real-valued valuative invariants. In some settings maximizing is more natural, while in others minimizing is more natural, so we refer to the generic case as optimizing valuative invariants.
3. Perfect matroid designs
A perfect matroid design is a matroid $M$ in which flats of the same rank have the same size, that is, there are integers $\alpha_{0},\alpha_{1},\ldots,\alpha_{r(M)}$ so that $|F|=\alpha_{r(F)}$ for all flats $F$ of $M$. (See [16] and [28].) Examples include projective and affine geometries and their truncations, as well as paving matroids of rank $r$ on $[n]$ for which the hyperplanes are the blocks of a Steiner system $S(r-1,k,n)$ (i.e., a set of $k$-subsets of $[n]$, the blocks, for which each $(r-1)$-subset of $[n]$ is a subset of exactly one block).
Nonisomorphic perfect matroid designs can have the same sequence $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$; for instance, for many proper prime powers $q$, there are nonisomorphic projective planes of order $q$, and all have the sequence $(1,q+1,q^{2}+q+1)$. By [7, Example 3.4], all perfect matroid designs with the sequence $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$ have the same $\mathcal{G}$-invariant and so share all valuative invariants. Perfect matroid designs that have the sequence $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$ can be identified in the manner indicated in Theorem 2.6 by maximizing the number of flats of rank 00 with $\alpha_{0}$ elements and then, for each $i\in[r-1]$, minimizing the number of flats of rank $i$ and sizes other than $\alpha_{i}$. For $i\in[r]$, the flats of rank $i$ are cyclic if and only if $\alpha_{i}>\alpha_{i-1}+1$. Also, the number of rank-$(i+1)$ flats that contain a given rank-$i$ flat $F$ is $(\alpha_{r}-\alpha_{i})/(\alpha_{i+1}-\alpha_{i})$ since each element in $E(M)-F$ is in exactly one rank-$(i+1)$ flat that contains $F$. It follows that for a perfect matroid design $M$ and chain $C\in\mathcal{C}(\mathcal{Z}(M))$, the number of chains $C^{\prime}\in\mathcal{C}(\mathcal{Z}(M))$ with $C\subseteq C^{\prime}$ and specified ranks for the flats in $C^{\prime}$ depends only on $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$, the ranks of the flats in $C$, and the ranks specified for $C^{\prime}$. It then follows by induction, starting with the largest chains $C\in\mathcal{C}(\mathcal{Z}(M)$, that $\mu(C,\hat{1})$ depends only on $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$ and the ranks of the flats in $C$. Thus, two perfect matroid designs that share the sequence $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$ yield the same point in $\Omega_{r,n}$. This gives the following result.
Theorem 3.1.
Perfect matroid designs are extremal. In particular, projective and affine geometries, as well as their truncations, are extremal.
Since it will be useful for our work on direct sums in Section 8, we give another result that uses a valuative invariant to distinguish perfect matroid designs from other matroids. The valuative invariant used below is defined in part (7) of Theorem 2.5. The ideas below are essentially, if not explicitly, in [7, Example 3.4].
Theorem 3.2.
If there are perfect matroid designs that have the sequence $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$ of sizes of flats, then they are precisely the rank-$r$ matroids on $\alpha_{r}$ elements that maximize the valuative invariant $N\mapsto F_{0,r}(N;\alpha_{0},\alpha_{1},\ldots,\alpha_{r})$. In particular, for $r\geq 4$ and a prime power $q$, among all simple rank-$r$ matroids on $\frac{q^{r}-1}{q-1}$ elements, $\operatorname{PG}(r-1,q)$ is the unique matroid that maximizes the valuative invariant
$N\mapsto F_{0,r}(N;0,1,q+1,q^{2}+q+1,\ldots,\frac{q^{r}-1}{q-1}).$
Also, among simple rank-$r$ matroids on $q^{r-1}$ elements, $\operatorname{AG}(r-1,q)$ is the unique matroid that maximizes the valuative invariant $N\mapsto F_{0,r}(N;0,1,q,q^{2},\ldots,q^{r-1})$.
Proof.
Let $M$ be a rank-$r$ matroid on $\alpha_{r}$ elements with $\alpha_{0}$ loops. If $F$ is a rank-$i$ flat of $M$ with $|F|=\alpha_{i}$, then the maximum number of rank-$(i+1)$ flats of size $\alpha_{i+1}$ that contain $F$ is $\frac{\alpha_{r}-\alpha_{i}}{\alpha_{i+1}-\alpha_{i}}$. Thus,
$F_{0,r}(M;\alpha_{0},\alpha_{1},\ldots,\alpha_{r})\leq\prod_{i=0}^{r-1}\frac{\alpha_{r}-\alpha_{i}}{\alpha_{i+1}-\alpha_{i}}.$
Also, equality holds if and only if each point has size $\alpha_{1}$ and is contained only in lines of size $\alpha_{2}$, and each line is contained only in planes of size $\alpha_{3}$, etc., so $M$ is a perfect matroid design with sequence $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$. ∎
4. Bose-Burton geometries
Bose-Burton geometries are matroids of the form $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(k-1,q)$, that is, the matroid obtained by deleting from $\operatorname{PG}(r-1,q)$ one of its rank-$k$ flats. The next theorem, from Bose and Burton [9], is a counterpart, for projective geometries, of Turán’s theorem; just as Turán’s theorem treats subgraphs of $K_{n}$ that do not have another particular complete graph as a subgraph, this theorem treats restrictions of $\operatorname{PG}(r-1,q)$ that do not have another particular projective geometry over $\operatorname{GF}(q)$ as a restriction.
Theorem 4.1.
If $M$ is a restriction of $\operatorname{PG}(r-1,q)$ that has no restriction isomorphic to $\operatorname{PG}(m-1,q)$, then
$|E(M)|\leq\frac{q^{r}-q^{r-m+1}}{q-1}.$
Furthermore, equality holds if and only if $M$ is isomorphic to $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(r-m,q)$.
That is one of many attractive properties of Bose-Burton geometries. If $k=r-1$, then $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(k-1,q)$ is the affine geometry $\operatorname{AG}(r-1,q)$, which, since it is a perfect matroid design, is extremal. The result below from [1] shows that all other Bose-Burton geometries except perhaps $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(r-3,q)$ are extremal since the conditions (1)–(3) in Theorem 4.2 can be cast as optimizing valuative invariants.
Theorem 4.2.
Let $q$ be a prime power. Fix $r,k\in\mathbb{N}$ with $r\geq 4$ and $31\leq k\leq r-$ 3. Let $M$ be a simple rank-$r$ matroid on $(q^{r}-q^{k})/(q-1)$ elements that satisfies conditions (1)–(3).
  • (1)
    If $F$ is a rank-$(r-k+1)$ flat of $M$, then $|F|<(q^{r-k+1}-1)/(q-1)$.
  • (2)
    If $H$ is a hyperplane of $M$, then $|H|=(q^{r-1}-q^{i})/(q-1)$ where $i\in\{k,k-1\}$.
  • (3)
    If $F$ is a rank-$(r-2)$ flat of $M$, then $|F|=(q^{r-1}-q^{i})/(q-1)$ where $i$ is in $\{k,k-1,k-2\}$ and $i$ can be $k-2$ only if $k>1$.
Then $M$ is isomorphic to $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(k-1,q)$.
Theorem 4.5 completes the proof that Bose-Burton geometries are extremal: we show that $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(r-3,q)$ is extremal. The proof uses the result below from [4].
Theorem 4.3.
Let $r>3$ be an integer and let $q$ be a prime power. Any simple rank-$r$ matroid that has no minor isomorphic to $U_{2,q+2}$, the $(q+2)$-point line, and that has at least $q^{r-1}$ points is representable over $\operatorname{GF}(q)$.
To treat the remaining Bose-Burton geometries, $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(r-3,q)$, we address $r\geq 4$ and $r=3$ separately, in part since, when $r\geq 4$ there is just one projective geometry of rank $r$ and order $q$ (which must be a prime power), but when $r=3$ there can be many projective planes of order $q$ and it is unknown whether $q$ is limited to prime powers.
Theorem 4.4.
Fix an integer $q\geq 2$. Let $M$ be a rank-$3$ simple matroid with $q^{2}+q$ elements in which each line has $q$ or $q+1$ points. The number of $q$-point lines in $M$ is at least $q+1$, and if it is $q+1$, then $M=N\backslash x$ for some projective plane $N$ of order $q$ and element $x\in E(N)$. Also, single-element deletions of projective planes of order $q$ are extremal.
Proof.
If all lines that contain $e\in E(M)$ had $q+1$ points, then $|E(M)-\{e\}|=q^{2}+q-1$ would be divisible by $q$, but it is not. Thus, each element is in at least one $q$-point line. Along with the factorization $q^{2}+q=q(q+1)$, that implies that at least $q+1$ lines have $q$ points each, as claimed. The factorization $q^{2}+q=q(q+1)$ also implies that $M$ has $q+1$ lines with $q$ points each if and only if each element of $M$ is in one such line. In that case the $q$-point lines must be pairwise disjoint and so, along with $E(M)$, they form a modular cut. The corresponding single-element extension, say $N$, has $q^{2}+q+1$ points and all lines have $q+1$ points, from which it follows that $N$ is a projective plane of order $q$.
There may be nonisomorphic projective planes of order $q$ if $q>2$, so to complete the proof that they are extremal, we must show that all single-element deletions of projective planes of order $q>2$ give the same point in $\Omega_{3,q^{2}+q}$. For that, it suffices to show that from $q$ alone we get the nested matroids and their coefficients in the linear combination of indicator functions of nested matroids that gives $\mathbb{I}(\mathcal{P}(M))$. The lattice of cyclic flats of $M$ consists of $\emptyset$, $q+1$ lines that each have $q$ points, $q^{2}$ lines that each have $q+1$ points, and $E(M)$. Thus, the lattice $\mathcal{C}(\mathcal{Z}(M))$ has, besides $\hat{1}$,
  • $q+1$ elements $C=\{\emptyset,L,E(M)\}$ covered by $\hat{1}$ for which $|L|=q$; for each, $\mu(C,\hat{1})=-1$,
  • $q^{2}$ elements $C=\{\emptyset,L,E(M)\}$ covered by $\hat{1}$ for which $|L|=q+1$; for each, $\mu(C,\hat{1})=-1$,
  • the least chain, $C=\{\emptyset,E(M)\}$; given the values above, including $\mu(\hat{1},\hat{1})=1$, we have $\mu(C,\hat{1})=q^{2}+q$.
Thus, all single-element deletions of projective planes of order $q$ give the same point in $\Omega_{3,q^{2}+q}$. ∎
We now turn to ranks $4$ and greater. The counting is slightly harder than for $r=3$, but only one matroid, up to isomorphism, satisfies the conditions below, so it is automatically extremal. If $F$ and $X$ are rank-$(r-2)$ flats in $\operatorname{PG}(r-1,q)$ with $r(X\cap F)=r-4$, then $r(X\cup F)=r$, so no hyperplane of $\operatorname{PG}(r-1,q)$ that contains $F$ also contains $X$. Thus, $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(r-3,q)$ satisfies condition (b) below, which minimizes the valuative invariant $F_{r-2,r-1}(M;q^{r-3}+q^{r-4},q^{r-2})$ in part (7) of Theorem 2.5.
Theorem 4.5.
Fix a prime power $q$. Assume that $M$ is a simple matroid of rank $r\geq 4$ on $q^{r-1}+q^{r-2}$ elements that satisfies the following conditions:
  • (1)
    if $2\leq i<r$ and $F$ is a rank-$i$ flat of $M$, then $|F|\in\{q^{i-1},q^{i-1}+q^{i-2}\}$, and
  • (2)
    no rank-$(r-2)$ flat with $q^{r-3}+q^{r-4}$ elements is a subset of a hyperplane with $q^{r-2}$ elements.
Then the minimum number of $q^{r-2}$-point hyperplanes in $M$ is $q+1$. Furthermore, $M$ has $q+1$ such hyperplanes if and only if each element of $M$ is in one such hyperplane, and in that case, $M$ is isomorphic to $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(r-3,q)$. Thus, for any $r\geq 4$, the matroid $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(r-3,q)$ is extremal.
Proof.
To prove the lower bound of $q+1$ on the number of $q^{r-2}$-point hyperplanes, first note that $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(r-3,q)$ has $q+1$ such hyperplanes, so the minimum is at most $q+1$. Note that if, for some $i\geq 1$, the element $e$ is in a rank-$i$ flat $F$ with $q^{i-1}$ elements and $i<r-1$, then having all rank-$(i+1)$ flats that contain $F$ have $q^{i}+q^{i-1}$ elements would yield an integer solution $k$ of the equation $q^{r-1}+q^{r-2}=q^{i-1}+kq^{i}$, and so of $q^{r-i}+q^{r-i-1}=1+kq$, which has no integer solution; thus, $e$ is in some rank-$(i+1)$ flat with $q^{i}$ elements. Thus, by induction, each $e\in E(M)$ is in at least one $q^{r-2}$-point hyperplane. The lower bound on the number of $q^{r-2}$-point hyperplanes now follows since the fewest number of sets of size $q^{r-2}$ whose union is $E(M)$ is $q+1$ since $q^{r-1}+q^{r-2}=q^{r-2}(q+1)$. Since each element is in at least one $q^{r-2}$-point hyperplane, that factorization also shows that $M$ has $q+1$ such hyperplanes if and only if each element of $M$ is in one such hyperplane.
Now assume that $M$ has $q+1$ hyperplanes of size $q^{r-2}$, so each element of $M$ is in one such hyperplane. We next use Theorem 4.3 to show that $M$ is representable over $\operatorname{GF}(q)$. To show that $M$ has no $U_{2,q+2}$-minor, by the Scum Theorem it suffices to show that at most $q+1$ hyperplanes contain any rank-$(r-2)$-flat $F$ of $M$. Each element is in one $q^{r-2}$-point hyperplane, so at most one such hyperplane contains $F$. First assume that $|F|=q^{r-3}$ and that $i$ hyperplanes that contain $F$ have $q^{r-2}$ elements, so $i\in\{0,1\}$. The number of $(q^{r-2}+q^{r-3})$-point hyperplanes that contain $F$ is then
$\frac{q^{r-1}+q^{r-2}-q^{r-3}-i(q^{r-2}-q^{r-3})}{q^{r-2}}.$
Canceling common factors shows that $q$ divides $q^{2}+q-1-i(q-1)$, so $q$ divides $i-1$. Thus, $i=1$ and the displayed quantity simplifies to $q$, so $F$ is in $q+1$ hyperplanes, as needed. Now assume that $|F|=q^{r-3}+q^{r-4}$. By conditions (1) and (2), each hyperplane that contains $F$ has size $q^{r-2}+q^{r-3}$, so the number of such hyperplanes is
$\frac{q^{r-1}+q^{r-2}-(q^{r-3}+q^{r-4})}{q^{r-2}+q^{r-3}-(q^{r-3}+q^{r-4})}=q+1,$
so $F$ is contained in $q+1$ hyperplanes. Thus, by the Scum Theorem and Theorem 4.3, $M$ is representable over $\operatorname{GF}(q)$ and so is isomorphic to a restriction of $\operatorname{PG}(r-1,q)$.
Finally, condition (1) implies that $M$ has no flat $F$ for which $M|F$ is isomorphic to $\operatorname{PG}(2,q)$, so, by Theorem 4.1, $M$ is isomorphic to $\operatorname{PG}(r-1,q)\backslash\operatorname{PG}(r-3,q)$. ∎
5. Cycle matroids of complete graphs, and their truncations
To prove that $M(K_{r+1})$ is extremal, we will use the following characterization of $M(K_{r+1})$ from [8, Theorem 3.2], which does not have the form set out in Theorem 2.6.
Theorem 5.1.
Assume that a rank-$r$ simple matroid $M$ on (${r+1\choose 2}$ elements has
  • (1)
    (${r+1\choose 3}$ lines with three points,
  • (2)
    no $5$-point planes, (${r+1\choose 4}$ planes with six points, no planes with more than six points, and
  • (3)
    no rank-$4$ flats with more than ten points.
Then $M$ is isomorphic to the cycle matroid $M(K_{r+1})$.
As noted in the introduction, all criteria in the next result optimize valuative invariants, so this shows that $M(K_{r+1})$ is extremal, as conjectured in [19].
Theorem 5.2.
[…]
Proof.
In this proof, we call $3$-point lines large lines. The statement clearly holds for $r\leq 2$, so now assume that $r\geq 3$.
[…]
Let $M$ have $k$ large lines. The number of triples $(L,e,f)$, where $L$ is a large line and $e$ and $f$ are distinct points on $L$, is $6k$. Since there are at most $\binom{r+1}{2}$ options for $e$, and then at most $r-1$ options for $L$, and then two options for $f$, we get $26k\leq\binom{r+1}{2}(r-1)$ 2, which simplifies to $k\leq\binom{r+1}{3}$, as the bound in the theorem asserts.
Now assume that $M$ has $\binom{r+1}{3}$ large lines. The counting argument above shows that each element $e$ of $M$ is in $r-1$ large lines, any $h-1$ of which span a flat of rank $h$. Also, $E(M)$ is the union of the $r-1$ large lines $L_{1},L_{2},\ldots,L_{r-1}$ that contain $e$ along with the $\binom{r-1}{2}$ singletons $\{a_{i,j}\}=\operatorname{cl}(L_{i}\cup L_{j})-(L_{i}\cup L_{j})$ for distinct $i,j\in[r-1]$. We continue to use the notation $e,L_{1},L_{2},\ldots,L_{r-1}$, and $a_{i,j}$ in the next three paragraphs as we deduce more about the $6$-point planes of $M$.
We claim that, for each $i\in[r-1]$, the planes $\operatorname{cl}(L_{i}\cup L_{j})$, for $j\in[r-1]-\{i\}$, account for all $6$-point planes that contain $L_{i}$. Since $\operatorname{cl}(L_{i}\cup L_{j})=\operatorname{cl}(L_{i}\cup\{a_{i,j}\})$, each other plane that contains $L_{i}$ is $\operatorname{cl}(L_{i}\cup\{a_{s,t}\})$ for some $\{s,t\}\subseteq[r-1]-\{i\}$. Assume, for the sake of contradiction, that $P=\operatorname{cl}(L_{i}\cup\{a_{s,t}\})$ has six elements. For any $f\in(L_{s}\cup L_{t})-\{e\}$, the rank-$4$ flat $\operatorname{cl}(P\cup\{f\})$ would be $\operatorname{cl}(L_{i}\cup L_{s}\cup L_{t})$, and $P\cap(L_{s}\cup L_{t})=\{e\}$. Since
$|\operatorname{cl}(L_{i}\cup L_{s}\cup L_{t})-P|=4=|(L_{s}\cup L_{t})-\{e\}|,$
the plane $P$ would have to be $L_{i}\cup\{a_{s,t},a_{i,s},a_{i,t}\}$. However, that is impossible since $\operatorname{cl}(L_{i}\cup\{a_{s,t},a_{i,s},a_{i,t}\})$ is the rank-$4$ flat $\operatorname{cl}(L_{i}\cup L_{s}\cup L_{t})$.
We claim that, for all $\{i,j\}\subseteq[r-1]$, the flats $\operatorname{cl}(L_{i}\cup L_{j}\cup L_{k})$ for $k\in[r-1]-\{i,j\}$ account for all $10$-element rank-$4$ flats that contain $L_{i}\cup L_{j}$. Assume that $F$ is another such rank-$4$ flat, say $F=\operatorname{cl}(L_{i}\cup L_{j}\cup\{a_{s,t}\})$ for some $\{s,t\}\subseteq[r-1]-\{i,j\}$. Now $F\cap(L_{s}\cup L_{t})=\{e\}$ and for any $f\in(L_{s}\cup L_{t})-\{e\}$, the rank-$5$ flat $F^{\prime}=\operatorname{cl}(F\cup\{f\})$ would be $\operatorname{cl}(L_{i}\cup L_{j}\cup L_{s}\cup L_{t})$. The four points in $(L_{s}\cup L_{t})-\{e\}$ are in $F^{\prime}-F$, and $|F^{\prime}-F|=5$, so at least three of $a_{s,i}$, $a_{s,j}$, $a_{t,i}$, and $a_{t,j}$ are in $F$; since $L_{i}\cup L_{j}\subsetneq F$, that gives the contradiction that the rank-$4$ flat $F$ contains the rank-$5$ flat $\operatorname{cl}(L_{i}\cup L_{j}\cup L_{s}\cup L_{t})$.
We now rule out having a $6$-point plane $P$ with $P\cap(L_{1}\cup L_{2}\cup\cdots\cup L_{r-1})=\{e\}$. Assume, to the contrary, that there is such a $6$-point plane $P$. Consider $a_{i,j}\in P$. For any $f\in(L_{i}\cup L_{j})-\{e\}$, the rank-$4$ flat $F=\operatorname{cl}(P\cup\{f\})$ would contradict what we proved in the previous paragraph since it contains $L_{i}\cup L_{j}$ and five points that are in none of the lines $L_{k}$ for $k\in[r-1]$,
Thus, every $6$-point plane that contains $e$ contains at least two large lines that contain $e$. Any point in such a plane can be $e$, so each $6$-point plane of $M$ contains four large lines.
Let $k$ be the number of $6$-point planes of $M$. Consider the triples $(P,L,e)$ where $P$ is a $6$-point plane, $L$ is a large line in $P$, and $e\in L$. The number of such triples is $k\cdot 4\cdot 3=12k$ by the conclusion above. By focusing on $e$ first, and then $L$, and finally $P$, we see that the number of triples is $\binom{r+1}{2}(r-1)(r-2)$, where $r-2$ accounts for choosing a second line containing $e$ to determine the plane $P$. Thus, $12k=\binom{r+1}{2}(r-1)(r-2)$, so $k=\binom{r+1}{4}$.
This shows that the criteria in Theorem 5.1 hold, so $M$ is $M(K_{r+1})$. ∎
We use Theorem 5.2 to prove the next result, which gives a characterization of $M(K_{r+1})$ by optimizing valuative invariants that suits our work in Section 8 on direct sums.
Theorem 5.3.
Let $M$ be a simple matroid of rank $r\geq 2$ on $\binom{r+1}{2}$ elements. Assume that for each $i\in[r-1]$ and each rank-$i$ flat $F$, either $|F|=\binom{i+1}{2}$ or $|F|\leq\binom{i}{2}+1$. Call a rank-$i$ flat $F$ large if $|F|=\binom{i+1}{2}$, and call a flag $F_{0}\subsetneq F_{1}\subsetneq F_{2}\subsetneq\cdots\subsetneq F_{r}$ of flats large if each $F_{i}$ is large. The number of large flags is at most $(r+1)!/2$, and if the number of large flags is $(r+1)!/2$, then $M$ is $M(K_{r+1})$. Thus, $M(K_{r+1})$ is extremal.
Proof.
We first prove, by induction on $r$, that $M$ has at most $r+1$ large hyperplanes. An inequality that we use to prove the induction step holds for $r\geq 5$, so we check all cases $r<5$ first. If $r=2$, then $M$ is $U_{2,3}$, so the assertion holds. For $r=3$, since $|E(M)|=6$, each point is in at most two large lines; thus, if there are $k$ large lines, then the number of pairs $(e,L)$ where $L$ is a large line and $e\in L$ is, on the one hand, $3k$, and, on the other, at most $6\cdot 2$, so $3k\leq 12$, that is, $k\leq 4$, as claimed.
We now show that if $r=4$, so $|E(M)|=10$, then $M$ has at most five large planes. By inclusion/exclusion, the intersection of any two large planes $P$ and $P^{\prime}$ is a line with two or three elements. First assume that $P\cap P^{\prime}=\{a,b\}$. Thus, $P\cup P^{\prime}=E(M)$, so any other large plane $P^{\prime\prime}$ must intersect each of $P$ and $P^{\prime}$ in a large line; let $P^{\prime\prime}\cap P=L$ and $P^{\prime\prime}\cap P^{\prime}=L^{\prime}$. Now $P^{\prime\prime}=L\cup L^{\prime}$, so $L\cap L^{\prime}=\emptyset$, and so $\{a,b\}\cap(L\cup L^{\prime})=\emptyset$. However, only one large line in $P$ can be disjoint from $\{a,b\}$, and similarly for $P^{\prime}$, so there is at most one large plane besides $P$ and $P^{\prime}$, consistent with the claim. Now assume that any two large planes of $M$ intersect in a large line. By counting, any large line is in at most two large planes. Since any large plane $P$ has at most four large lines and each other large plane must intersect $P$ in one of those lines, and each such line is in at most one large plane besides $P$, there are at most five large planes, as needed. It is relevant for the last step of the proof to note that we have shown that the bound of five large planes is met only when each pair of those large planes intersects in a large line.
Now assume that $r\geq 5$ and that the bound on the number of large hyperplanes holds for matroids of rank $r-1$. Let $H_{1},H_{2},\ldots,H_{t}$ be the large hyperplanes of $M$. For any $\{i,j\}\subseteq[t]$, we have $|H_{i}\cup H_{j}|\leq|E(M)|=\binom{r+1}{2}$, so
$|H_{i}\cap H_{j}|$ $\,=|H_{i}|+|H_{j}|-|H_{i}\cup H_{j}|$
$\,\geq 2\binom{r}{2}-\binom{r+1}{2}$
$\,=\binom{r-1}{2}-1$
$\,>\binom{r-2}{2}+1$
where the last line holds since $r\geq 5$. Thus, $H_{i}\cap H_{j}$ must be a large rank-$(r-2)$ flat.
We next claim that the $t-1$ intersections $H_{i}\cap H_{t}$, for $i\in[t-1]$, are all different. Assume instead that $H_{i}\cap H_{t}=F=H_{j}\cap H_{t}$ where $1\leq i<j<t$. As just shown, $F$ is a large rank-$(r-2)$ flat. The sets $H_{i}-F,H_{j}-F,H_{t}-F$ partition $(H_{i}\cup H_{j}\cup H_{t})-F$, and each has size $\binom{r}{2}-\binom{r-1}{2}=r-1$, so
$|H_{i}\cup H_{j}\cup H_{k}|=\binom{r-1}{2}+3(r-1)=\binom{r+1}{2}+r-2,$
which gives the contradiction $|H_{i}\cup H_{j}\cup H_{t}|>|E(M)|$, and so proves the claim.
By the induction assumption applied to $M|H_{t}$, we get $t-1\leq r$, so $t\leq r+1$, as claimed. With that, the bound of $(r+1)!/2$ on the number of large flags $F_{0}\subsetneq F_{1}\subsetneq\cdots\subsetneq F_{r}$ follows since there are at most $r+1$ choices for $F_{r-1}$, which, as we showed, gives at most $r$ choices for $F_{r-2}$, and so on, and, ending with three options for $F_{1}$.
Assume that $M$ has $(r+1)!/2$ large flags. Then $M$ must have $r+1$ large hyperplanes, say $H_{1},H_{2},\ldots,H_{r+1}$, and each $H_{i}$ must contain $r$ large rank-$(r-2)$ flats, which, as we saw, are the intersections $H_{i}\cap H_{j}$ for $j\in[r+1]-\{i\}$. Applying that to each $M|H_{i}$ shows that the intersection of any three large hyperplanes of $M$ is a large rank-$(r-3)$ flat. By iterating this, we see that if $I\subsetneq[r+1]$, then $\bigcap_{i\in I}H_{i}$ is a large rank-$(r-|I|)$ flat. Thus, the intersections of $r-2$ large hyperplanes account for $\binom{r+1}{r-2}=\binom{r+1}{3}$ different large lines of $M$. Since $M$ has $\binom{r+1}{3}$ large lines, $M$ is $M(K_{r+1})$ by Theorem 5.2. ∎
The proof above gives another characterization of $M(K_{r+1})$ by optimizing valuative invariants: a matroid that satisfies the assumptions in Theorem 5.3 has at most $r+1$ large hyperplanes, and the only such matroid with $r+1$ large hyperplanes is $M(K_{r+1})$.
We now turn to truncations of $M(K_{r+1})$ of rank at least $5$. In [13], de Mier and Noy showed that these truncations are Tutte unique. For that, they first developed a numerical characterization of these matroids similar to that of $M(K_{r+1})$ in Theorem 5.1, adding an assumption on the number of independent sets. We characterize truncations of $M(K_{r+1})$ of rank at least $5$ as optimizing certain valuative invariants and so show that these matroids are extremal. The proof below does not rest on prior characterizations. Theorem 5.4 gives yet another characterization of $M(K_{r+1})$, although it is limited to $r\geq 5$.
We use $\operatorname{Tr}^{i}(N)$ for the matroid obtained from $N$ by truncating $i$ times, where $i\leq r(N)$. Thus, $\operatorname{Tr}^{i}(N)$ has rank $r(N)-i$, and its independent sets are those of $N$ that have size at most $r(N)-i$.
Theorem 5.4.
Fix $r,k\in\mathbb{N}\cup\{0\}$ with $r-k\geq 5$. Let $M$ be a simple rank-$(r-k)$ matroid on $\binom{r+1}{2}$ elements in which $|F|\leq\binom{r(F)+1}{2}$ for all flats $F\neq E(M)$ and no planes have five elements. Then at most $\binom{r+1}{3}$ lines of $M$ have three elements. If $\binom{r+1}{3}$ lines of $M$ have three elements, then $M$ has at most as many independent sets as $\operatorname{Tr}^{k}(M(K_{r+1}))$. If, in addition, $M$ has the same number of independent sets as $\operatorname{Tr}^{k}(M(K_{r+1}))$, then $M$ is $\operatorname{Tr}^{k}(M(K_{r+1}))$. Thus, $\operatorname{Tr}^{k}(M(K_{r+1}))$ is extremal.
Proof.
[…]
We next identify the points and large lines of $M$ with those of $M(K_{r+1})$. Fix a point, which we will call $a_{r,r+1}$, and let $L_{1},L_{2},\ldots,L_{r-1}$ be the large lines that contain $a_{r,r+1}$. For integers $i$ and $j$ with $1\leq i<j<r$, let the singleton set $\operatorname{cl}(L_{i}\cup L_{j})-(L_{i}\cup L_{j})$ be $\{a_{i,j}\}$. Arbitrarily label the points on $L_{1}-\{a_{r,r+1}\}$ as $a_{1,r}$ and $a_{1,r+1}$. For each $i$ with $2\leq i<r$, let $a_{i,r}$ be the point in $L_{i}\cap\operatorname{cl}(\{a_{1,r},a_{1,i}\})$, and let $a_{i,r+1}$ be the point in $L_{i}\cap\operatorname{cl}(\{a_{1,r+1},a_{1,i}\})$, which is $L_{i}-\{a_{r,r+1},a_{i,r}\}$. To show that this labeling agrees with the large lines of $M(K_{r+1})$, we want to show that
  • (a)
    for $1<i<j<r$, both $\{a_{i,r},a_{j,r},a_{i,j}\}$ and $\{a_{i,r+1},a_{j,r+1},a_{i,j}\}$ are lines,
  • (b)
    for $1\leq i<j<k<r$, the set $\{a_{i,j},a_{i,k},a_{j,k}\}$ is a line,
and that we have identified all large lines of $M$. For the set $\{a_{i,r},a_{j,r},a_{i,j}\}$ in assertion (a), note that the plane $\operatorname{cl}(\{a_{1,r},a_{i,r},a_{j,r}\})$ is contained in the rank-$4$ flat $\operatorname{cl}(L_{1}\cup L_{i}\cup L_{j})$ and it contains $a_{1,i}$ and $a_{1,j}$ but none of $L_{1}$, $L_{i}$, and $L_{j}$. Therefore $a_{i,j}$ must be the sixth point of the plane $\operatorname{cl}(\{a_{1,r},a_{i,r},a_{j,r}\})$. Thus, the plane $\operatorname{cl}(\{a_{1,r},a_{i,r},a_{j,r}\})$ intersects the plane $\operatorname{cl}(L_{i}\cup J_{j})$ in the set $\{a_{i,r},a_{j,r},a_{i,j}\}$, which then must be a line, as needed. The argument to show that $\{a_{i,r+1},a_{j,r+1},a_{i,j}\}$ is a line is similar. For assertion (b), by assertion (a) and the labeling, the plane $P_{r}=\operatorname{cl}(\{a_{i,j},a_{j,k},a_{j,r}\})$ contains both $a_{i,r}$ (which is in $\operatorname{cl}(\{a_{i,j},a_{j,r}\})$) and $a_{k,r}$ (which is in $\operatorname{cl}(\{a_{j,k},a_{j,r}\})$), and so $P_{r}$ contains $a_{i,k}$, which is in $\operatorname{cl}(\{a_{i,r},a_{k,r}\})$; likewise, the plane $P_{r+1}=\operatorname{cl}(\{a_{i,j},a_{j,k},a_{j,r+1}\})$ contains both $a_{i,r+1}$ and $a_{k,r+1}$, and so it contains $a_{i,k}$. Since the two distinct planes $P_{r}$ and $P_{r+1}$ intersect in $\{a_{i,j},a_{i,k},a_{j,k}\}$, that set is a line. Finally, the $r-1$ large lines $L_{1},L_{2},\ldots,L_{r-1}$ that contain $a_{r,r+1}$, the $2\binom{r-1}{2}$ other large lines in the planes $\operatorname{cl}(L_{i}\cup L_{j})$, and the $\binom{r-1}{3}$ large lines identified in item (b) account for all $\binom{r+1}{3}$ lines of $M$. Thus, we can identify the elements and large lines of $M$ with those of $\operatorname{Tr}^{k}(M(K_{r+1}))$. With that identification, we can now assume that $M$ and $\operatorname{Tr}^{k}(M(K_{r+1}))$ have the same ground set.
We next show that sets that are independent in $M$ are independent in $\operatorname{Tr}^{k}(M(K_{r+1}))$; equivalently, every circuit $C$ of $\operatorname{Tr}^{k}(M(K_{r+1}))$ is dependent in $M$. This holds if $C$ spans $\operatorname{Tr}^{k}(M(K_{r+1}))$ since then $|C|>r(M)$. Now assume that $|C|\leq r-k$, so $C$ is the edge set of a cycle of $K_{r+1}$. In $K_{r+1}$, fix an edge $e$ in $C$, let $v$ run through the vertices in that cycle that are not incident with $e$, and consider the $3$-cycles determined by $e$ and $v$: we see that $C$ is in the closure, say $F$, of $|C|-2$ large lines of $\operatorname{Tr}^{k}(M(K_{r+1}))$ that contain $e$. By the identification above, we have $C\subseteq F$ in $M$. Now $r_{M}(F)=|C|-1$, so $C$ is dependent in $M$.
It follows that $M$ has at most as many independent sets as $\operatorname{Tr}^{k}(M(K_{r+1}))$, and if $M$ and $\operatorname{Tr}^{k}(M(K_{r+1}))$ have the same number of independent sets, then they have the same independent sets and so are the same matroid. ∎
6. Dowling geometries
We first recall the idea behind Dowling geometries. Let $\{b_{1},b_{2},\ldots,b_{r}\}$ be a basis of the projective geometry $\operatorname{PG}(r-1,q)$ over the finite field $\operatorname{GF}(q)$. Let $Q_{r}(\operatorname{GF}^{*}(q))$ be the restriction of $\operatorname{PG}(r-1,q)$ to the set $D=\bigcup_{\{i,j\}\subseteq[r]}\operatorname{cl}(\{b_{i},b_{j}\})$. One can check that the sets $F_{I}=\bigcup_{\{i,j\}\subseteq I}\operatorname{cl}(\{b_{i},b_{j}\})$, for $I\subseteq[r]$ with $|I|\geq 2$, are the non-singleton modular flats of $Q_{r}(\operatorname{GF}^{*}(q))$, that is, the flats $F$ for which $r(F)+r(X)=r(F\cup X)+r(F\cap X)$ for all flats $X$. Furthermore, the $3$-point lines in the flats $F_{I}$ with $|I|=3$ geometrically encode the group table of the multiplicative group $\operatorname{GF}^{*}(q)$ of units of $\operatorname{GF}(q)$. Dowling [17] showed how to replace the group $\operatorname{GF}^{*}(q)$ by an arbitrary finite group.
There are many ways to define Dowling geometries, including by group-labeled partial partitions of $[r]$, as in Dowling [17], and by group-labelled graphs, as in [23]. We will use cyclic flats and their ranks. A matroid is determined by its cyclic flats and their ranks. In a matroid $M$ with no loops, each connected component of a cyclic flat $F$ with $|F|>1$ is a connected flat, that is, $M|X$ is connected. We first give the connected flats of Dowling geometries, and use them to give the cyclic flats.
Let $G$ be a finite group, written with multiplicative notation. Fix $r\geq 3$. The ground set of the rank-$r$ Dowling geometry over $G$, denoted $Q_{r}(G)$, is
$\{b_{1},b_{2},\ldots,b_{r}\}\cup\{a_{(i,j)}\,:\,a\in G\text{ and }1\leq i<j\leq r\}.$
Hence $Q_{r}(G)$ has $r+|G|\binom{r}{2}$ elements, among which are $\binom{r}{2}$ copies of each element of $G$, each one indexed by a pair $(i,j)$ with $1\leq i<j\leq r$. For $i>j$ we set $a_{(i,j)}=(a^{-1})_{(j,i)}$, where $a^{-1}$ is the inverse of $a$ in $G$. The connected flats $F$ with $|F|>1$ are of two types:
  • for $I\subseteq[r]$ with $|I|\geq 2$, the large flat
    $\{b_{i}\,:\,i\in I\}\cup\{a_{(i,j)}\,:\,a\in G\text{ and }\{i,j\}\subseteq I\}$
    has rank $|I|$ and support $I$;
  • for a subset $J=\{j_{1}<j_{2}<\cdots<j_{t}\}$ of $[r]$ with $t\geq 3$ and (not necessarily distinct) $a_{2},a_{3},\ldots,a_{t}\in G$, the small flat
[…]
    has rank $t-1$ and support $J$.
Thus, a large flat with support $I$ has size $|I|+|G|\binom{|I|}{2}$, and a small flat with support $J$ has size $\binom{|J|}{2}$. The cyclic flats of $Q_{r}(G)$ are the unions of at most one large flat and any number of small flats for which the supports are pairwise disjoint; the rank of a cyclic flat is the sum of the ranks of the large and small flats that it contains. Figure 3 illustrates the construction with $Q_{3}(G)$ where $G$ is the sign group, $\{1,-1\}$. In that example, the set $I=\{1,2\}$ gives the large line $\{b_{1},b_{2},1_{(1,2)},-1_{(1,2)}\}$, and the set $J=[3]$ and the group elements $a_{2}=a_{3}=-1$ give the dotted line $\{-1_{(1,2)},1_{(2,3)},-1_{(1,3)}\}$.
$-1_{(1,2)}$$b_{1}$$1_{(1,2)}$$-1_{(2,3)}\,\,\,\,$$b_{2}$    $\,\,\,\,1_{(2,3)}$$-1_{(1,3)}$$b_{3}$$1_{(1,3)}$
Figure 3. The Dowling geometry $Q_{3}(G)$ where $G$ is the sign group, $\{1,-1\}$.
[…]
We will use the following characterization of Dowling geometries. The case of $r\geq 4$ was proven in [5]; the extension to $r=3$, along with a different proof, appeared in [3], which also gave a geometric criterion for distinguishing rank-$3$ Dowling geometries based on groups from those based on other quasigroups. That criterion cannot be cast in terms of valuative invariants since all Dowling geometries of the same rank based on groups (or, if $r=3$, quasigroups) of the same order have the same $\mathcal{G}$-invariant [7, Proposition 4.8].
Theorem 6.1.
A simple matroid $M$ of rank $r\geq 3$ is a Dowling geometry (based on a quasigroup if $r=3$) if and only if $M$ has elements $b_{1},b_{2},\dots,b_{r}$ that satisfy the following properties:
  1. (1)
    $E(M)$ is the union of the $\binom{r}{2}$ lines $\operatorname{cl}(\{b_{i},b_{j}\})$, for $\{i,j\}\subsetneq[r]$,
  2. (2)
    $|\operatorname{cl}(\{b_{i},b_{j}\})|>2$ for all $\{i,j\}\subsetneq[r]$, and
  3. (3)
    if $\{i,j,k\}\subsetneq[r]$, $e\in\operatorname{cl}(\{b_{i},b_{j}\})-\{b_{i},b_{j}\}$, and $f\in\operatorname{cl}(\{b_{i},b_{k}\})-\{b_{i},b_{k}\}$, then $|\operatorname{cl}(\{e,f\})|>2$.
The next result characterizes Dowling geometries of rank $r\geq 3$ other than $M(K_{r+1})$ by optimizing a sequence of valuative invariants. Since all Dowling geometries of rank $r$ based on groups (or quasigroups if $r=3$) of order $g$ have the same $\mathcal{G}$-invariant, we cannot distinguish between them via valuative invariants, but they are extremal since, as we show in Theorem 6.4 below, they give the same point in the polytope of unlabeled matroids.
Theorem 6.2.
[…]
  • (a)
    The matroid $M$ has at most $r$ large hyperplanes.
  • (b)
    If $M$ has $r$ large hyperplanes, then $M$ has a basis $B$ for which $\operatorname{cl}(X)$ is large for all $X\subseteq B$.
  • (c)
    If $M$ has $r$ large hyperplanes, then the greatest number of $3$-point lines in $M$ is $\binom{r}{3}g^{2}$, and any such matroid $M$ with $\binom{r}{3}g^{2}$ three-point lines is a rank-$r$ Dowling geometry based on a group (or quasigroup if $r=3$) of order $g$.
Proof.
We first prove conclusion (a) by induction on $r$. Let $M$ have the greatest possible number of large hyperplanes, and let $H_{1},H_{2},\ldots,H_{t}$ be the large hyperplanes of $M$. Since rank-$r$ Dowling geometries have $r$ large hyperplanes, we have $t\geq r$.
First assume that $r=3$. If $H_{i}\cap H_{j}=\emptyset$ for some $\{i,j\}\subsetneq[3]$, then we would get the contradiction $|H_{1}\cup H_{2}\cup H_{3}|\geq 2(g+2)+g=3g+4>|E(M)|$. The same contradiction (with different counting) would occur if $H_{1}\cap H_{2}\cap H_{3}\neq\emptyset$. Thus, $H_{i}\cap H_{j}$ is a singleton for all $\{i,j\}\subsetneq[3]$, while $H_{1}\cap H_{2}\cap H_{3}=\emptyset$. Inclusion/exclusion now gives $|H_{1}\cup H_{2}\cup H_{3}|=3g+3$, so $H_{1}\cup H_{2}\cup H_{3}=E(M)$. If $t>3$, then, since $|H_{i}\cap H_{4}|\leq 1$ for all $i\in[3]$ while $|H_{4}|>3$, we would get the contradiction $|H_{1}\cup H_{2}\cup H_{3}\cup H_{4}|>|E(M)|$. Thus, $t=3$. Also, the union of the three singletons $H_{i}\cap H_{j}$, for $\{i,j\}\subsetneq[3]$, is a basis, and the line spanned by any two of its elements is large. Thus, conclusions (a) and (b) hold when $r=3$.
An inequality in the general case requires $r>4$, so we next treat $r=4$. In that case, no line $L$ can be in three large planes, for otherwise we would have the contradiction
$|E(M)|\geq 3(3g+3-|L|)+|L|=9g+9-2|L|>6g+4=|E(M)|.$
By inclusion/exclusion and what we just proved, the least that $|H_{1}\cup H_{2}\cup H_{3}\cup H_{4}|$ could be would occur if each intersection $H_{i}\cap H_{j}$, for $\{i,j\}\subsetneq[4]$, is a large line, and each intersection $H_{i}\cap H_{j}\cap H_{k}$, for $\{i,j,k\}\subsetneq[4]$, is a singleton. Since the intersection of three large lines in a plane is empty, the intersection of all four planes is empty, so
$|H_{1}\cup H_{2}\cup H_{3}\cup H_{4}|\geq 4(3g+3)-6(g+2)+4=6g+4=|E(M)|.$
It follows that $H_{1}\cup H_{2}\cup H_{3}\cup H_{4}=E(M)$, that each intersection $H_{i}\cap H_{j}$, for $\{i,j\}\subsetneq[4]$, is a large line, and that each intersection $H_{i}\cap H_{j}\cap H_{k}$, for $\{i,j,k\}\subsetneq[4]$, is a singleton. Also, the union of the four singleton sets $H_{i}\cap H_{j}\cap H_{k}$, for $\{i,j,k\}\subsetneq[4]$, is a basis $B$ that satisfies conclusion (b) above. If $t>4$, then the same argument applied to $H_{1}$, $H_{2}$, $H_{3}$, and $H_{5}$ would show that $H_{5}$ must contain the large lines in $H_{1}\cup H_{2}\cup H_{3}$ that are not among $H_{i}\cap H_{j}$ for $\{i,j\}\subsetneq[3]$; this forces $H_{5}$ to be $H_{4}$. Thus, $t=4$, so both conclusion (a) and (b) hold.
Now assume that $r>4$ and that the bound on the number of large hyperplanes holds for lower ranks. We claim that, for $1\leq i<j\leq t$, the flat $H_{i}\cap H_{j}$ is a large rank-$(r-2)$ flat. Since $|H_{i}\cup H_{j}|\leq r+g\binom{r}{2}$, we have
$|H_{i}\cap H_{j}|$ $\,=|H_{i}|+|H_{j}|-|H_{i}\cup H_{j}|$
$\,\geq 2\Bigl{(}r-1+g\binom{r-1}{2}\Bigr{)}-\Bigr{(}r+g\binom{r}{2}\Bigl{)}$
$\,=g\Bigl{(}2\binom{r-1}{2}-\binom{r}{2}\Bigl{)}+\,r-2$
$\,=g\Bigl{(}\binom{r-2}{2}-1\Bigl{)}+\,r-2$
$\,>g\binom{r-3}{2}+r-2,$
(the final inequality requires $r>4$). Thus, $|H_{i}\cap H_{j}|$ exceeds the bound on the size of rank-$(r-3)$ flats by at least two, so $H_{i}\cap H_{j}$ is a large rank-$(r-2)$ flat, as claimed.
We next claim that the $t-1$ intersections $H_{i}\cap H_{t}$, for $i\in[t-1]$, are all different. Assume instead that $H_{i}\cap H_{t}=F=H_{j}\cap H_{t}$ where $1\leq i<j<t$. As just shown, $F$ is a large rank-$(r-2)$ flat. The sets $H_{i}-F,H_{j}-F,H_{t}-F$ partition $(H_{i}\cup H_{j}\cup H_{t})-F$, and each has size $r-1+g\binom{r-1}{2}-(r-2)-g\binom{r-2}{2}=1+g(r-2)$, so
$|H_{i}\cup H_{j}\cup H_{k}|$ $\,=r-2+g\binom{r-2}{2}+3(1+g(r-2))$
$\,=r+1+g\binom{r-1}{2}+2g(r-2)$
$\,=r+1+g\binom{r}{2}+g(r-3),$
which gives the contradiction $|H_{i}\cup H_{j}\cup H_{t}|>|E(M)|$, and so proves the claim.
The induction hypothesis applied to $M|H_{r}$ now gives $t-1\leq r-1$, so $t\leq r$, and so conclusion (a) holds.
For the induction step in the proof that a basis $B=\{b_{1},b_{2},\ldots,b_{r}\}$ satisfying statement (b) exists, it will be useful to know several properties of such bases. By counting, the union of the $\binom{r}{2}$ large lines $\operatorname{cl}(\{b_{i},b_{j}\})$ is $E(M)$. We claim that if $L$ is a line of $M$ with $|L|\geq 3$, then $L\subseteq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$ for some $\{i,j,k\}\subseteq[r]$. Assume that this is false. Then $L\cap B=\emptyset$ since if $b_{i}\in L$ and $\operatorname{cl}(\{b_{j},b_{k}\})$ contains another element of $L$, then $L\subseteq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$. Similarly, for any two elements of $L$, the $2$-subsets of $B$ whose closures contain those elements are disjoint. Since $|L|\geq 3$, there are disjoint $2$-subsets $B_{1},B_{2},B_{3}$ of $B$ and elements $e_{1},e_{2},e_{3}\in L$ with $e_{i}\in\operatorname{cl}(B_{i})$ for $i\in[3]$. Say $B_{3}=\{b_{h},b_{k}\}$. Since $e_{i}\in\operatorname{cl}(B_{i})\subseteq\operatorname{cl}(B-\{b_{k}\})$ for $i\in[2]$, we get $L\subseteq\operatorname{cl}(B-\{b_{k}\})$; thus, $b_{k}\in\operatorname{cl}(B-\{b_{k}\})$ since $\{b_{h},e_{3}\}\subseteq\operatorname{cl}(B-\{b_{k}\})$. However $b_{k}\not\in\operatorname{cl}(B-\{b_{k}\})$ since $B$ is a basis. This contradiction proves our claim.
For any large line $L$ of $M$, we have $L\subseteq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$ for some $\{i,j,k\}\subseteq[r]$, and, as shown in the second paragraph of the proof, the large plane $\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$ has three large lines, $\operatorname{cl}(\{b_{i},b_{j}\})$, $\operatorname{cl}(\{b_{i},b_{k}\})$, and $\operatorname{cl}(\{b_{j},b_{k}\})$. Thus, the large lines $\operatorname{cl}(\{b_{i},b_{j}\})$, for $\{i,j\}\subsetneq[r]$, account for all large lines of $M$. Thus, when a basis $B=\{b_{1},b_{2},\ldots,b_{r}\}$ as in conclusion (b) exists, it is unique.
We now prove conclusion (b) by induction on $r$. This was shown above for $r\in\{3,4\}$. For the induction step, assume that $r>4$ and that conclusion (b) holds for matroids of rank less than $r$. Thus, conclusion (b) holds for each restriction $M|H_{i}$ with $i\in[r]$, as well as for $M|(H_{i}\cap H_{j})$ when $\{i,j\}\subsetneq[r]$. Thus, $M|H_{1}$ has a basis $\{b_{1},b_{2},\ldots,b_{r-1}\}$ that satisfies conclusion (b). Now $M|H_{2}$ also has such a basis, as does $M|(H_{1}\cap H_{2})$. By the uniqueness of the basis of $M|(H_{1}\cap H_{2})$ satisfying conclusion (b) and noting that the large lines of $M|(H_{1}\cap H_{2})$ are large lines of $M|H_{1}$ and $M|H_{2}$ as well, it follows that, up to relabeling, the basis of $M|H_{2}$ that satisfies conclusion (b) is $\{b_{2},b_{3},\ldots,b_{r-1},b_{r}\}$ for some element $b_{r}$. We claim that the basis of $M$ that satisfies conclusion (b) is $B=\{b_{1},b_{2},\ldots,b_{r}\}$. To prove this, it suffices to show that the line $\operatorname{cl}(\{b_{1},b_{r}\})$ is large since any other line spanned by two elements of $B$ is in at least one of $M|H_{1}$ and $M|H_{2}$ and so is large. Consider the hyperplane $H=\operatorname{cl}(B-\{b_{2}\})$. We know that the lines $\operatorname{cl}(\{b_{i},b_{j}\})$, for $\{i,j\}\subsetneq[r]-\{2\}$ and $\{i,j\}\neq\{1,r\}$, are large since each is a large line of at least one of $M|H_{1}$ and $M|H_{2}$; thus, $|H|\geq r-1+g\bigl{(}\binom{r-1}{2}-1\bigl{)}$, so $H$ is a large hyperplane. By the induction hypothesis, $H$ has a basis that satisfies conclusion (b), and, as we showed above, that basis accounts for all large lines of $M|H$. Thus, that basis must be $B-\{b_{2}\}$, and so the line $\operatorname{cl}(\{b_{1},b_{r}\})$ is large, as needed. Thus, conclusion (b) holds.
We now turn to conclusion (c). As shown above, for the basis $B=\{b_{1},b_{2},\ldots,b_{r}\}$ in conclusion (b) and any $3$-point line $L$ of $M$, there is a subset $\{i,j,k\}$ of $[r]$ for which $L\subseteq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$. Note that no $3$-point line contains an element of $B$. The bound $\binom{r}{3}g^{2}$ follows since $\binom{r}{3}$ is the number of large planes $\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$, and if $L\subseteq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$, then $L$ is determined by its intersections with any two of the long lines in $\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$, each of which has $g$ elements that are not in $B$.
Now assume that the number of $3$-point lines in $M$ is $\binom{r}{3}g^{2}$. Properties (1) and (2) of Theorem 6.1 follow from conclusion (b). Since $M$ has $\binom{r}{3}g^{2}$ lines with three points, each choice of a subset $\{i,j,k\}\subseteq[r]$ along with elements $e\in\operatorname{cl}(\{b_{i},b_{j}\})-\{b_{i},b_{j}\}$ and $f\in\operatorname{cl}(\{b_{i},b_{k}\})-\{b_{i},b_{k}\}$ must give a $3$-point line $\operatorname{cl}(\{e,f\})$. Thus, property (3) of Theorem 6.1 also holds, so, as claimed, $M$ is a Dowling geometry. ∎
The theorem and proof above easily yield the following alternative characterization of Dowling geometries by optimizing valuative invariants. This characterization covers the case $r=2$ also, which is generally not considered since the group plays no significant role because $Q_{2}(G)$ is just the $(g+2)$-point line, but in Section 8 we will see that including $r=2$ gives us a slightly stronger result about direct sums.
Theorem 6.3.
[…]
Proof.
The statement trivially holds if $r=2$, so assume that $r>3$. Dowling geometries in $\mathcal{C}$ have $r!\,(g+2)/2$ large flags since a flag includes one of the $r$ large hyperplanes, and one of the $r-1$ large rank-$(r-2)$ flats that it contains, and so on through one of the three large lines in a large plane, and then one of the $g+2$ points in that large line. Furthermore, since, by Theorem 6.2, matroids in $\mathcal{C}$ have at most $r$ large hyperplanes, each of which contains at most $r-1$ large rank-$(r-2)$ flats, etc., if a matroid in $\mathcal{C}$ has $r!\,(g+2)/2$ large flags, then it must have $r$ large hyperplanes. With that, the conclusions of Theorem 6.2 give the result. ∎
The next result completes the proof that Dowling geometries are extremal.
Theorem 6.4.
Fix integers $r\geq 3$ and $g\geq 2$. All Dowling geometries of rank $r$ based on groups (or quasigroups if $r=3$) of order $g$ yield the same point in the polytope $\Omega_{r,r+g\binom{r}{2}}$ of unlabeled matroids. Thus, Dowling geometries are extremal.
Proof.
Let $G$ have order $g$. To prove the result, we show that everything involved in finding the point in the polytope $\Omega_{r,r+g\binom{r}{2}}$ given by $Q_{r}(G)$ uses only $r$ and $g$.
The connected components of a cyclic flat $F\in\mathcal{Z}(Q_{r}(G))$ are connected non-singleton flats, and at most one is a large flat. Call the components large or small according to whether they are large or small as flats. Let $[F]$, the type of $F$, be $(S,T)$ where $T$ is the set of supports of the small components of $F$, and if $F$ has a large component, then $S$ is its support, otherwise $S=\emptyset$. The type $[C]$ of a chain $C\in\mathcal{C}(\mathcal{Z}(Q_{r}(G)))$ is the set of the types of its members. The type $[F]=(S,T)$ gives $r(F)$ and $|F|$:
$r(F)=|S|+\sum_{X\in T}(|X|-1)\quad\text{ and }\quad|F|=|S|+g\,\binom{|S|}{2}+\sum_{X\in T}\binom{|X|}{2}.$
Thus, from the type $[C]$ of $C\in\mathcal{C}(\mathcal{Z}(M))$, we get the isomorphism type of the nested matroid that $C$ contributes to Equation (2.1) and its symmetrized counterpart.
By the description of the connected and cyclic flats of $Q_{r}(G)$ in the third paragraph of this section, for each type $(S,T)$, the number of cyclic flats of $Q_{r}(G)$ of type $(S,T)$ depends only on $g$ and $(S,T)$, specifically, it is $g^{t}$ where $t=\sum_{X\in T}(|X|-1)$ since such a flat is completely determined by $(S,T)$ along with, for each $H\in T$, say with least element $h$, the choice of one of $g$ elements $a_{(h,k)}\in\operatorname{cl}(\{b_{h},b_{k}\})-\{b_{h},b_{k}\}$ for each $k\in H-\{h\}$.
We claim that, for a proper cyclic flat $F$ of $Q_{r}(G)$ with $[F]=(S,T)$, the number of cyclic flats $F^{\prime}$ that cover $F$ in $\mathcal{Z}(Q_{r}(G))$ and that have a given type, say $(S^{\prime},T^{\prime})$, also depends just on $[F]$, $g$, and $(S^{\prime},T^{\prime})$. This follows by examining the covers $F^{\prime}$ of $F$.
  • If $S=\emptyset$ and neither $i$ nor $j$ is in any set in $T$, then one cover has type ($(\{i,j\},T)$.
  • If $S\neq\emptyset$ and $i$ is in neither $S$ nor any set in $T$, then one cover has type $(S\cup\{i\},T)$.
  • For each $H\in T$, one cover has type $(S\cup H,T-\{H\})$.
  • If $H=\{i,j,k\}\subseteq[r]$ and $H$ is disjoint from $S$ and the sets in $T$, then $g^{2}$ covers have type $(S,T\cup\{H\})$ since choosing one $a_{(i,j)}$ in $\operatorname{cl}(\{b_{i},b_{j}\})-\{b_{i},b_{j}\}$ and one $a_{(i,k)}$ in $\operatorname{cl}(\{b_{i},b_{k}\})-\{b_{i},b_{k}\}$ determines the cover.
  • If $\{H,H^{\prime}\}\subseteq T$, then $g$ covers have type $(S,(T-\{H,H^{\prime}\})\cup\{H\cup H^{\prime}\})$ since choosing one $a_{(h,h^{\prime})}\in\operatorname{cl}(\{b_{h},b_{h^{\prime}}\})-\{b_{h},b_{h^{\prime}}\}$, for one $h\in H$ and $h^{\prime}\in H^{\prime}$, determines the cover.
  • If $H\in T$ and $i$ is in neither $S$ nor any set of $T$, then $g$ covers of $F$ have type $(S,(T-\{H\})\cup\{H\cup\{i\})$ since choosing one $a_{(h,i)}\in\operatorname{cl}(\{b_{h},b_{i}\})-\{b_{h},b_{i}\}$, for one $h\in H$, determines the cover.
It follows that the number of chains in $\mathcal{Z}(Q_{r}(G))$ with a given type depends only on $r$ and $g$. Also, for a given chain in $\mathcal{Z}(Q_{r}(G))$, the number chains that contain it, and the multiset of their types, depends only on $r$ and $g$. It then follows by induction on the number of flats in chains, where maximal chains give the base case, that for any chain $C\in\mathcal{C}(\mathcal{Z}(M))$, the Möbius value $\mu(C,\hat{1})$ is determined by the type $[C]$. Thus, the point in the polytope $\Omega_{r,r+g\binom{r}{2}}$ that $Q_{r}(G)$ gives rise to is determined by $r$ and $g$ alone. ∎
7. Spikes
Spikes have played many important roles in matroid theory in recent decades (see [23]). The following generalization of spikes was introduced in [6]: for $r,s,t\in\mathbb{N}$ with $r\geq 3$, $s\geq r-1$, and $t\geq 2$, an $(r,s,t+1)$-spike with tip $a$ is a rank-$r$ simple matroid whose ground set is the union of $s$ lines $L_{1},L_{2},\ldots,L_{s}$, each having $t+1$ elements and containing $a$, for which, for any $k\in[r-1]$, the union of any $k$ of these lines has rank $k+1$. An $r$-spike is an $(r,r,3)$-spike. For example, the Fano plane $F_{7}$ is a $3$-spike and (atypically) any element can be the tip; for a given point $a\in F_{7}$, relaxing any one or more of the four circuit-hyperplanes that do not contain $a$ gives more examples of $3$-spikes. Also, there is only one $(r,r-1,t+1)$-spike up to isomorphism, namely, the parallel connection at $a$ of $r-1$ lines, each having $t+1$ elements.
In [6], $(r,s,t+1)$-spikes with $r\geq 5$ were shown to be distinguishable from all other matroids (but not necessarily from each other) by their Tutte polynomials, and it was shown which $(r,s,t+1)$-spikes have the same Tutte polynomial; those results imply that certain generalized spikes are Tutte unique. We identify some $(r,s,t+1)$-spikes that are extremal. (We focus exclusively on spikes with tips. Peter Nelson (private communications) has treated $r$-spikes without tips.)
Brylawski [11] showed that two paving matroids of the same rank on the same number of elements have the same Tutte polynomial if and only if the multisets of sizes of their dependent hyperplanes are equal; the same is true for the $\mathcal{G}$-invariant [7, Theorem 3.7]. For $(t+1)$-point lines $L_{1},L_{2},\ldots,L_{s}$, consider all matroids formed by taking a parallel connection of $L_{1}$ and $L_{2}$, and then taking a parallel connection of that with $L_{3}$, and so on, and truncating to rank $3$; all of these paving matroids have the same $\mathcal{G}$-invariant, but not all are $(3,s,t+1)$-spikes. Thus, $(3,s,t+1)$-spikes cannot be characterized using valuative invariants, so we focus on $r\geq 4$.
In any $(r,s,t+1)$-spike $M$ with $(t+1)$-point lines $L_{1},L_{2},\ldots,L_{s}$, any $3$-subset of any $L_{i}$ is a circuit, as is any $4$-set $C$ for which $|C\cap(L_{i}-\{a\})|=2=|C\cap(L_{j}-\{a\})|$ for some $\{i,j\}\subseteq[s]$. All other circuits in $M$ have size $r$ or $r+1$. An $r$-set $X$ with $a\not\in X$ that contains one element from each of $r$ lines among $L_{1},L_{2},\ldots,L_{s}$ is either a basis or a circuit. Thus, there can be two types of $4$-circuits if $r=4$. The hyperplane $H$ that is spanned by an $r$-circuit $X$ described above intersects each set $L_{i}-\{a\}$ in at most one element, and if $s=r$, then $H$ is a circuit-hyperplane. For any $r$, $s$, and $t$, there is a unique $(r,s,t+1)$-spike, up to isomorphism, that has no $r$-circuits, or, for $r=4$, just $\binom{s}{2}\bigl{(}\binom{t}{2}\bigr{)}^{2}$, from the $(2t+1)$-point planes; this is the free $(r,s,t+1)$-spike. The free $(r,s,t+1)$-spike is the truncation to rank $r$ of the parallel connection of $s$ lines at $a$, each having $t+1$ elements. The free $(r,s,t+1)$-spike with $r\geq 4$ is Tutte unique (see [6]).
Let $C$ be an $r$-circuit in $\operatorname{PG}(r-1,2)$, let $a$ be an element of $\operatorname{PG}(r-1,2)$ that is not in the hyperplane $\operatorname{cl}(C)$, and let $X$ be the union of the lines $L_{e}=\operatorname{cl}(\{a,e\})$ for $e\in C$. The restriction $M=\operatorname{PG}(r-1,2)|X$ is an $r$-spike, and $C$ is a circuit-hyperplane of $M$. It is easy to check that the symmetric difference of $C$ with $k$ of the sets $L_{e}-\{a\}$, for $e\in C$, is a circuit-hyperplane if and only if $k$ is even, otherwise the symmetric difference is a basis of $M$. Thus, $M$ has $2^{r-1}$ circuit-hyperplanes. It is easy to see that $2^{r-1}$ is the maximum number of circuit-hyperplanes in an $r$-spike. Using the ideas about symmetric differences above, it is not hard to construct an isomorphism between $M$ and any $r$-spike with $2^{r-1}$ circuit-hyperplanes. This justifies calling the $r$-spike with $2^{r-1}$ circuit-hyperplanes the binary $r$-spike. The binary $r$-spike with $r\geq 5$ is Tutte unique (see [6]).
In the next result we show how to identify $(r,s,t+1)$-spikes by optimizing valuative invariants, and we show that the free $(r,s,t+1)$-spike, the $(r,r-1,t+1)$-spike, and the binary $r$-spike are extremal. The assumption below about large lines being contained only in large planes amounts to, for all $w\neq 2t+1$, minimizing the valuative invariant $F_{2,3}(M;t+1,w)$ in part (7) of Theorem 2.5.
Theorem 7.1.
Fix integers $r\geq 4$, $t\geq 2$, and $s\geq r-1$. Let $M$ be a rank-$r$ matroid on $st+1$ elements. Assume that for $i<r$, each rank-$i$ flats has at most $(i-1)t+1$ elements. For $i<r$, call rank-$i$ flats with $(i-1)t+1$ elements large. Assume that each plane that contains a large line is large. Then $M$ has at most $s$ large lines, and if $M$ has $s$ large lines, then $M$ is an $(r,s,t+1)$-spike. Also, the free $(r,s,t+1)$-spike, the binary $r$-spike, and the $(r,r-1,t+1)$-spike are extremal.
Proof.
By the sizes of large lines and planes, coplanar large lines intersect in a point. Let $L$ be a large line of $M$. Any plane $P$ with $L\subsetneq P$ is large, so $|P-L|=t$. Now $|E(M)-L|=(s-1)t$ and $\pi=\{P-L\,:\,P\text{ is a plane, }L\subsetneq P\}$ is a partition of $E(M)-L$, so $\pi$ has $s-1$ blocks; say $\pi=\{B_{i}\,:\,i\in[s-1]\}$. The union of the planes $L\cup B_{i}$ and $L\cup B_{j}$, where $\{i,j\}\subseteq[s-1]$, has rank $4$ and size $13t+$ 1, so $L\cup B_{i}\cup B_{j}$ is a flat. Any large line $L^{\prime}\neq L$ of $M$ is $B_{i}\cup\{a\}$, for some $a\in L$ and $i\in[s-1]$, since otherwise $L^{\prime}$ would contain points in at least two blocks, say $B_{i}$ and $B_{j}$ where $\{i,j\}\subseteq[s-1]$; but then the line $L^{\prime}$ would be in the rank-$4$ flat $L\cup B_{i}\cup B_{j}$ and, since $t\geq 2$, it would have at least two points in one of $B_{i}$ or $B_{j}$, which would give the contradiction that $L^{\prime}\subseteq L\cup B_{i}$ or $L^{\prime}\subseteq L\cup B_{j}$. Thus, there are at most $s$ large lines.
Since $L$ was an arbitrary large line, the argument above shows that any two large lines of $M$ are coplanar. That implies that all large lines $L^{\prime}\neq L$ contain the same point of $L$ since two large lines $L^{\prime}$ and $L^{\prime\prime}$ intersect in some point, which is therefore in the planes $L\cup L^{\prime}$ and $L\cup L^{\prime\prime}$ and so in their intersection, $L$. It follows that if $M$ has $s$ large lines, then $M$ is an $(r,s,t+1)$-spike.
The free $(r,s,t+1)$-spike is extremal since, among $(r,s,t+1)$-spikes, it is the only one that minimizes the number of $r$-circuits. The binary $r$-spike is extremal since, among $r$-spikes, it is the only one that maximizes the number of circuit-hyperplanes. Finally, the $(r,r-1,t+1)$-spike is extremal since it is the only spike with its parameters. ∎
We give a second characterization of $(r,r-1,t+1)$-spikes so that in the next section we can show that direct sums of such spikes, where $t$ is fixed, are extremal. The notion of an $(r,r-1,t+1)$-spike makes sense even when $r=2$, in which case a $(2,1,t+1)$-spike is a $(t+1)$-point line. What we gain by this mild extension is a slightly stronger result about direct sums in the next section. While Theorem 7.1 is limited to ranks $4$ and higher, Theorem 7.2 applies for ranks $2$ and higher.
Theorem 7.2.
Fix integers $r,t\in\mathbb{N}-\{1\}$. Let $M$ be a rank-$r$ matroid on $(r-1)t+1$ elements. Assume that for each $i\in[r]$, each rank-$i$ flat has at most $(i-1)t+1$ elements. For $i\in[r]$, call rank-$i$ flats with $(i-1)t+1$ elements large, and call a flag of $M$ large if it is counted by $F_{0,r}(N;0,1,t+1,2t+1,\ldots,(r-1)t+1)$. Then $M$ has at most $(r-1)!(t+1)$ large flags, and $M$ has $(r-1)!(t+1)$ large flags if and only if $M$ is an $(r,r-1,t+1)$-spike. Also, if $r>2$, then $M$ has $(r-1)!(t+1)$ large flags if and only if $M$ has $r-1$ large hyperplanes, and that is the maximum number of large hyperplanes under the assumptions on $M$.
Proof.
The statement is easy to see when $r=2$, so we focus on $r\geq 3$. We first treat the bound on the number of large hyperplanes. A rank-$3$ matroid on $12t+$ 1 points has at most two $(t+1)$-point lines, so the bound holds for $r=3$. Now assume that $r>3$ and that the bound holds for matroids of smaller rank. Let $H_{1},H_{2},\ldots,H_{s}$ be the large hyperplanes of $M$. A simpler counterpart of the inclusion/exclusion arguments used in the last two sections shows that if $\{i,j\}\subseteq[s]$, then $|H_{i}\cap H_{j}|\geq(r-3)t+1$. Since $H_{i}\cap H_{j}$ is a flat of rank at most $r-2$, the inequality and the assumed bound on the sizes of flats of each rank imply that $r(H_{i}\cap H_{j})=r-2$ and $|H_{i}\cap H_{j}|=(r-3)t+1$, and from that we get $H_{i}\cup H_{j}=E(M)$. The equality $H_{i}\cup H_{j}=E(M)$ shows that all intersections $H_{i}\cap H_{j}$, for $\{i,j\}\subseteq[s]$, are distinct since $H_{i}$ and $H_{j}$ are the only hyperplanes that contain $H_{i}\cap H_{j}$. By the induction hypothesis applied to $M|H_{s}$, we get $s-1\leq r-2$, so $s\leq r-1$, as needed.
By iterating the argument above, we get that the intersection of any $i$ large hyperplanes is a large rank-$(r-i)$ flat and all such intersections are different. In particular, if $M$ has $r-1$ large hyperplanes, then $M$ has $(r-1)!(t+1)$ large flags and $\binom{r-1}{r-2}=r-1$ large lines; also, since two $(t+1)$-point lines in a $(2t+1)$-point plane intersect in a point, it follows that all large hyperplanes, and so all large lines, intersect in one point, so $M$ is an $(r,r-1,t+1)$-spike.
Arguments like those given in earlier sections show that, when $r>2$, having $r-1$ large hyperplanes is equivalent to having $(r-1)!(t+1)$ large flags. ∎
8. Direct sums of some extremal matroids
Ferroni and Fink [19] conjectured that direct sums of extremal matroids are extremal. While it is far from settling that conjecture, in Theorem 8.2 below, we show that direct sums of extremal matroids that have certain common features that are captured by valuative invariants are extremal. Examples of extremal matroids with the relevant common features that we treat are (a) cycle matroids of complete graphs, (b) projective geometries of rank at least four and the same order, (c) affine geometries of rank at least four and the same order, (d) Dowling geometries based on a group that is the only group of its order, and (e) $(r,r-1,t+1)$-spikes where $t$ is fixed.
We start with another special case of the conjecture; we treat direct sums of uniform matroids. Ferroni and Fink [19] observed that uniform matroids are extremal since they maximize the number of bases. We use a result that was noted after Theorem 2.5: the function that gives the number of cyclic flats of a matroid is a valuative invariant.
Theorem 8.1.
Direct sums of uniform matroids are extremal.
Proof.
Let $M$ be $U_{r_{1},n_{1}}\oplus U_{r_{2},n_{2}}\oplus\cdots\oplus U_{r_{k},n_{k}}$ where $r_{1}\leq r_{2}\leq\cdots\leq r_{k}$ and, if $r_{i}=r_{i+1}$, then $n_{i}\geq n_{i+1}$. Let $r=r_{1}+r_{2}+\cdots+r_{k}$ and $n=n_{1}+n_{2}+\cdots+n_{k}$. If $M$ has $s$ loops and $t$ coloops, then we can focus on matroids that have $s$ loops and $t$ coloops by maximizing the valuative invariant $N\mapsto f_{0}(N;s,0)+f_{r}(N;n,t)$ within the class of rank-$r$ matroids on $n$ elements (the maximum is $2$). Thus, without loss of generality, we can assume that $M$ has neither loops nor coloops, so $0<r_{i}<n_{i}$ for each $i\in[k]$. Consider all matroids of rank $r$ on $n$ elements with neither loops nor coloops. The sum of the coefficients of the terms $x^{i}y^{j}$, for $i+j<k$, in the Tutte polynomial of a matroid is a valuative invariant, and by minimizing it, we can focus on matroids that have at least $k$ connected components. The number of cyclic flats in a matroid is a valuative invariant, and the matroids that minimize it, within the class of matroids that have at least $k$ connected components, have exactly $k$ connected components and their $2^{k}$ cyclic flats are precisely the unions of those components, so each component has no cyclic flats other than $\emptyset$ and its ground set; thus, each component is uniform. Now we must distinguish between $M$ and all other direct sums $N=U_{r^{\prime}_{1},n^{\prime}_{1}}\oplus U_{r^{\prime}_{2},n^{\prime}_{2}}\oplus\cdots\oplus U_{r^{\prime}_{k},n^{\prime}_{k}}$ where $r^{\prime}_{1}\leq r^{\prime}_{2}\leq\cdots\leq r^{\prime}_{k}$ and, if $r^{\prime}_{i}=r^{\prime}_{i+1}$, then $n^{\prime}_{i}\geq n^{\prime}_{i+1}$, and $r=r^{\prime}_{1}+r^{\prime}_{2}+\cdots+r^{\prime}_{k}$ and $n=n^{\prime}_{1}+n^{\prime}_{2}+\cdots+n^{\prime}_{k}$. The lexicographically first rank sequence that has a nonzero coefficient in $\mathcal{G}(M)$ is
[…]
Let $\underline{s}(M)=x_{1}\ldots x_{n}$ be this rank sequence. Let the lexicographically first rank sequence that has a nonzero coefficient in $\mathcal{G}(N)$ be $\underline{s}(N)=y_{1}\ldots y_{n}$. Since $\underline{s}(M)$ encodes the rank and size of each connected component of $M$, and likewise for $\underline{s}(N)$, if $M$ and $N$ are not isomorphic, then there is a least $i$ for which $x_{i}\neq y_{i}$. First assume that $x_{i}>y_{i}$, so $x_{i}=1$ and $y_{i}=0$. Let $p=x_{1}+\cdots+x_{i-1}=y_{1}+\cdots+y_{i-1}$. Since $\underline{s}(M)$ and $\underline{s}(N)$ agree in the first $i-1$ places, both $M$ and $N$ have sets of size $i-1$ and rank $p$ that can be ordered in such a way that, for each $j\in[i-1]$, the rank of the set of the first $j$ elements is $x_{1}+\cdots+x_{j}=y_{1}+\cdots+y_{j}$; however, since $x_{i}=1$ and $y_{i}=0$, at least one such $(i-1)$-set in $N$ can be extended to a set of size $i$ and rank $p$, while no such $(i-1)$-set in $M$ has an extension to a set of size $i$ and rank $p$. Thus, the coefficient $g_{\underline{s}(N)}(M)$ of $[\underline{s}(N)]$ in $\mathcal{G}(M)$ is zero. Eliminate such matroids $N$ from consideration by, for each, minimizing the valuative invariant $M^{\prime}\mapsto g_{\underline{s}(N)}(M^{\prime})$. Now assume that $x_{i}<y_{i}$. Switching the roles of $M$ and $N$ in the argument above shows that $g_{\underline{s}(M)}(N)=0$. Eliminate such matroids $N$ by maximizing $M^{\prime}\mapsto g_{\underline{s}(M)}(M^{\prime})$. The only matroid left is $M$, so $M$ is extremal. ∎
We next give a result on direct sums of extremal matroids that, while far from proving the conjecture of Ferroni and Fink, applies to direct sums of many of the matroids that we treated earlier in this paper. Looking ahead to Corollary 8.3 may make the notation below easier to digest. Characterizations of the type specified in this result are illustrated by Theorem 5.3 for $M(K_{r+1})$. In many applications, $\mathcal{M}$ contains, up to isomorphism, one matroid of each rank (e.g,. $\mathcal{M}$ can be the set of cycle matroids of complete graphs of rank two and greater). For the general case, an example to keep in mind is Dowling geometries, where many matroids maximize the number of large flags, but only Dowling geometries also maximize the number of $3$-point lines.
Theorem 8.2.
Let $\alpha_{0}=0,\alpha_{1}=1,\alpha_{2},\alpha_{3},\ldots$ be integers for which
  • (1)
    $\alpha_{2}\geq 3$ and $\alpha_{i}+\alpha_{j}\leq\alpha_{i+j-1}+1$ for all $i,j\in\mathbb{N}$.
Fix a set $A\subseteq[\alpha_{2}]-\{1\}$ with $2,\alpha_{2}\in A$. Let $\mathcal{M}$ be a set of matroids of rank at least $2$ that satisfies the following conditions:
  • (2)
    for all $M\in\mathcal{M}$ and $i\leq r(M)$, some rank-$i$ flat of $M$ has size $\alpha_{i}$,
  • (3)
    for each $M\in\mathcal{M}$ and flat $F$ of $M$, if $2<r(F)<r(M)$, then either $|F|=\alpha_{r(F)}$ or $|F|\leq\alpha_{r(F)-1}+1$, while if $r(F)=2$, then $|F|\in A$, and
  • (4)
    for $r\geq 2$, a rank-$r$ matroid $M$ is in $\mathcal{M}$ if and only if, among rank-$r$ matroids that satisfy items (2)-(3), the valuative invariant $N\mapsto F_{0,r}(N;\alpha_{0},\alpha_{1},\alpha_{2},\ldots,\alpha_{r})$ is maximized by $M$.
If, for some sequence of valuative invariants $f_{1},f_{2},\ldots,f_{k}$ that are additive on direct sums (i.e., $f_{i}(M\oplus N)=f_{i}(M)+f_{i}(N)$), the matroids $M_{1},M_{2},\ldots,M_{t}$ are, up to isomorphism, the only matroids in $\mathcal{M}$ of their ranks that maximize $f_{1}$, and then $f_{2}$, and so on, through $f_{k}$, then $M_{1}\oplus M_{2}\oplus\cdots\oplus M_{t}$ is extremal.
Proof.
Call rank-$i$ flats of size $\alpha_{i}$ in a matroid $M\in\mathcal{M}$ large, and call a flag of $M$ large if it has flats of each rank from 00 to $r(M)$, all of which are large. Thus, if $M\in\mathcal{M}$, then $F_{0,r}(M;\alpha_{0},\alpha_{1},\alpha_{2},\ldots,\alpha_{r})$ is the number of large flags of $M$. It follows from conditions (1)-(2) that each matroid in $\mathcal{M}$ is connected.
Fix any matroids $M_{1},M_{2},\ldots,M_{t}$ in $\mathcal{M}$ on disjoint ground sets. We can assume that $r(M_{i})\geq r(M_{i+1})$ for $i\in[t-1]$. Set $M=M_{1}\oplus M_{2}\oplus\cdots\oplus M_{t}$. We start with some observations about $M$.
By condition (1), for $i\geq 2$, no rank-$i$ flat in $M$ that has elements in two or more of $M_{1},M_{2},\ldots,M_{t}$ has size more than $\alpha_{i-1}+1$. Thus, $M$ satisfies the bounds in condition (3) and any rank-$i$ flat of $M$ of size $\alpha_{i}$ is a flat of some $M_{j}$. Therefore, we can focus on direct sums that satisfy the bounds in condition (3).
For $s\in[t+1]$, let $p_{s}$ be the sum of the first $s-1$ terms among $r(M_{1}),r(M_{2}),\ldots,r(M_{t})$, so $p_{1}=0$ and $p_{t+1}=r(M)$. If $i\in[r(M)-1]$, then $i=p_{s}+u$ for unique $s\in[t]$ and $u\in[r(M_{s})-1]\cup\{0\}$, and the largest rank-$i$ flats in $M$ have size
$\beta_{i}=\alpha_{r(M_{1})}+\alpha_{r(M_{2})}+\cdots+\alpha_{r(M_{s-1})}+\alpha_{u}.$
Set $\beta_{r(M)}=|E(M)|$. We extend the terminology of large flats and large flags to $M$ in the obvious way, using the sequence $\beta_{0},\beta_{1},\ldots,\beta_{r(M)}$. Note that $\beta_{i}\leq\alpha_{i}$ for all $i$, with equality if and only if $i\leq r(M_{1})$. To get the large flags of $M$, start with a large flag of a matroid among $M_{1},M_{2},\ldots,M_{t}$ of largest rank, then take the union of the ground set of that matroid with, in turn, the nonempty sets in a large flag of a matroid of largest rank among $M_{1},M_{2},\ldots,M_{t}$ that has not yet been selected, and repeat until we reach $E(M)$. Condition (1) gives $\beta_{k+1}-\beta_{k}=1$ for $t$ integers $k\in[r(M)]$, namely, $k=p_{s}$ with $s\in[t]$. Since $r(M_{i})\geq r(M_{i+1})$ for $i\in[t-1]$, if $F_{0,r(M)}(M;\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)})\neq 0$ for a sequence $\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)}$, then for each $s\in[t]$, we have $\gamma_{k}-\gamma_{k-1}=1$ for at least $s$ integers $k\in[p_{s}+1]$. Thus, if, for some $s\in[t]$, a sequence $\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)}$ has $\gamma_{i}=\beta_{i}$ for $0\leq i\leq p_{s}$ and $\gamma_{p_{s}+1}-\gamma_{p_{s}}>1$, then $F_{0,r(M)}(M;\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)})=0$.
Consider matroids of rank $r(M)$ and size $|E(M)|=\beta_{r(M)}$. We claim that optimizing the valuative invariants below, in the order listed, isolates the direct sums $N_{1}\oplus N_{2}\oplus\cdots\oplus N_{t}$ where $N_{i}\in\mathcal{M}$ and, if the components are indexed so that $r(N_{i})\geq r(N_{i+1})$ for $i\in[t-1]$, then $r(M_{i})=r(N_{i})$, for all $i\in[t]$.
  • (a)
    To focus on simple matroids, maximize $N\mapsto f_{1}(N;1)$.
  • (b)
    As justified above, minimize $N\mapsto f_{2}(N;j)$ for each $j\not\in A$; also, for all $i,j$ with $3\leq i\leq r(M)$ and $\alpha_{i-1}+1<j<\alpha_{i}$, minimize $N\mapsto f_{i}(N;j)$.
  • (c)
    To focus on matroids with at least $t$ connected components, minimize the sum of the coefficients of the terms $x^{i}y^{j}$ with $i+j<t$ in the Tutte polynomial $T(N;x,y)$.
  • (d)
    For each sequence $\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)}$ where, for some $s\in[t]$, we have $\gamma_{i}=\beta_{i}$ for $0\leq i\leq p_{s}$ but $\gamma_{p_{s}+1}-\gamma_{p_{s}}>1$, minimize $N\mapsto F_{0,r(M)}(N;\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)})$.
  • (e)
    Maximize $N\mapsto F_{0,r(M)}(N;\beta_{0},\beta_{1},\ldots,\beta_{r(M)})$.
Since $\beta_{k+1}-\beta_{k}=1$ if and only if $k=p_{s}$ with $s\in[t]$, matroids that satisfy (c) and (e) have $t$ connected components. Let $N=N_{1}\oplus N_{2}\oplus\cdots\oplus N_{t}$ satisfy conditions (a)-(e), so each $N_{i}$ is connected. Let $F_{0}\subsetneq F_{1}\subsetneq\cdots\subsetneq F_{r(M)}$ be a flag in $N$ that $F_{0,r(M)}(N;\beta_{0},\ldots,\beta_{r(M)})$ counts. Since $\beta_{i+1}-\beta_{i}>1$ for $i\in[r(M_{1})-1]$, we have $F_{r(M_{1})}\subseteq E(N_{j})$ for some $j\in[t]$; we may assume that $j=1$. If $E(N_{1})\neq F_{r(M_{1})}$, then $|X-F_{r(M_{1})}|>1$ for some flat $X$ of $N_{1}$ that covers $F_{r(M_{1})}$ (otherwise more than $t$ sets $F_{i+1}-F_{i}$ would be singletons, so more than $t$ terms $\beta_{i+1}-\beta_{i}$ would be $1$); thus, $F_{0,r(M)}(N;\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)})>0$ for some sequence $\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)}$ for which $\gamma_{i}=\beta_{i}$ for $0\leq i\leq r(M_{1})$ and $\gamma_{r(M)_{1}+1}-\gamma_{r(M_{1})}=|X-F_{r(M_{1})}|>1$, but the minimization in item (d) prevents that. Thus, $E(N_{1})=F_{r(M_{1})}$. The same ideas show that, up to reindexing the matroids $N_{j}$, we have $r(N_{i})=r(M_{i})$ for all $i\in[t]$, and $E(N_{i})=F_{p_{i+1}}-F_{p_{i}}$ when $i\in[t]$. For $F_{0,r(M)}(N;\beta_{0},\beta_{1},\ldots,\beta_{r(M)})$ to be maximal, each $F_{0,r(N_{i})}(N_{i};\alpha_{0},\alpha_{1},\ldots,\alpha_{r(N_{i})})$, for $i\in[t]$, must be maximal, and so each $N_{i}$ is in $\mathcal{M}$, as needed.
We now use the hypothesis that there are valuative invariants $f_{1},f_{2},\ldots,f_{k}$ for which, among all matroids in $\mathcal{M}$, the matroids $M_{1},M_{2},\ldots,M_{t}$ are, up to isomorphism, the unique matroids of their ranks that maximize $f_{1}$, and then $f_{2}$, and so on, through $f_{k}$. By the assumption that each $f_{i}$ is additive on direct sums, $f_{i}(N)$ is maximized precisely when each $f_{i}(N_{j})$ is maximized. By assumption, that is when $N_{j}$ is isomorphic to $M_{j}$. Thus, maximizing $f_{1}$, and then $f_{2}$, and so on through $f_{k}$ leaves one matroid, up to isomorphism, namely, $M$, as claimed. ∎
The assumption that $f_{1},f_{2},\ldots,f_{k}$ are additive can be replaced by any condition that ensures that each $f_{i}$ is maximized on a matroid precisely when it is maximized on the restrictions to its connected components.
Corollary 8.3.
Direct sums of cycle matroids of complete graphs of rank at least two are extremal. For a fixed prime power $q$, direct sums of projective geometries of order $q$ and rank at least four are extremal, and the same holds for affine geometries. If there is only one projective or affine plane of order $q$, up to isomorphism, then planes can be included in the previous statement. For a group $G$ for which $G$ is the only group of order $|G|$ (e.g., prime order), direct sums of Dowling geometries of rank at least two based on $G$ are extremal. For a fixed integer $t\geq 2$, direct sums of $(r,r-1,t)$-spikes of rank at least two are extremal.
Proof.
For all cases except Dowling geometries, the set $\mathcal{M}$ in Theorem 8.2 has, up to isomorphism, a unique matroid of each rank. The first statement follows directly from Theorems 5.3 and 8.2 with $\alpha_{i}=\binom{i+1}{2}$ and $A=\{2,3\}$. The assertion about projective geometries follows Theorems 3.2 and 8.2 with $\alpha_{i}=\frac{q^{i}-1}{q-1}$, for $i\geq 1$, and $A=\{2,q+1\}$; likewise for affine geometries, with $\alpha_{i}=q^{i-1}$, for $i\geq 1$, and $A=\{2,q\}$. For Dowling geometries, the result follows from Theorems 6.3, 6.4, and 8.2 with $\alpha_{i}=g\binom{i}{2}+i$ and $A=\{2,3,|G|+2\}$ by using a single additional valuative invariant $f_{1}$, namely, the number of $3$-point lines. For $(r,r-1,t)$-spikes, the result follows from Theorems 7.2 and 8.2 with $\alpha_{i}=1+(i-1)t$ for $i\geq 1$ and $A=\{2,t+1\}$. ∎
Acknowledgments
The author is grateful to Luis Ferroni, Alex Fink, and Peter Nelson for helpful and encouraging e-mail exchanges while this material was being developed.