### v2

# Characterizations of certain matroids by maximizing valuative invariants

Abstract.
Luis Ferroni and Alex Fink recently introduced a polytope of all unlabeled matroids of rank $r$ on $n$ elements, and they showed that the vertices of this polytope come from matroids that can be characterized by maximizing a sequence of valuative invariants. We prove that a number of the matroids that they conjectured to yield vertices indeed do (these include cycle matroids of complete graphs, projective geometries, and Dowling geometries), and we give additional examples (including truncations of cycle matroids of complete graphs, Bose-Burton geometries, and binary and free spikes with tips). We prove a special case of a conjecture of Ferroni and Fink by showing that direct sums of uniform matroids yield vertices of their polytope, and we prove a similar result for direct sums whose components are in certain restricted classes of matroids.
Key words and phrases:
Matroid, extremal matroid theory, valuative invariant, polytope of matroids
1991 Mathematics Subject Classification:
Primary: 05B35
1. Introduction
By a result of Derksen and Fink [16], when the base polytope of a matroid is written as an indicator function, the base polytope of any matroid of rank $r$ on $n$ elements is a unique linear combination of base polytopes of nested matroids of rank $r$ on $n$ elements and the coefficients are integers. Matroid invariants assign the same value to isomorphic matroids. Valuative invariants are matroid invariants $f$ that take values in an abelian group and for which, for any matroid $M$, the image $f(M)$ can be found by taking the integer linear combination of the base polytopes of nested matroids that give the base polytope of $M$ and replacing each nested matroid by its image under $f$. (See Section 2.3 for more precise statements.) Valuative invariants include many well-known invariants, such as the Tutte polynomial, but they encompass far more than deletion-contraction invariants. Ferroni and Fink [20] introduced two polytopes of matroids; one is the convex hull of the coordinate vectors of the base polytopes of matroids of rank $r$ and size $n$ relative to the base polytopes of nested matroids of rank $r$ and size $n$; the other, $\Omega_{r,n}$, is a counterpart for unlabeled matroids (that is, isomorphism types of matroids). They showed that the vertices of $\Omega_{r,n}$ come from matroids, or sets of matroids that give the same point in $\Omega_{r,n}$, that can be characterized by sequences of maximization (or, equivalently, minimization) problems of valuative invariants. We illustrate such characterizations, which are the focus of this paper, with Theorem 5.2 below, which shows that, as Ferroni and Fink conjectured, the cycle matroid $M(K_{r+1})$ of the complete graph $K_{r+1}$ yields a vertex.
Assume that $M$ is a simple matroid of rank $r$ on $\binom{r+1}{2}$ elements, that all lines have $2$ or $3$ elements, that all planes have $3$, 44, or $6$ elements, and that, for all $h$ with $4\leq h<r$, all rank-$h$ flats of $M$ have at most $\binom{h+1}{2}$ elements. Then the number of $3$-point lines in $M$ is at most $\binom{r+1}{3}$; also, if the number of $3$-point lines in $M$ is $\binom{r+1}{3}$, then $M$ is $M(K_{r+1})$.
To see how this conforms to the description of the vertices of $\Omega_{r,\binom{r+1}{2}}$, we use the fact that the number of flats of a given rank and size in a matroid is a valuative invariant, so maximizing the number of flats of rank $1$ and size $1$ limits the scope to simple matroids. Then we can minimize the number of lines of sizes other than two or three to say that all lines in the matroids of interest have two or three points, and likewise we can then impose the remaining conditions on the sizes of the flats of each rank. Relative to those conditions, we show that the greatest possible number of $3$-point lines is $\binom{r+1}{3}$, and, for the final link, we show that the maximum is achieved precisely by $M(K_{r+1})$. Thus, a sequence of optimization problems for valuative invariants hones in on $M(K_{r+1})$, so $M(K_{r+1})$ gives rise to a vertex of $\Omega_{r,\binom{r+1}{2}}$.
As that example suggests, these problems have the flavor of many that have long been studied, but with a new twist, namely, we are limited to optimizing valuative invariants. Thus, above we can not assume at the outset that the number of $3$-point lines is $\binom{r+1}{3}$; we can do that only after showing that that value is the maximum among matroids that satisfy the prior constraints. Identifying vertices of Ferroni and Fink’s polytopes $\Omega_{r,n}$ thus opens up a new and interesting chapter in extremal matroid theory. Appropriately, they call the matroids that yield vertices in their polytope extremal matroids.
After reviewing the necessary background in Section 2, we show that several matroids that Ferroni and Fink conjectured to be extremal indeed are: $M(K_{r+1})$ (Theorems 5.2 and 5.3) and Dowling geometries $Q_{r}(G)$ based on groups (or quasigroups, if $r=3$; Theorems 6.2, 6.3, and 6.4) are extremal. They conjectured that projective geometries are extremal; in Section 3 we prove that, more broadly, all perfect matroid designs (i.e., matroids in which flats of the same rank have the same size) are extremal; this applies to projective geometries, affine geometries, their truncations, and much more. Among matroids that Ferroni and Fink did not mention, we show that truncations of $M(K_{r+1})$ to rank at least five (Theorem 5.4), Bose-Burton geometries (Theorems 4.3 and 4.5), and binary and free spikes with tips (Theorem 7.1) are extremal. Ferroni and Fink conjecture that direct sums of extremal matroids are extremal; we do not have a proof of that conjecture, but provide evidence for it by showing that direct sums of uniform matroids are extremal (Theorem 8.1) and giving a general result (Theorem 8.2) that shows that direct sums of cycle matroids of complete graphs are extremal, and likewise for projective geometries of the same order, affine geometries of the same order, Dowling geometries based on a group that is the only group of its order, and certain spikes (Corollary 8.3).
2. Background
For basic matroid theory, we recommend [24]; we follow the notation used there. All matroids considered in this paper are finite. We use $\mathbb{N}$ for the set of positive integers, $\mathbb{N}_{0}$ for the set of nonnegative integers, and $[n]$ for the set $\{1,2,\ldots,n\}$.
2.1. A few items about matroids
A set $X$ in a matroid $M$ is cyclic if it is a union of circuits, or, equivalently, $M|X$ has no coloops. We are most interested in cyclic flats, that is, flats that are cyclic. The set of cyclic flats of $M$, which is denoted $\mathcal{Z}(M)$, is a lattice under inclusion. The least element of $\mathcal{Z}(M)$ is $\operatorname{cl}_{M}(\emptyset)$, the set of loops of $M$, and the greatest element is $E(M)-\operatorname{cl}_{M^{*}}(\emptyset)$ since $\operatorname{cl}_{M^{*}}(\emptyset)$ is the set of coloops of $M$. Routine arguments give the equality $r(X)=\min\{r(F)+|X-F|\,:\,F\in\mathcal{Z}(M)\}$ for all $X\subseteq E(M)$, so knowing $E(M)$, the cyclic flats of $M$, and the rank of each cyclic flat determines $M$.
We use the term cover in the lattice of flats of a matroid $M$ and in its lattice of cyclic flats $\mathcal{Z}(M)$. In general, $y$ is a cover of $x$, or $y$ covers $x$, in an ordered set if $x<y$ and there is no $z$ with $x<z<y$. If $Y$ covers $X$ in the lattice of flats, then $r(Y)=r(X)+1$, but if $Y$ covers $X$ in $\mathcal{Z}(M)$, then r(Y)r(X)r(Y)-r(X) can be any integer in [r(M)r(X)][r(M)-r(X)].
[…]
Matroids are labeled: the ground set matters, and, for instance, which sets are bases matters. However, sometimes all we care about is the isomorphism type of the matroid. In that case, we call the isomorphism class of a matroid, or a preferred representative of that class, an unlabeled matroid. The next item illustrates this with a relevant class of matroids.
A matroid $M$ is nested if $\mathcal{Z}(M)$ is a chain, that is, for any cyclic flats $F$ and $G$ of $M$, either $F\subseteq G$ or $G\subseteq F$. Nested matroids are known by many names, including Schubert matroids. Often all we care about is the isomorphism type of a nested matroid. Natural representatives of isomorphism classes of nested matroids are matroids on $[n]$ in which the nonempty cyclic flats are initial intervals $[s]$ of $[n]$. The number of unlabeled nested matroids of rank $r$ on $n$ elements is $\binom{n}{r}$, as we now justify. Consider sequences of matroids $M_{0},M_{1},\ldots,M_{n}$, where $M_{0}$ is empty, and for each $i\in[n]$, we get $M_{i}$ by adding the element $i$ to $M_{i-1}$ either (a) as a coloop or (b) via free extension; then $M_{n}$ is a nested matroid on $[n]$, and if option (a) is applied $r$ times, then $M_{n}$ has rank $r$. Since $E(M)\cup\{e\}$ is the unique cyclic flat of $M+e$ that contains $e$, consistent with what we said about a natural representative, the set of nonempty cyclic flats of $M_{n}$ is
[…]
The number of elements in $[s]$ that are added as coloops is the rank of $[s]$. There are $\binom{n}{r}$ options for which elements are added as coloops, so there are $\binom{n}{r}$ unlabeled rank-$r$ nested matroids on $n$ elements. We can encode $M_{n}$ as a lattice path $P=s_{1}s_{2}\ldots s_{n}$ where $s_{i}=N$ (a north step) if $i$ is a coloop of $M_{i}$, otherwise $s_{i}=E$ (an east step); if $r(M_{n})=r$, then the lattice path $P$ goes from $(0,0)$ to (nr,r)(n-r,r). For instance, the lattice path $NENEENNE$ encodes the unlabeled rank-$4$ nested matroid on eight elements in which the cyclic flats have the following size-rank pairs $(|X|,r(X))$: $(0,0)$, $(2,1)$, $(5,2)$, and $(8,4)$; the corresponding sets in the natural representative are $\emptyset$, $[2]$, $[5]$, and $[8]$.
[…]
We will review the matroids that may be less familiar (Dowling geometries, spikes, and Bose-Burton geometries) at the beginnings of the sections in which they are treated.
2.2. The Tutte polynomial
The Tutte polynomial is the most well-known enumerative invariant in matroid theory (see, e.g., [13, 19]). For a matroid $M$, its Tutte polynomial is defined to be
$T(M;x,y)=\sum_{A\subseteq E(M)}(x-1)^{r(M)-r(A)}(y-1)^{|A|-r(A)},$
which is a generating function for the $2^{|E(M)|}$ pairs $(|A|,r(A))$ for $A\subseteq E(M)$. Among the most basic properties of the Tutte polynomial are
  • (i)
[…]
  • (ii)
    the direct sum factorization: $T(M\oplus N;x,y)=T(M;x,y)T(N;x,y)$,
[…]
Lemma 2.1.
A matroid $M$ is connected if and only if the coefficient of at least one of $x$ or $y$ in its Tutte polynomial $T(M;x,y)$ is nonzero.
Corollary 2.2.
A matroid $M$ has at least $k$ connected components if and only if all terms $x^{i}y^{j}$ in $T(M;x,y)$ with $i+j<k$ have zero coefficient.
The Tutte polynomial is the universal deletion/contraction invariant, as the following result by Oxley and Welsh [25] (extending a result of Brylawski [12]) makes precise.
Theorem 2.3.
Let $R$ be a commutative ring with unity. For any $u,v,\sigma,\tau\in R$ with $\sigma$ and $\tau$ nonzero, there is a unique function $t$ from the set of all matroids into $R$ that has the following properties.
  1. (1)
    If $E(M)=\emptyset$, then $t(M)=1$.
  2. (2)
    If $e$ is a loop of $M$, then $t(M)=v\cdot t(M\backslash e)$.
  3. (3)
    If $e$ is a coloop of $M$, then $t(M)=u\cdot t(M/e)$.
  4. (4)
    If $e$ is neither a loop nor a coloop, then $t(M)=\sigma\cdot t(M\backslash e)+\tau\cdot t(M/e)$.
Furthermore, $t$ is the following evaluation of the Tutte polynomial $T(M;x,y)$:
$t(M)=\sigma^{|E(M)|-r(M)}\tau^{r(M)}T(M;u/\tau,v/\sigma).$
A matroid $M$ is Tutte unique if the equality $T(M;x,y)=T(N;x,y)$ holds only when $N$ is isomorphic to $M$. Many earlier characterizations of matroids by numerical invariants that we will use, such as the characterization of $M(K_{r+1})$ in Theorem 5.1, were proven to show that these matroids are Tutte unique, but in most cases the characterizations do not fit the pattern of maximizing a sequence of valuative invariants that is required in order to show that the matroids yield vertices of the polytope of unlabeled matroids. Some Tutte-unique matroids do not yield vertices of the polytope of unlabeled matroids (e.g., the unlabeled nested matroid that is encoded by the lattice path $NENE$; see [20, Examples 3.3 and 4.5]), and some matroids that yield vertices are not Tutte unique (e.g., Dowling geometries of the same rank based on nonisomorphic groups of the same order). However, there are strong connections between these two notions. For more about Tutte-unique matroids, see [8].
2.3. Matroid base polytopes, valuative invariants, and the $\mathcal{G}$-invariant
We discuss matroid base polytopes very lightly; readers wanting more on this topic are referred to [2, 16, 22]. To be able to use simpler notation, when discussing matroid base polytopes it is common to take the ground set of a matroid to be $[n]$. Given a matroid $M$ on $[n]$, a basis $B$ of $M$ can be encoded by its characteristic vector $\mathbf{v}_{B}$ in $\mathbb{R}^{n}$, that is, the vector in which entry $i$ is $1$ if and only if $i\in B$, otherwise entry $i$ is 0. The convex hull in $\mathbb{R}^{n}$ of the vectors $\mathbf{v}_{B}$, as $B$ ranges over all bases of $M$, is the matroid base polytope $\mathcal{P}(M)$.
For a subset $X$ of $\mathbb{R}^{n}$, its indicator function $\bar{\mathbb{I}}(X):\mathbb{R}^{n}\to\mathbb{R}$ is given by
$\bar{\mathbb{I}}(X)(x)=\begin{cases}1,&\text{ if }x\in X,\\ 0,&\text{ if }x\not\in X.\end{cases}$
We focus on indicator functions of base polytopes of matroids. We call $\bar{\mathbb{I}}(\mathcal{P}(M))$ the indicator function of $M$. The next result is due to Derksen and Fink [16, Theorem 5.4].
Theorem 2.4.
Let $M$ be a rank-$r$ matroid on $[n]$ and let $N_{1},N_{2},\ldots,N_{t}$ be the nested matroids of rank $r$ on $[n]$. There are unique integers $a_{1},a_{2},\ldots,a_{t}$ for which
(2.1) $\bar{\mathbb{I}}(\mathcal{P}(M))=a_{1}\cdot\bar{\mathbb{I}}(\mathcal{P}(N_{1}))+a_{2}\cdot\bar{\mathbb{I}}(\mathcal{P}(N_{2}))+\cdots+a_{t}\cdot\bar{\mathbb{I}}(\mathcal{P}(N_{t})).$
[…]
$^{\prime}aa$$cc^{\prime}$$^{\prime}bb$$dd^{\prime}$$M$
[…]
$^{\prime}aa$$cc^{\prime}$$dd^{\prime}$$^{\prime}bb$$M^{\prime}$
[…]
The form of Equation (2.1) motivates one of several equivalent formulations of the notion of a valuation. First, by a matroid invariant, we mean a function that is defined on the set of matroids and that assigns the same value to isomorphic matroids. A matroid invariant $f$ taking values in an abelian group $A$ is valuative, or is a valuation, if, for any matroids $M_{1},M_{2},\ldots,M_{t}$ of rank $r$ on $[n]$ and any integers $a_{1},a_{2},\ldots,a_{t}$, if
$a_{1}\cdot\bar{\mathbb{I}}(\mathcal{P}(M_{1}))(x)+a_{2}\cdot\bar{\mathbb{I}}(\mathcal{P}(M_{2}))(x)+\cdots+a_{t}\cdot\bar{\mathbb{I}}(\mathcal{P}(M_{t}))(x)=0$
for all $x\in\mathbb{R}^{n}$, then $a_{1}\cdot f(M_{1})+a_{2}\cdot f(M_{2})+\cdots+a_{t}\cdot f(M_{t})$ is the identity of $A$. Thus, by Equation (2.1), knowing the values of a valuative invariant $f$ on (unlabeled) nested matroids completely determines $f$.
For example, the indicator functions of the matroids in Figures 1 and 2, which must be different, are written in terms of different nested matroids, and for the matroid $M^{\prime}$ in Figure 2, we see a Möbius value, 2-2, that does not occur for the matroid $M$ in Figure 1. However, any valuative invariant $f$ gives the same values on these two matroids, namely,
$4\cdot f(NENENEEE)-2\cdot f(NENNEEEE)-f(NNEENEEE),$
where $f(P)$ is the value of $f$ at the unlabeled nested matroid that is encoded by the lattice path $P$; the first term is from the four coatoms in $\mathcal{C}(\mathcal{Z}(M))$ (or $\mathcal{C}(\mathcal{Z}(M^{\prime}))$) and the other terms are from the atoms, some of which, besides $\emptyset$ and $E(M)$, include a cyclic flat of rank $1$ and size $2$ (giving $NENNEEEE$) and the others include a cyclic flat of rank $2$ and size $4$ (giving $NNEENEEE$).
Derksen and Fink [16] proved a counterpart of Theorem 2.3 for valuative invariants; the universal valuative invariant is the $\mathcal{G}$-invariant, to which we turn. For a rank-$r$ matroid $M$ and permutation $\pi=e_{1}e_{2}\ldots e_{n}$ of $E(M)$, the associated rank sequence is the sequence $\underline{r}(\pi)=r_{1}r_{2}\ldots r_{n}$ where $r_{1}=r(\{e_{1}\})$ and for $j\geq 2$,
$r_{j}=r(\{e_{1},e_{2},\ldots,e_{j}\})-r(\{e_{1},e_{2},\ldots,e_{j-1}\}).$
Thus, $r$ entries of $\underline{r}(\pi)$ are $1$ and nrn-r are 0; we call such sequences $(n,r)$-sequences. The rank sequence $\underline{r}(\pi)$ shows the rank increases as one adjoins the elements of $E(M)$, one at a time, to $\emptyset$ in the order given by $\pi$. The set $\{e_{i}\,:\,r_{i}=1\}$ is a basis of $M$.
Fix integers $n$ and $r$ with $0\leq r\leq n$. For each $(n,r)$-sequence $\underline{r}$, let $\underline{r}]$ be a formal symbol. Let $\mathcal{G}(n,r)$ be the vector space, over a field $F$ of characteristic zero, that consists of all formal linear combinations, with coefficients in $F$, of the symbols $\underline{r}]$. For a matroid $M$ of rank $r$ on $n$ elements, its $\mathcal{G}$-invariant and its coefficients $g_{\underline{r}}(M)$ are defined by
(2.2) $\mathcal{G}(M)=\sum_{\underline{r}}g_{\underline{r}}(M)[\underline{r}]=\sum_{\pi}[\underline{r}(\pi)]$
where the second sum is over all $n!$ permutations $\pi$ of $E(M)$. Thus, $g_{\underline{r}}(M)$ is the number of permutations $\pi$ of $E(M)$ for which $\underline{r}(\pi)=\underline{r}$. One can think of $\mathcal{G}(M)$ as a generating function for rank sequences.
As mentioned above, Derksen and Fink [16] showed that the $\mathcal{G}$-invariant is the universal valuative invariant: one obtains any valuative invariant $f$ with values in an abelian group $A$ by assigning values in $A$ to each symbol $\underline{r}]$ and extending linearly, that is, getting $f(M)$ by replacing each symbol $\underline{r}(\pi)]$ in Equation (2.2) by its image $f([\underline{r}(\pi)])$ in $A$. For example, for $A=\mathbb{Z}$ (the case of interest for this paper) and a fixed $(n,r)$-sequence $\underline{r}$, by mapping $\underline{r}]$ to $1$ and all other rank sequences $\underline{r}^{\prime}]$ to 0, we see that the function mapping $M$ to $g_{\underline{r}}(M)$ is valuative. Also, integer linear combinations of valuations with values in $\mathbb{Z}$ are valuations. Map each $\underline{r}]$ to $\underline{r}^{*}]$ where we obtain $\underline{r}^{*}$ from $\underline{r}$ by replacing each $1$ by 0, and each 0 by $1$, and reversing the order of the entries; it is routine to check that $\mathcal{G}(M)$ then maps to $\mathcal{G}(M^{*})$. Thus, valuative invariants of $M^{*}$ are also valuative invariants of $M$.
Derksen [15] showed that the Tutte polynomial is a valuative invariant. From this and the definition of a valuation, it follows easily that, for fixed nonnegative integers $i$ and $j$, the map that sends a matroid $M$ to the coefficient of $x^{i}y^{j}$ in its Tutte polynomial $T(M;x,y)$ is a valuative invariant. The following theorem collects these results along with valuative invariants that were identified in [9, Theorem 5.6] and that we will use heavily.
Theorem 2.5.
Each of the following is a valuative invariant:
  1. (1)
    the map $M\mapsto\mathcal{G}(M)$,
  2. (2)
    the map $M\mapsto\mathcal{G}(M^{*})$,
  3. (3)
    for any rank sequence $\underline{r}$, the map $M\mapsto g_{\underline{r}}(M)$,
  4. (4)
    the map $M\mapsto T(M;x,y)$,
  5. (5)
    for $i,j\in\mathbb{N}_{0}$, the map $M\mapsto t_{ij}$ where $t_{ij}$ is the coefficient of $x^{i}y^{j}$ in $T(M;x,y)$,
  6. (6)
    the map $M\mapsto|\mathcal{I}(M)|$, where $\mathcal{I}(M)$ is the set of independent sets of $M$,
  7. (7)
    for any $h,k,s_{h},s_{h+1},\ldots,s_{k}\in\mathbb{N}_{0}$ for which $0\leq h\leq k\leq r(M)$, the map $M\mapsto F_{h,k}(M;s_{h},\ldots,s_{k})$, where $F_{h,k}(M;s_{h},\ldots,s_{k})$ is the number of flags $F_{h}\subsetneq F_{h+1}\subsetneq\cdots\subsetneq F_{k}$ of flats of $M$ with $r(F_{i})=i$ and $|F_{i}|=s_{i}$ for all $i$ with $h\leq i\leq k$,
  8. (8)
    the map $M\mapsto f_{k}(M;s)$, where $f_{k}(M;s)$ is the number of rank-$k$ size-$s$ flats of $M$, and
  9. (9)
    the map $M\mapsto f_{k}(M;s,c)$, where $f_{k}(M;s,c)$ is the number of rank-$k$ size-$s$ flats $X$ of $M$ for which $M|X$ has $c$ coloops.
Item (8) is a special case of item (7). By combining item (2) with other items, it follows that the number of circuits of $M$ of a given size is a valuative invariant, as is the number of cyclic sets of a given size and nullity. The number of cyclic flats of $M$ of a given rank and size is a valuative invariant; set $c=0$ in item (9). The number of circuit-hyperplanes is a valuative invariant; in a rank-$r$ matroid $M$, it is $f_{r-1}(M;r,0)$. Item (6) is a well-known evaluation of the Tutte polynomial, as is the number of bases.
2.4. Extremal matroids of the polytope of unlabeled matroids
To associate analogs of matroid base polytopes to unlabeled matroids, Ferroni and Fink use the symmetrized indicator function $\mathbb{I}(X):\mathbb{R}^{n}\to\mathbb{R}$ of a subset $X$ of $\mathbb{R}^{n}$, which is given by
$\mathbb{I}(X)(x_{1},x_{2},\ldots,x_{n})=\frac{1}{n!}\sum_{\sigma\in S_{n}}\bar{\mathbb{I}}(X)(x_{\sigma(1)},x_{\sigma(2)},\ldots,x_{\sigma(n)})$
where $S_{n}$ is the symmetric group on $[n]$. (The bar over $\mathbb{I}$ is dropped in the symmetrized indicator function.) Note that if $M$ and $N$ are isomorphic but unequal matroids on $[n]$, then $\bar{\mathbb{I}}(\mathcal{P}(M))\neq\bar{\mathbb{I}}(\mathcal{P}(N))$ but $\mathbb{I}(\mathcal{P}(M))=\mathbb{I}(\mathcal{P}(N))$. Thus, letting the symmetrized indicator function of an unlabeled matroid be the symmetrized indicator function of any (labeled) representative of its isomorphism class makes sense. Theorem 2.4 carries over to unlabeled matroids, but we must take into account that some terms may merge. For instance, for the matroid $M$ in Figure 1, for $\bar{\mathbb{I}}(\mathcal{P}(M))$, there are two nested matroids whose cyclic flats are $\emptyset$, $E(M)$, and a set of size two and rank $1$ (i.e., $A$ or $B$), and they give different terms in Equation (2.1), each with coefficient 1-1; however, for $\mathbb{I}(\mathcal{P}(M))$, they give the same term, namely, the symmetrized indicator function of the unlabeled nested matroid that is encoded by the lattice path $NENNEEEE$, and its coefficient is 2-2. Theorem 2.4 implies that the symmetrized indicator function of any unlabeled rank-$r$ matroid on $[n]$ can be written in exactly one way as an integer linear combination of the symmetrized indicator functions of unlabeled rank-$r$ nested matroids on $[n]$. Fix a linear order on the set $\mathcal{N}$ of all $\binom{n}{r}$ unlabeled rank-$r$ nested matroids on $[n]$. Let $U$ be the set of coordinate vectors of the symmetrized indicator functions of unlabeled rank-$r$ matroids on $[n]$ relative to those of the matroids in $\mathcal{N}$, so $U\subsetneq\mathbb{R}^{\binom{n}{r}}$. The convex hull of $U$ is the polytope $\Omega_{r,n}$ of unlabeled rank-$r$ matroids on $[n]$. A matroid that gives a vertex in $\Omega_{r,n}$ is called extremal. The starting point for our work is the following result from [20, Lemma 4.3].
Theorem 2.6.
Fix integers $r$ and $n$ with $0\leq r\leq n$. Let $f_{1},f_{2},\ldots,f_{t}$ be real-valued valuative invariants. Let the set $S_{0}$ consist of one matroid of each isomorphism type among rank-$r$ matroids on $n$ elements, and, for each $i\in[t]$, set
$S_{i}=\{M\in S_{i-1}\,:\,f_{i}(M)\geq f_{i}(N)\text{ for all }N\in S_{i-1}\}.$
If all matroids in $S_{t}$ yield the same point in the polytope $\Omega_{r,n}$, then that point is a vertex of $\Omega_{r,n}$, so all matroids in $S_{t}$ are then extremal.
Theorem 2.6 motivates characterizing certain matroids by maximizing a sequence of real-valued valuative invariants. Such characterizations are the focus of this paper. Note that $f$ is valuative if and only if f-f is, so we can also minimize real-valued valuative invariants. In some settings maximizing is more natural, while in others minimizing is more natural, so we refer to the generic case as optimizing valuative invariants.
As noted at the beginning of [20, Section 4.3], a matroid $M$ is extremal for $\Omega_{r,n}$ if and only if $M^{*}$ is extremal for $\Omega_{n-r,n}$. Valuative invariants of $M$ are valuative invariants of $M^{*}$, so a characterization of $M$ by optimizing valuative invariants also gives such a characterization of $M^{*}$.
3. Perfect matroid designs
A perfect matroid design is a matroid $M$ in which flats of the same rank have the same size, that is, there are integers $\alpha_{0},\alpha_{1},\ldots,\alpha_{r(M)}$ so that $|F|=\alpha_{r(F)}$ for all flats $F$ of $M$. (See [17] and [30].) Examples include (finite) projective and affine geometries and their truncations, as well as paving matroids of rank $r$ on $[n]$ for which the hyperplanes are the blocks of a Steiner system S(r1,k,n)S(r-1,k,n) (i.e., a set of $k$-subsets (blocks) of $[n]$ for which each (r1)(r-1)-subset of $[n]$ is a subset of exactly one block).
Nonisomorphic projective planes of order $q$ show that nonisomorphic perfect matroid designs can have the same sequence of flat sizes. In contrast, any perfect matroid design that has the same sequence of flat sizes as the projective geometry $\operatorname{PG}(r-1,q)$ with $r\geq 4$ satisfies the axioms for projective geometries and so is isomorphic to $\operatorname{PG}(r-1,q)$, and likewise for the affine geometry $\operatorname{AG}(r-1,q)$. (See, e.g., [3].)
By [9, Example 3.4], the $\mathcal{G}$-invariant of a perfect matroid design $M$ is determined by its sequence $\alpha_{0},\alpha_{1},\ldots,\alpha_{r(M)}$ of flat sizes. Thus, $f_{i}(M;\alpha_{i})$, the number of rank-$i$ flats, is determined by $\alpha_{0},\alpha_{1},\ldots,\alpha_{r(M)}$. Also, since perfect matroid designs with the same sequence of flat sizes have the same $\mathcal{G}$-invariant, valuative invariants cannot distinguish between them; indeed, as we show next as part of proving that perfect matroid designs are extremal, they determine the same point in the polytope of unlabeled matroids. The valuative invariant used below is defined in Theorem 2.5.
Theorem 3.1.
Perfect matroid designs are extremal. If perfect matroid designs having the sequence of flat sizes $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$ exist, then they are precisely the rank-$r$ matroids on $\alpha_{r}$ elements that maximize the valuative invariant $N\mapsto F_{0,r}(N;\alpha_{0},\alpha_{1},\ldots,\alpha_{r})$. In particular, for $r\geq 4$ and a prime power $q$, among all simple rank-$r$ matroids on $\frac{q^{r}-1}{q-1}$ elements, $\operatorname{PG}(r-1,q)$ is the unique matroid that maximizes the valuative invariant
$N\mapsto F_{0,r}(N;0,1,q+1,q^{2}+q+1,\ldots,\frac{q^{r}-1}{q-1}),$
and among simple rank-$r$ matroids on $q^{r-1}$ elements, $\operatorname{AG}(r-1,q)$ is the unique matroid that maximizes the valuative invariant $N\mapsto F_{0,r}(N;0,1,q,q^{2},\ldots,q^{r-1})$.
Proof.
Recall that each element that is not in a flat $F$ is in exactly one flat that covers $F$. Thus, if $0\leq i<r$ and $F$ is a rank-$i$ flat of $M$ with $|F|=\alpha_{i}$, then at most $\frac{\alpha_{r}-\alpha_{i}}{\alpha_{i+1}-\alpha_{i}}$ flats cover $F$ and have size $\alpha_{i+1}$, and there are $\frac{\alpha_{r}-\alpha_{i}}{\alpha_{i+1}-\alpha_{i}}$ such covers if and only if all covers of $F$ have size $\alpha_{i+1}$. It follows that
$F_{0,r}(M;\alpha_{0},\alpha_{1},\ldots,\alpha_{r})\leq\prod_{i=0}^{r-1}\frac{\alpha_{r}-\alpha_{i}}{\alpha_{i+1}-\alpha_{i}},$
with equality if and only if $M$ is a perfect matroid design with the sequence $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$ of flat sizes.
[…]
It is easy to see that, for $i\in[r]$, the rank-$i$ flats are cyclic if and only if $\alpha_{i}>\alpha_{i-1}+1$. Also, for a rank-$i$ flat $F$ of $M$, the contraction $M/F$ is a perfect matroid design and its sequence of flat sizes is $0,\alpha_{i+1}-\alpha_{i},\ldots,\alpha_{r}-\alpha_{i}$. Thus, the number of rank-$j$ flats of $M$ that contain $F$ is determined by $0,\alpha_{i+1}-\alpha_{i},\ldots,\alpha_{r}-\alpha_{i}$, and so by $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$. Statement (a) above now follows. It also follows that, for a chain $C\in\mathcal{C}(\mathcal{Z}(M))$, the number of chains $C^{\prime}\in\mathcal{C}(\mathcal{Z}(M))$ with $C\subseteq C^{\prime}$ and specified ranks for the flats in $C^{\prime}$ depends only on $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$, the ranks of the flats in $C$, and the ranks specified for $C^{\prime}$. Statement (b) above now follows by a routine induction, starting with the largest chains $C\in\mathcal{C}(\mathcal{Z}(M)$. Thus, all perfect matroid designs that share the sequence $\alpha_{0},\alpha_{1},\ldots,\alpha_{r}$ yield the same point in $\Omega_{r,\alpha_{r}}$ and so are extremal. ∎
4. Bose-Burton geometries
[…]
Theorem 4.1.
If $M$ is a restriction of $\operatorname{PG}(r-1,q)$ that has no restriction isomorphic to $\operatorname{PG}(m-1,q)$, then
$|E(M)|\leq\frac{q^{r}-q^{r-m+1}}{q-1}.$
[…]
Theorem 4.2.
Let $q$ be a prime power. Fix $r,k\in\mathbb{N}$ with $r\geq 4$ and $1\leq k\leq r-3$. Let $M$ be a simple rank-$r$ matroid on $(q^{r}-q^{k})/(q-1)$ elements that satisfies conditions (1)–(3).
  • (1)
    If $F$ is a rank-(rk+1)(r-k+1) flat of $M$, then $|F|<(q^{r-k+1}-1)/(q-1)$.
  • (2)
    If $H$ is a hyperplane of $M$, then $|H|=(q^{r-1}-q^{i})/(q-1)$ where $i\in\{k,k-1\}$.
  • (3)
    If $F$ is a rank-(r2)(r-2) flat of $M$, then $|F|=(q^{r-1}-q^{i})/(q-1)$ where $i$ is in $\{k,k-1,k-2\}$ and $i$ can be k2k-2 only if $k>1$.
[…]
Theorem 4.3.
[…]
Proof.
[…]
Now assume that each line of a rank-$3$ simple matroid $M$ on $q^{2}+q$ elements has size $q$ or $q+1$. If all lines that contain $e\in E(M)$ had $q+1$ points, then $|E(M)-\{e\}|=q^{2}+q-1$ would be divisible by $q$, but it is not. Thus, each element is in at least one $q$-point line. Along with the factorization $q^{2}+q=q(q+1)$, that implies that at least $q+1$ lines have $q$ points each, as claimed. That factorization also implies that $M$ has $q+1$ lines with $q$ points each if and only if each element of $M$ is in one such line. In that case the $q$-point lines must be pairwise disjoint and so, along with $E(M)$, they form a modular cut. The corresponding single-element extension, say $N$, has $q^{2}+q+1$ points and all lines have $q+1$ points, from which it follows that $N$ is a projective plane of order $q$.
There may be nonisomorphic projective planes of order $q$ if $q\geq 9$, so to complete the proof that their single-element deletions are extremal, we show that, for $q\geq 9$, all single-element deletions of all projective planes of order $q$ give the same point in $\Omega_{3,q^{2}+q}$. For a single-element deletion $M$ of a projective plane of order $q$, its indicator function $\bar{\mathbb{I}}(\mathcal{P}(M))$ depends on $M$, but, as we show next, from $q$ alone we get the relevant unlabeled nested matroids, their multiplicities, and their coefficients, so the symmetrized indicator function $\mathbb{I}(\mathcal{P}(M))$, and hence the corresponding point in $\Omega_{3,q^{2}+q}$, depends only on $q$, not on $M$. The lattice of cyclic flats of $M$ consists of $\emptyset$, $q+1$ lines that each have $q$ points, $q^{2}$ lines that each have $q+1$ points, and $E(M)$. Thus, the lattice $\mathcal{C}(\mathcal{Z}(M))$ has, besides $\hat{1}$,
  • $q+1$ elements $C=\{\emptyset,L,E(M)\}$ covered by $\hat{1}$ where $r(L)=2$ and $|L|=q$; for each, $\mu(C,\hat{1})=-1$,
  • $q^{2}$ elements $C=\{\emptyset,L,E(M)\}$ covered by $\hat{1}$ where $r(L)=2$ and $|L|=q+1$; for each, $\mu(C,\hat{1})=-1$,
  • the least chain, $C=\{\emptyset,E(M)\}$; given the values above, including $\mu(\hat{1},\hat{1})=1$, we have $\mu(C,\hat{1})=q^{2}+q$.
Thus, all single-element deletions of projective planes of order $q$ give the same point in $\Omega_{3,q^{2}+q}$. ∎
We treat $r\geq 4$ in Theorem 4.5 below. The proof uses the result below from [5].
Theorem 4.4.
Let $r\geq 4$ be an integer and let $q$ be a prime power. Any simple rank-$r$ matroid that has no minor isomorphic to $U_{2,q+2}$, the $(q+2)$-point line, and that has at least $q^{r-1}$ points is representable over $\operatorname{GF}(q)$.
Theorem 4.5.
Fix a prime power $q$. Assume that $M$ is a simple matroid of rank $r\geq 4$ on $q^{r-1}+q^{r-2}$ elements that satisfies the following conditions:
  • (1)
    if $2\leq i<r$ and $A$ is a rank-$i$ flat of $M$, then $|A|\in\{q^{i-1},q^{i-1}+q^{i-2}\}$, and
  • (2)
    $F_{r-2,r-1}(M;q^{r-3}+q^{r-4},q^{r-2})=0$, that is, no rank-(r2)(r-2) flat with $q^{r-3}+q^{r-4}$ elements is a subset of a hyperplane with $q^{r-2}$ elements.
[…]
Proof.
We first show that $\X\operatorname{PG}(r-1,q)\backslash$ X, where $X$ is a rank-(r2)(r-2) flat of $\operatorname{PG}(r-1,q)$, has the properties above. For any flat $F$ of $\operatorname{PG}(r-1,q)$ where $F\not\subseteq X$, if $r(F\cup X)=r-1$, then $r(F\cap X)=r(F)-1$, and otherwise $r(F\cap X)=r(F)-2$; these two cases yield the options for |FX||F-X| in condition (1). Now let $F$ be a rank-(r2)(r-2) flat of $\operatorname{PG}(r-1,q)$ for which $|F-X|=q^{r-3}+q^{r-4}$, so $r(X\cap F)=r-4$, and so $r(X\cup F)=r$. Thus, if $H$ is a hyperplane of $\operatorname{PG}(r-1,q)$ and $F\subsetneq H$, then $X\not\subseteq H$, so $r(H\cap X)=r-3$, and so $|H-X|=q^{r-2}+q^{r-3}$; thus, condition (2) holds. The $q^{r-2}$-point hyperplanes of $\X\operatorname{PG}(r-1,q)\backslash$ X are the sets HXH-X where $H$ ranges over the hyperplanes of $\operatorname{PG}(r-1,q)$ that contain $X$, so there are $q+1$ such hyperplanes.
Now assume that $M$ satisfies conditions (1) and (2). To prove the lower bound of $q+1$ on the number of $q^{r-2}$-point hyperplanes, note that if, for some $i\geq 1$, the element $e$ is in a rank-$i$ flat $F$ with $q^{i-1}$ elements and i<r1i<r-1, then having all rank-$(i+1)$ flats that contain $F$ have $q^{i}+q^{i-1}$ elements would yield an integer solution $k$ of the equation $q^{r-1}+q^{r-2}=q^{i-1}+kq^{i}$, and so of $q^{r-i}+q^{r-i-1}=1+kq$, which has no integer solution; thus, $e$ is in some rank-$(i+1)$ flat with $q^{i}$ elements. Thus, by induction, each $e\in E(M)$ is in at least one $q^{r-2}$-point hyperplane. The lower bound on the number of $q^{r-2}$-point hyperplanes now follows since the fewest number of sets of size $q^{r-2}$ whose union is $E(M)$ is $q+1$ since $q^{r-1}+q^{r-2}=q^{r-2}(q+1)$. Since each element is in at least one $q^{r-2}$-point hyperplane, that factorization also shows that $M$ has $q+1$ such hyperplanes if and only if each element of $M$ is in one such hyperplane.
Now assume that $M$ has $q+1$ hyperplanes of size $q^{r-2}$, so each element of $M$ is in one such hyperplane. We next use Theorem 4.4 to show that $M$ is representable over $\operatorname{GF}(q)$. To show that $M$ has no $U_{2,q+2}$-minor, by the Scum Theorem it suffices to show that each rank-(r2)(r-2)-flat $F$ of $M$ is contained in at most $q+1$ hyperplanes. Each element is in one $q^{r-2}$-point hyperplane, so at most one such hyperplane contains $F$. First assume that $|F|=q^{r-3}$ and that $i$ hyperplanes that contain $F$ have $q^{r-2}$ elements, so $i\in\{0,1\}$. The number of $(q^{r-2}+q^{r-3})$-point hyperplanes that contain $F$ is then
$\frac{q^{r-1}+q^{r-2}-q^{r-3}-i(q^{r-2}-q^{r-3})}{q^{r-2}}.$
Canceling common factors shows that $q$ divides $q^{2}+q-1-i(q-1)$, so $q$ divides i1i-1. Thus, $i=1$ and the displayed quantity simplifies to $q$, so $F$ is in $q+1$ hyperplanes, as needed. Now assume that $|F|=q^{r-3}+q^{r-4}$. By conditions (1) and (2), each hyperplane that contains $F$ has size $q^{r-2}+q^{r-3}$, so the number of such hyperplanes is
$\frac{q^{r-1}+q^{r-2}-(q^{r-3}+q^{r-4})}{q^{r-2}+q^{r-3}-(q^{r-3}+q^{r-4})}=q+1,$
so $F$ is contained in $q+1$ hyperplanes. Thus, by the Scum Theorem and Theorem 4.4, $M$ is representable over $\operatorname{GF}(q)$ and so is isomorphic to a restriction of $\operatorname{PG}(r-1,q)$.
[…]
5. Cycle matroids of complete graphs, and their truncations
To prove that $M(K_{r+1})$ is extremal, we use the characterization of $M(K_{r+1})$ below from [10, Theorem 3.2], which does not have the form set out in Theorem 2.6 since the hypotheses include the numbers of $3$-point lines and $6$-point planes without establishing that they are the greatest numbers of such flats.
Theorem 5.1.
The cycle matroid $M(K_{r+1})$ is the unique rank-$r$ simple matroid on $r+1\choose 2$} elements that has
  • (1)
    $r+1\choose 3$} lines with three points,
  • (2)
    no $5$-point planes, $r+1\choose 4$} planes with six points, no planes with more than six points, and
  • (3)
    no rank-$4$ flats with more than ten points.
As discussed in Section 1, all criteria in the next result optimize valuative invariants, so this shows that $M(K_{r+1})$ is extremal, as conjectured in [20]. This result sets the pattern for later results in the paper in that since it is immediate that the matroid of interest, $M(K_{r+1})$ in this case, satisfies the conditions in the result, we do not explicitly state or prove that.
Theorem 5.2.
Assume that $M$ is a simple rank-$r$ matroid on $\binom{r+1}{2}$ elements, that all lines have $2$ or $3$ elements, that all planes have $3$, 44, or $6$ elements, and that, for all $h$ with $4\leq h<r$, all rank-$h$ flats of $M$ have at most $\binom{h+1}{2}$ elements. Then at most $\binom{r+1}{3}$ lines of $M$ have three elements, and if $\binom{r+1}{3}$ lines of $M$ have three elements, then $M$ is $M(K_{r+1})$. Thus, $M(K_{r+1})$ is extremal.
Proof.
In this proof, we call $3$-point lines large lines. The statement clearly holds for $r\leq 2$, so now assume that $r\geq 3$.
Let $F$ be a rank-$h$ flat of $M$ where $3\leq h\leq r$. Let an element $e\in F$ be in $m$ large lines in $F$, say $L_{i}$, where $i\in[m]$. We claim that $m\leq h-1$. If $h=3$, then $1+2m\leq|F|\leq 6$, so $m\leq 2$, as needed. Now consider $h=4$. No three large lines are coplanar by the case $h=3$. Since no plane of $M$ has five elements, each plane $\operatorname{cl}(L_{i}\cup L_{j})$ contains one point $a_{i,j}$ that is in neither $L_{i}$ nor $L_{j}$. The $1+2m$ elements in the lines $L_{j}$, for $j\in[m]$, and the m1m-1 distinct points $a_{i,m}$, for $i\in[m-1]$, give $3m\leq|F|\leq 10$, so $m\leq 3$, as needed. Now consider $h>4$. As above, each difference $\operatorname{cl}(L_{i}\cup L_{j})-(L_{i}\cup L_{j})$ is a singleton, say $\{a_{i,j}\}$. We claim that, apart from the obvious equality $a_{i,j}=a_{j,i}$, all $a_{i,j}$ are distinct. If $a_{i,j}=a_{i,k}$, then the plane $\operatorname{cl}(L_{i}\cup\{a_{i,j}\})$ would contain $L_{i}$, $L_{j}$, and $L_{k}$, contrary to the case $h=3$. If $a_{i,j}=a_{s,t}$ with $\{i,j\}\cap\{s,t\}=\emptyset$, then the rank-$4$ flat $\operatorname{cl}(L_{i}\cup L_{s}\cup\{a_{i,j}\})$ would contain $L_{i}$, $L_{j}$, $L_{s}$, and $L_{t}$, contrary to the case $h=4$. Since all $a_{i,j}$ are distinct, the rank-$h$ flat $F$ contains at least $1+2m+\binom{m}{2}$ elements, so $1+2m+\binom{m}{2}\leq\binom{h+1}{2}$, that is, $\binom{m+2}{2}\leq\binom{h+1}{2}$, so $m\leq h-1$, as claimed.
Let $M$ have $k$ large lines. The number of pairs $(L,e)$, where $L$ is a large line and $e\in L$, is $3k$. There are $\binom{r+1}{2}$ points $e$ and each is in at most r1r-1 large lines, so $3k\leq\binom{r+1}{2}(r-1)$, that is, $k\leq\binom{r+1}{3}$, as the bound in the theorem asserts.
Now assume that $M$ has $\binom{r+1}{3}$ large lines. The counting arguments above show that each element $e$ of $M$ is in r1r-1 large lines, any h1h-1 of which span a flat of rank $h$. Also, $E(M)$ is the union of the large lines $L_{1},L_{2},\ldots,L_{r-1}$ that contain $e$ along with the $\binom{r-1}{2}$ singletons $\{a_{i,j}\}=\operatorname{cl}(L_{i}\cup L_{j})-(L_{i}\cup L_{j})$ for distinct $i,j\in[r-1]$. This notation applies throughout the next three paragraphs, which deduce more about the $6$-point planes of $M$.
We claim that, for each $i\in[r-1]$, the planes $\operatorname{cl}(L_{i}\cup L_{j})$, for $j\in[r-1]-\{i\}$, account for all $6$-point planes that contain $L_{i}$. Since $\operatorname{cl}(L_{i}\cup L_{j})=\operatorname{cl}(L_{i}\cup\{a_{i,j}\})$, each other plane that contains $L_{i}$ is $\operatorname{cl}(L_{i}\cup\{a_{s,t}\})$ for some $\{s,t\}\subseteq[r-1]-\{i\}$. Since $\operatorname{cl}(L_{i}\cup\{a_{s,t}\})$ is in the rank-$4$ flat $\operatorname{cl}(L_{i}\cup L_{s}\cup L_{t})$, and the two $6$-point planes $\operatorname{cl}(L_{i}\cup L_{s})$ and $\operatorname{cl}(L_{i}\cup L_{t})$ contain nine of the ten points in $\operatorname{cl}(L_{i}\cup L_{s}\cup L_{t})$, the plane $\operatorname{cl}(L_{i}\cup\{a_{s,t}\})$ is $L_{i}\cup\{a_{s,t}\}$ and so has only four points, which proves the claim.
Similarly, rank-$4$ flats that contain $L_{i}\cup L_{j}$ besides those of the form $\operatorname{cl}(L_{i}\cup L_{j}\cup L_{k})$ have the form $F=\operatorname{cl}(L_{i}\cup L_{j}\cup\{a_{s,t}\})$ for some $\{s,t\}\subseteq[r-1]-\{i,j\}$. Now $F$ is contained in the rank-$5$ flat $F^{\prime}=\operatorname{cl}(L_{i}\cup L_{j}\cup L_{s}\cup L_{t})$, and $\operatorname{cl}(L_{i}\cup L_{j}\cup L_{s})$ and $\operatorname{cl}(L_{i}\cup L_{j}\cup L_{t})$ contain $14$ of the $15$ points in $F^{\prime}$, so $|F|=7$. Thus, for all $\{i,j\}\subseteq[r-1]$, the flats $\operatorname{cl}(L_{i}\cup L_{j}\cup L_{k}),$ for $k\in[r-1]-\{i,j\}$, are the only $10$-element rank-44 flats that contain $L_{i}\cup L_{j}$.
We claim that if $P$ is a $6$-point plane with $\{e,a_{i,j}\}\subsetneq P$, then $P=\operatorname{cl}(L_{i}\cup L_{j})$. If not, then $P\cap\operatorname{cl}(L_{i}\cup L_{j})=\{e,a_{i,j}\}$. Set $F=P\cup\operatorname{cl}(L_{i}\cup L_{j})$. Now $|F|=10$ and $r(F)=4$, so $F$ is a flat, and so $F=\operatorname{cl}(L_{i}\cup L_{j}\cup L_{k})$, for some $k\in[r-1]-\{i,j\}$, by the previous paragraph. Thus, $L_{k}\subsetneq P$ and $L_{h}\not\subseteq P$ for all $h\in[r-1]-\{k\}$, contrary to what we proved two paragraphs above. Thus, the $6$-point planes that contain $e$ are the planes $\operatorname{cl}(L_{i}\cup L_{j})$, for $\{i,j\}\subseteq[r-2]$. Any point in a $6$-point plane can be $e$, so each such plane contains four large lines.
Let $k$ be the number of $6$-point planes of $M$. Consider the triples $(P,L,e)$ where $P$ is a $6$-point plane, $L$ is a large line in $P$, and $e\in L$. The number of such triples is $k\cdot 4\cdot 3=12k$ by the conclusion above. By focusing on $e$ first, and then $L$, and finally $P$, we see that the number of triples is $\binom{r+1}{2}(r-1)(r-2)$, where r2r-2 accounts for choosing a second line containing $e$ to determine the plane $P$. Thus, $12k=\binom{r+1}{2}(r-1)(r-2)$, so $k=\binom{r+1}{4}$.
This shows that the criteria in Theorem 5.1 hold, so $M$ is $M(K_{r+1})$. ∎
The next result gives two more characterizations of $M(K_{r+1})$ by optimizing valuative invariants. The second characterization below suits our work in Section 8 on direct sums. The proof uses Theorem 5.2.
Theorem 5.3.
Let $M$ be a simple matroid of rank $r\geq 2$ on $\binom{r+1}{2}$ elements. Assume that for all $i\in[r-1]$ and rank-$i$ flats $F$, either $|F|=\binom{i+1}{2}$ or $|F|\leq\binom{i}{2}+1$. Call a rank-$i$ flat $F$ large if $|F|=\binom{i+1}{2}$, and call a flag $F_{0}\subsetneq F_{1}\subsetneq F_{2}\subsetneq\cdots\subsetneq F_{r}$ of flats large if each $F_{i}$ is large. Each conclusion below shows that $M(K_{r+1})$ is extremal.
  • (a)
    The number of large hyperplanes is at most $r+1$, and if it is $r+1$, then $M$ is $M(K_{r+1})$.
  • (b)
    The number of large flags is at most $(r+1)!/2$, and if it is $(r+1)!/2$, then $M$ is $M(K_{r+1})$.
Proof.
We prove the bound in (a) by induction on $r$. An inequality that we use in the induction step holds for $r\geq 5$, so we check all cases $r<5$ first. If $r=2$, then $M$ is $U_{2,3}$, so the assertion holds. For $r=3$, since $|E(M)|=6$, each point is in at most two large lines; thus, if there are $k$ large lines, then the number of pairs $(e,L)$ where $L$ is a large line and $e\in L$ is, on the one hand, $3k$, and, on the other, at most $6\cdot 2$, so $3k\leq 12$, that is, $k\leq 4$, as needed.
Now consider $r=4$, so $|E(M)|=10$. Note that the intersection of any two large planes $P$ and $P^{\prime}$ is a line with two or three elements. First assume that $P\cap P^{\prime}=\{a,b\}$. Thus, $P\cup P^{\prime}=E(M)$, so any other large plane $Q$ must intersect each of $P$ and $P^{\prime}$ in a large line; let $Q\cap P=L$ and $Q\cap P^{\prime}=L^{\prime}$. Now $Q=L\cup L^{\prime}$, so $L\cap L^{\prime}=\emptyset$, and so $\{a,b\}\cap(L\cup L^{\prime})=\emptyset$. However, only one large line in $P$ can be disjoint from $\{a,b\}$, and similarly for $P^{\prime}$, so there is at most one large plane besides $P$ and $P^{\prime}$, consistent with the bound. Now assume that any two large planes of $M$ intersect in a large line. By counting, a large line is in at most two large planes. Since any large plane $P$ has at most four large lines and each other large plane must intersect $P$ in one of those lines, and each such line is in at most one large plane besides $P$, there are at most five large planes, as needed. It is relevant for the last step of the proof to note that if $M$ has five large planes, then any two of them intersect in a large line.
Now assume that $r\geq 5$ and that the bound on the number of large hyperplanes holds in the case of rank r1r-1. Let $H_{1},H_{2},\ldots,H_{t}$ be the large hyperplanes of $M$. For any $\{i,j\}\subseteq[t]$, we have $|H_{i}\cup H_{j}|\leq|E(M)|=\binom{r+1}{2}$, so
$|H_{i}\cap H_{j}|$ $\,=|H_{i}|+|H_{j}|-|H_{i}\cup H_{j}|$
$\,\geq 2\binom{r}{2}-\binom{r+1}{2}$
$\,=\binom{r-1}{2}-1$
$\,>\binom{r-2}{2}+1$
where the last line holds since $r\geq 5$. Therefore $H_{i}\cap H_{j}$ is a large rank-(r2)(r-2) flat. Since $H_{i}-(H_{i}\cap H_{j})$ and $H_{j}-(H_{i}\cap H_{j})$ each contain $\binom{r}{2}-\binom{r-1}{2}=r-1$ out of the $\binom{r+1}{2}-\binom{r-1}{2}=2r-1$ elements in $E(M)-(H_{i}\cap H_{j})$, it follows that $H_{i}$ and $H_{j}$ are the only large hyperplanes that contain $H_{i}\cap H_{j}$, so all $\binom{t}{2}$ intersections $H_{i}\cap H_{j}$, with $\{i,j\}\subseteq[t]$, are different. Thus, $H_{i}\cap H_{t}$, for $i\in[t-1]$, are t1t-1 different large hyperplanes of $M|H_{t}$, so the induction assumption applied to $M|H_{t}$ gives $t-1\leq r$, so $t\leq r+1$, which completes the proof of the bound stated in (a).
The bound of $(r+1)!/2$ on the number of large flags $F_{0}\subsetneq F_{1}\subsetneq\cdots\subsetneq F_{r}$ now follows since there are at most $r+1$ options for $F_{r-1}$, and then at most $r$ options for $F_{r-2}$, and so on, ending with at most three options for $F_{1}$.
The argument shows that in order for $M$ to have $(r+1)!/2$ large flags, it must have $r+1$ large hyperplanes, so proving the second part of (a) will also complete the proof of (b). Let $H_{1},H_{2},\ldots,H_{r+1}$ be the large hyperplanes of $M$. As we showed, each $H_{i}$ contains $r$ large rank-(r2)(r-2) flats, namely, $H_{i}\cap H_{j}$ for $j\in[r+1]-\{i\}$. With that, a routine induction shows that if $I\subsetneq[r+1]$, then $\bigcap_{i\in I}H_{i}$ is a large rank-(r|I|)(r-|I|) flat. Thus, the only large hyperplanes that contain $\bigcap_{i\in I}H_{i}$ are $H_{i}$ with $i\in I$, so all such intersections are different. Therefore the intersections of r2r-2 large hyperplanes account for $\binom{r+1}{r-2}=\binom{r+1}{3}$ different large lines of $M$, and so $M$ is $M(K_{r+1})$ by Theorem 5.2. ∎
We now turn to truncations of $M(K_{r+1})$ of rank at least $5$. In [14], de Mier and Noy showed that these truncations are Tutte unique. For that, they first developed a numerical characterization of these matroids similar to that of $M(K_{r+1})$ in Theorem 5.1, adding an assumption on the number of independent sets. We characterize truncations of $M(K_{r+1})$ of rank at least $5$ as optimizing certain valuative invariants and so show that these matroids are extremal. The proof below does not rest on prior characterizations. Theorem 5.4 gives yet another characterization of $M(K_{r+1})$ (set $k=0$), although it is limited to $r\geq 5$.
Let $\operatorname{Tr}^{k}(N)$, where $0\leq k\leq r(N)$, denote the matroid obtained from $N$ by truncating $k$ times. Thus, $\operatorname{Tr}^{k}(N)$ has rank r(N)kr(N)-k, and its independent sets are those of $N$ that have size at most r(N)kr(N)-k.
Theorem 5.4.
Fix $r,k\in\mathbb{N}\cup\{0\}$ with $r-k\geq 5$. Let $M$ be a simple rank-(rk)(r-k) matroid on $\binom{r+1}{2}$ elements in which $|F|\leq\binom{r(F)+1}{2}$ for all flats $F\neq E(M)$ and no planes have five elements. Then at most $\binom{r+1}{3}$ lines of $M$ have three elements. If $\binom{r+1}{3}$ lines of $M$ have three elements, then $M$ has at most as many independent sets as $\operatorname{Tr}^{k}(M(K_{r+1}))$. If, in addition, $M$ has the same number of independent sets as $\operatorname{Tr}^{k}(M(K_{r+1}))$, then $M$ is $\operatorname{Tr}^{k}(M(K_{r+1}))$. Thus, $\operatorname{Tr}^{k}(M(K_{r+1}))$ is extremal.
Proof.
As earlier, we call $3$-point lines large lines. The argument in the second paragraph of the proof of Theorem 5.2 applies and shows that for any $e\in E(M)$ and $h\in[r-k-1]$, the element $e$ is in at most h1h-1 large lines in any rank-$h$ flat. Let $e$ be in $i$ large lines of $M$, say $L_{1},L_{2},\ldots,L_{i}$. Now $|L_{1}\cup L_{2}\cup\cdots\cup L_{i}|=2i+1$, and, by the argument in the proof of Theorem 5.2, the sets $\operatorname{cl}(L_{s}\cup L_{t})-(L_{s}\cup L_{t})$, for $1\leq s<t\leq i$, are disjoint singletons. Thus, $1+2i+\binom{i}{2}\leq|E(M)|=\binom{r+1}{2}$, from which we get $i\leq r-1$. With that, the argument in the third paragraph in the proof of Theorem 5.2 applies and shows that $M$ has at most $\binom{r+1}{3}$ large lines, as asserted in the theorem.
Now assume that $M$ has $\binom{r+1}{3}$ large lines. As in the proof of Theorem 5.2, it follows that each element $e$ of $M$ is in r1r-1 large lines, any h1h-1 of which span a rank-$h$ flat if h<rkh<r-k. Furthermore, $E(M)$ is the union of the r1r-1 large lines $L_{1},L_{2},\ldots,L_{r-1}$ that contain $e$ along with the $\binom{r-1}{2}$ singletons $\{a_{i,j}\}=\operatorname{cl}(L_{i}\cup L_{j})-(L_{i}\cup L_{j})$ for $\{i,j\}\subsetneq[r-1]$. As in the proof of Theorem 5.2, for each $i\in[r-1]$, the planes $\operatorname{cl}(L_{i}\cup L_{j})$, for $j\in[r-1]-\{i\}$, account for all $6$-point planes that contain $L_{i}$. Consider $\{i,j\}\subsetneq[r-1]$ and $f\in L_{i}-\{e\}$. Since $f$ is in a large line in the plane $\operatorname{cl}(L_{i}\cup L_{j})$, our previous conclusion, applied to $f$ and the large lines that contain it, shows that $f$ is in two large lines in the plane $\operatorname{cl}(L_{i}\cup L_{j})$. From this, it follows that each $6$-point plane of $M$ that contains at least one large line contains four large lines.
We next identify the elements and large lines of $M$ with those of $M(K_{r+1})$. Fix an element of $M$, label it $a_{r,r+1}$, and let $L_{1},L_{2},\ldots,L_{r-1}$ be the large lines that contain $a_{r,r+1}$. For $i$ and $j$ with $1\leq i<j<r$, let the singleton $\operatorname{cl}(L_{i}\cup L_{j})-(L_{i}\cup L_{j})$ be $\{a_{i,j}\}$. Arbitrarily label the elements in $L_{1}-\{a_{r,r+1}\}$ as $a_{1,r}$ and $a_{1,r+1}$. For each $i$ with $2\leq i<r$, let $L_{i}\cap\operatorname{cl}(\{a_{1,r},a_{1,i}\})=\{a_{i,r}\}$ and $L_{i}\cap\operatorname{cl}(\{a_{1,r+1},a_{1,i}\})=\{a_{i,r+1}\}$, so $L_{i}=\{a_{r,r+1},a_{i,r},a_{i,r+1}\}$. To show that this labeling agrees with the large lines of $M(K_{r+1})$, where $K_{r+1}$ has vertices $v_{1},v_{2},\ldots,v_{r+1}$ and we want $a_{i,j}$ to correspond to the edge $\{v_{i},v_{j}\}$, we want to show that
  • (a)
    for $1<i<j<r$, both $\{a_{i,r},a_{j,r},a_{i,j}\}$ and $\{a_{i,r+1},a_{j,r+1},a_{i,j}\}$ are lines,
  • (b)
    for $1\leq i<j<k<r$, the set $\{a_{i,j},a_{i,k},a_{j,k}\}$ is a line,
and that we have identified all large lines of $M$. For the set $\{a_{i,r},a_{j,r},a_{i,j}\}$ in assertion (a), note that the plane $\operatorname{cl}(\{a_{1,r},a_{i,r},a_{j,r}\})$ is contained in the rank-$4$ flat $\operatorname{cl}(L_{1}\cup L_{i}\cup L_{j})$ and it contains $a_{1,i}$ and $a_{1,j}$ but none of $L_{1}$, $L_{i}$, and $L_{j}$. Therefore $a_{i,j}$ must be the sixth element of the plane $\operatorname{cl}(\{a_{1,r},a_{i,r},a_{j,r}\})$. Thus, $\{a_{i,r},a_{j,r},a_{i,j}\}$ is the intersection of the planes $\operatorname{cl}(\{a_{1,r},a_{i,r},a_{j,r}\})$ and $\operatorname{cl}(L_{i}\cup J_{j})$, and so is a line. A similar argument shows that $\{a_{i,r+1},a_{j,r+1},a_{i,j}\}$ is a line. For (b), by (a) and the labeling, the plane $P_{r}=\operatorname{cl}(\{a_{i,j},a_{j,k},a_{j,r}\})$ contains both $a_{i,r}$ (which is in $\operatorname{cl}(\{a_{i,j},a_{j,r}\})$) and $a_{k,r}$ (which is in $\operatorname{cl}(\{a_{j,k},a_{j,r}\})$), and so $P_{r}$ contains $a_{i,k}$, which is in $\operatorname{cl}(\{a_{i,r},a_{k,r}\})$; likewise, the plane $P_{r+1}=\operatorname{cl}(\{a_{i,j},a_{j,k},a_{j,r+1}\})$ contains both $a_{i,r+1}$ and $a_{k,r+1}$, and so it contains $a_{i,k}$. Thus, $\{a_{i,j},a_{i,k},a_{j,k}\}$ is $P_{r}\cap P_{r+1}$ and so is a line. Finally, the r1r-1 large lines $L_{1},L_{2},\ldots,L_{r-1}$, the $2\binom{r-1}{2}$ other large lines in the planes $\operatorname{cl}(L_{i}\cup L_{j})$, and the $\binom{r-1}{3}$ large lines identified in (b) account for all $\binom{r+1}{3}$ lines of $M$. With this identification, we can now assume that $M$ and $\operatorname{Tr}^{k}(M(K_{r+1}))$ have the same ground set and the same $3$-point lines.
We next show that sets that are independent in $M$ are independent in $\operatorname{Tr}^{k}(M(K_{r+1}))$; equivalently, every circuit $C$ of $\operatorname{Tr}^{k}(M(K_{r+1}))$ is dependent in $M$. This holds if $C$ spans $\operatorname{Tr}^{k}(M(K_{r+1}))$ since then $|C|>r(M)$. It also holds if $|C|=3$ by our work above. Now assume that $4\leq|C|\leq r-k$. In $K_{r+1}$, the set $C$ contains the edges of a cycle; fix an edge $e$ in $C$, let $v$ run through the vertices in that cycle that are not incident with $e$, and consider the $3$-cycles determined by $e$ and $v$, which are large lines of both $M$ and $\operatorname{Tr}^{k}(M(K_{r+1}))$ that contain $e$. Each edge of $C$ is in either one of these lines or the closure of two of them, so, in both $M$ and $\operatorname{Tr}^{k}(M(K_{r+1}))$, the set $C$ is in the closure of the union of |C|2|C|-2 large lines that contain $e$. That union has rank |C|1|C|-1 in $M$, so $C$ is dependent in $M$.
By what we just showed, $M$ has at most as many independent sets as $\operatorname{Tr}^{k}(M(K_{r+1}))$; furthermore, if $M$ and $\operatorname{Tr}^{k}(M(K_{r+1}))$ have the same number of independent sets, then they have the same independent sets and so are the same matroid. ∎
6. Dowling geometries
We first recall the idea behind Dowling geometries. Let $\{b_{1},b_{2},\ldots,b_{r}\}$ be a basis of the projective geometry $\operatorname{PG}(r-1,q)$ over the finite field $\operatorname{GF}(q)$. Let $Q_{r}(\operatorname{GF}^{*}(q))$ be the restriction of $\operatorname{PG}(r-1,q)$ to the set $\bigcup_{\{i,j\}\subseteq[r]}\operatorname{cl}(\{b_{i},b_{j}\})$. The independent sets, circuits, flats, etc., in $Q_{r}(\operatorname{GF}^{*}(q))$ can be written using only the group $\operatorname{GF}^{*}(q)$ of units of $\operatorname{GF}(q)$. Dowling [18] showed how to replace $\operatorname{GF}^{*}(q)$ by an arbitrary finite group.
There are many ways to define Dowling geometries, including via group-labeled partial partitions (see [18]) and group-labeled graphs (see [24]). Cyclic flats play a key role when we show that Dowling geometries of the same rank based on groups of the same order give the same point in the polytope of unlabeled matroids (Theorem 6.4), so to define Dowling geometries, we give their cyclic flats and the rank of each cyclic flat. In a matroid $M$ with no loops, for each connected component $X$ of a nonempty cyclic flat, $|X|>1$ and $X$ is a connected flat, i.e., $X$ is a flat and $M|X$ is connected. We give the connected flats $X$ of Dowling geometries with $|X|>1$, and use them to give the cyclic flats.
Let $G$ be a finite group, written with multiplicative notation. Fix $r\geq 3$. The ground set of the rank-$r$ Dowling geometry $Q_{r}(G)$ over $G$ is
$\{b_{1},b_{2},\ldots,b_{r}\}\cup\{a_{(i,j)}\,:\,a\in G\text{ and }1\leq i<j\leq r\}.$
Hence $Q_{r}(G)$ has $r+|G|\binom{r}{2}$ elements, among which are $\binom{r}{2}$ copies of each element of $G$, each one indexed by an ordered pair $(i,j)$ with $1\leq i<j\leq r$. For $i>j$, let $a_{(i,j)}=(a^{-1})_{(j,i)}$, where $a^{-1}$ is the inverse of $a$ in $G$. The connected flats $F$ with $|F|>1$ are of two types:
  • (L)
    for $I\subseteq[r]$ with $|I|\geq 2$, the large flat
    $\{b_{i}\,:\,i\in I\}\cup\{a_{(i,j)}\,:\,a\in G\text{ and }\{i,j\}\subseteq I\}$
    has rank $|I|$ and support $I$;
  • (S)
    for a subset $J=\{j_{1}<j_{2}<\cdots<j_{t}\}$ of $[r]$ with $t\geq 3$ and (not necessarily distinct) $a_{2},a_{3},\ldots,a_{t}\in G$, the small flat
    $\{(a_{h})_{(j_{1},j_{h})}\,:\,2\leq h\leq t\}\cup\{(a_{h}^{-1}a_{i})_{(j_{h},j_{i})}\,:\,2\leq h<i\leq t\}$
    has rank t1t-1 and support $J$.
Thus, a large flat with support $I$ has size $|I|+|G|\binom{|I|}{2}$, and a small flat with support $J$ has size $\binom{|J|}{2}$. The cyclic flats of $Q_{r}(G)$ are the unions of at most one large flat and any number of small flats for which all supports are pairwise disjoint; the rank of a cyclic flat is the sum of the ranks of the large and small flats that it contains. Figure 3 illustrates the construction with $Q_{3}(G)$ where $G$ is the sign group, $\{1,-1\}$. In that example, the set $I=\{1,2\}$ gives the large line $\{b_{1},b_{2},1_{(1,2)},-1_{(1,2)}\}$, and the set $J=[3]$ and the group elements $a_{2}=a_{3}=-1$ give the dotted line $\{-1_{(1,2)},1_{(2,3)},-1_{(1,3)}\}$.
$1_{(1,2)}$$b_{1}$$1_{(1,2)}$$1_{(2,3)}\,\,\,\,$$b_{2}$    $1_{(2,3)}$$1_{(1,3)}$$b_{3}$$1_{(1,3)}$
Figure 3. The Dowling geometry $Q_{3}(G)$ where $G$ is the sign group, $\{1,-1\}$.
It is easy to check that small flats are the sets $F$ for which, for some $J\subseteq[r]$ with $|J|\geq 3$, (i) $F$ consists of one element from each of the $\binom{|J|}{2}$ sets $\{a_{(i,j)}\,:\,a\in G\}$ for $\{i,j\}\subsetneq J$, and (ii) if $a_{(i,j)},a^{\prime}_{(j,k)}\in F$, where $\{i,j,k\}\subseteq J$, then $(aa^{\prime})_{(i,k)}\in F$. Also, the restriction of $Q_{r}(G)$ to a small rank-$h$ flat is isomorphic to $M(K_{h+1})$ and the restriction to a large rank-$h$ flat is isomorphic to $Q_{h}(G)$. If $|G|=1$, then $Q_{r}(G)$ is isomorphic to $M(K_{r+1})$, which we treated in Section 5. The results in this section apply to Dowling geometries $Q_{r}(G)$ with $|G|>1$. Only when $|G|>1$ is the set $\{b_{1},b_{2},\ldots,b_{r}\}$, which is a basis of $Q_{r}(G)$, determined by the isomorphism type of $Q_{r}(G)$.
We can define $Q_{3}(G)$ when $G$ is a quasigroup (a Latin square), that is, $G$ has a binary operation, and the equations $ax=b$ and $ya=b$ have unique solutions for all $a,b\in G$. Quasigroups may lack an identity and inverses, so the identification $a_{(i,j)}=(a^{-1})_{(j,i)}$ mentioned above for groups makes no sense; thus, we write $a_{(i,j)}$ only when $i<j$. The small lines of $Q_{3}(G)$ are $\{a_{(1,2)},a^{\prime}_{(2,3)},(aa^{\prime})_{(1,3)}\}$ for $a,a^{\prime}\in G$. For $r=3$, our results apply to $Q_{3}(G)$ where $G$ is any quasigroup with $|G|>1$.
We will use the following characterization of Dowling geometries. The case of $r\geq 4$ was proven in [6]; the extension to $r=3$, along with a different proof, appeared in [4], which also gave a geometric criterion for distinguishing rank-$3$ Dowling geometries based on groups from those based on other quasigroups. That criterion cannot be cast in terms of valuative invariants since all Dowling geometries of the same rank based on groups (or, if $r=3$, quasigroups) of the same order have the same $\mathcal{G}$-invariant [9, Proposition 4.8].
Theorem 6.1.
A simple matroid $M$ of rank $r\geq 3$ is a Dowling geometry (based on a quasigroup if $r=3$) if and only if $M$ has elements $b_{1},b_{2},\dots,b_{r}$ for which
  1. (1)
    $E(M)$ is the union of the $\binom{r}{2}$ lines $\operatorname{cl}(\{b_{i},b_{j}\})$, for $\{i,j\}\subsetneq[r]$,
  2. (2)
    $|\operatorname{cl}(\{b_{i},b_{j}\})|>2$ for all $\{i,j\}\subsetneq[r]$, and
  3. (3)
    if $\{i,j,k\}\subseteq[r]$, $e\in\operatorname{cl}(\{b_{i},b_{j}\})-\{b_{i},b_{j}\}$, and $f\in\operatorname{cl}(\{b_{i},b_{k}\})-\{b_{i},b_{k}\}$, then $|\operatorname{cl}(\{e,f\})|>2$.
The next result characterizes Dowling geometries of rank $r\geq 3$ other than $M(K_{r+1})$ by optimizing a sequence of valuative invariants. Since all Dowling geometries of rank $r$ based on groups (or quasigroups if $r=3$) of order $g$ have the same $\mathcal{G}$-invariant, we cannot distinguish between them via valuative invariants, but they are extremal since, as we show in Theorem 6.4 below, they give the same point in the polytope of unlabeled matroids.
Theorem 6.2.
Fix integers $r\geq 3$ and $g\geq 2$. Let $M$ be a simple matroid of rank $r$ with $|E(M)|=r+g\binom{r}{2}$ in which, for all $h\in[r-1]$, each rank-$h$ flat has at most $h+g\binom{h}{2}$ elements. Call a rank-$h$ flat $F$ large if $|F|=h+g\binom{h}{2}$. Assume that for any rank-$h$ flat $F$ that is not large, if $2<h<r$, then $|F|\leq h+g\binom{h-1}{2}$, and if $h=2$, then $|F|\leq 3$. Then the following conclusions hold.
  • (a)
    The matroid $M$ has at most $r$ large hyperplanes.
  • (b)
    If $M$ has $r$ large hyperplanes, then $M$ has a basis $B$ for which $\operatorname{cl}(X)$ is large for all $X\subseteq B$.
  • (c)
    If $M$ has $r$ large hyperplanes, then the greatest number of $3$-point lines in $M$ is $\binom{r}{3}g^{2}$, and any such matroid $M$ with $\binom{r}{3}g^{2}$ three-point lines is a rank-$r$ Dowling geometry based on a group (or quasigroup if $r=3$) of order $g$.
Proof.
We first prove (a) by induction on $r$. In addition to the base case $r=3$, we treat $r=4$ separately since an inequality that we use for the induction step applies only for $r>4$. Let $H_{1},H_{2},\ldots,H_{t}$ be the large hyperplanes of $M$.
First consider $r=3$. If $t\geq 3$ and $H_{i}\cap H_{j}=\emptyset$ for some $\{i,j\}\subsetneq[3]$, then we would get the contradiction $|H_{1}\cup H_{2}\cup H_{3}|\geq 2(2+g)+g=4+3g>|E(M)|$. The same contradiction (with different counting) would occur if $H_{1}\cap H_{2}\cap H_{3}\neq\emptyset$. Thus, $H_{i}\cap H_{j}$ is a singleton for all $\{i,j\}\subsetneq[3]$, while $H_{1}\cap H_{2}\cap H_{3}=\emptyset$. Inclusion/exclusion now gives $|H_{1}\cup H_{2}\cup H_{3}|=3+3g$, so $H_{1}\cup H_{2}\cup H_{3}=E(M)$. Thus, any large line contains at least two points in one of $H_{1},H_{2},H_{3}$ and so is that large line, so $t=3$. Also, the union of the three singletons $H_{i}\cap H_{j}$, for $\{i,j\}\subsetneq[3]$, is a basis, and the line spanned by any two of its elements is large. Thus, (a) and (b) hold when $r=3$.
Now consider $r=4$. Focus on two large planes $H_{1}$ and $H_{2}$, and let $P$ be any other plane. Since $|H_{1}|+|H_{2}|=6+6g=|E(M)|+2$, we have $|H_{1}\cap H_{2}|\geq 2$, so $|H_{1}\cap H_{2}|$ is $2$, 33, or $2+g$ by the assumption on the sizes of lines. First assume that $|H_{1}\cap H_{2}|=2$, so $H_{1}\cup H_{2}=E(M)$. Thus, $P=(P\cap H_{1})\cup(P\cap H_{2})$, so $|P|\leq 2(2+g)<3+3g$. Thus, $H_{1}$ and $H_{2}$ are the only large planes.
Now assume that $|H_{1}\cap H_{2}|=3$. Since $P\subseteq(P\cap H_{1})\cup(P\cap H_{2})\cup(E(M)-(H_{1}\cup H_{2}))$ and $|E(M)-(H_{1}\cup H_{2})|=1$, we get $|P|\leq 2(2+g)+1$. If $g>2$, then $5+2g<3+3g$, and so $H_{1}$ and $H_{2}$ are the only large planes. Now assume that $g=2$, so large planes have nine elements and lines have at most four elements. If $|P|=9$, then $P\cap H_{1}$ and $P\cap H_{2}$ are disjoint large lines, and so are disjoint from $H_{1}\cap H_{2}$. Since $|H_{1}\cap H_{2}|=3$, at most one large line of $H_{1}$ is disjoint from $H_{1}\cap H_{2}$, and likewise for $H_{2}$, so there is at most one large plane besides $H_{1}$ and $H_{2}$. Thus, there are at most three large planes.
We may now assume that any two large planes of $M$ intersect in a large line. No large line is in three large planes, for otherwise the union of the three large planes would contain $2+g+3(1+2g)=5+7g$ elements, but $|E(M)|=4+6g$. Since any large plane contains at most three large lines, each of which is in at most one more large plane, there are at most four large planes, so (a) holds when $r=4$.
Assume that $M$ has four large planes, $H_{i}$ for $i\in[4]$. By the argument above, $H_{i}\cap H_{j}$ is a large line for all $\{i,j\}\subsetneq[4]$ and no two such intersections are equal, so each $H_{i}$ contains three large lines. Using what we showed in the case $r=3$, it follows that $H_{i}\cap H_{j}\cap H_{k}$ is a singleton for all $\{i,j,k\}\subsetneq[4]$, all four such intersections are different, and their union is a basis, so (b) holds when $r=4$.
Now assume that $r>4$ and that the bound on the number of large hyperplanes holds for lower ranks. We claim that, if $\{i,j\}\subseteq[t]$, then $H_{i}\cap H_{j}$ is a large rank-(r2)(r-2) flat. Since $|H_{i}\cup H_{j}|\leq r+g\binom{r}{2}$, we have
$|H_{i}\cap H_{j}|$ $\,=|H_{i}|+|H_{j}|-|H_{i}\cup H_{j}|$
$\,\geq 2\Bigl{(}r-1+g\binom{r-1}{2}\Bigr{)}-\Bigr{(}r+g\binom{r}{2}\Bigl{)}$
$\,=r-2+g\Bigl{(}2\binom{r-1}{2}-\binom{r}{2}\Bigl{)}$
$\,=r-2+g\Bigl{(}\binom{r-2}{2}-1\Bigl{)}$
$\,>r-2+g\binom{r-3}{2}$,
(the final inequality requires $r>4$). By the hypotheses, $|F|\leq r-3+g\binom{r-3}{2}$ for a flat $F$ if r(F)<r2r(F)<r-2, and $|F|\leq r-2+g\binom{r-3}{2}$ if r(F)=r2r(F)=r-2 but $F$ is not large, so $H_{i}\cap H_{j}$ is a large rank-(r2)(r-2) flat, as claimed.
Each of $H_{i}-(H_{i}\cap H_{j})$ and $H_{j}-(H_{i}\cap H_{j})$ contains
$r-1+g\binom{r-1}{2}-(r-2)-g\binom{r-2}{2}=1+g(r-2)$
out of the
$r+g\binom{r}{2}-(r-2)-g\binom{r-2}{2}=2+g(2r-3)$
elements in $E(M)-(H_{i}\cap H_{j})$, so $H_{i}$ and $H_{j}$ are the only large hyperplanes that contain $H_{i}\cap H_{j}$. Therefore all $\binom{t}{2}$ intersections $H_{i}\cap H_{j}$, with $\{i,j\}\subseteq[t]$, are different. In particular, $H_{i}\cap H_{t}$, for $i\in[t-1]$, are t1t-1 different large hyperplanes of $M|H_{t}$, so the induction assumption applied to $M|H_{t}$ gives $t-1\leq r-1$, so $t\leq r$, so (a) holds.
For the induction step in the proof that a basis $B=\{b_{1},b_{2},\ldots,b_{r}\}$ satisfying statement (b) exists, it will be useful to know some properties of such bases. Counting shows that $\bigcup_{\{i,j\}\subsetneq[r]}\operatorname{cl}(\{b_{i},b_{j}\})=E(M)$. We claim that if $L$ is a line and $|L|\geq 3$, then we have $L\subseteq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$ for some $\{i,j,k\}\subseteq[r]$. Fix $\{e_{1},e_{2},e_{3}\}\subseteq L$. There are $2$-subsets $B_{1},B_{2},B_{3}$ of $B$ for which $e_{i}\in\operatorname{cl}(B_{i})$ for each $i\in[3]$. Now $e_{3}\in\operatorname{cl}(B_{1}\cup B_{2})\cap\operatorname{cl}(B_{3})$, so $r(B_{1}\cup B_{2}\cup B_{3})\leq 5$, so $B_{1}$, $B_{2}$, and $B_{3}$ are not pairwise disjoint, so we may assume that $B_{1}\cap B_{2}\neq\emptyset$. Thus, $L$ is contained in $\operatorname{cl}(B_{1}\cup B_{2})$, which is either a line $\operatorname{cl}(\{b_{i},b_{j}\})$ or a plane $\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$, thus proving the claim. It follows that the lines $\operatorname{cl}(\{b_{i},b_{j}\})$ are the only large lines of $M$ since, for any large line $L$ of $M$, we have $L\subseteq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$ for some $\{i,j,k\}\subseteq[r]$, and the large plane $\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$ has three large lines, $\operatorname{cl}(\{b_{i},b_{j}\})$, $\operatorname{cl}(\{b_{i},b_{k}\})$, and $\operatorname{cl}(\{b_{j},b_{k}\})$, as shown in the second paragraph of the proof. So, when a basis $B=\{b_{1},b_{2},\ldots,b_{r}\}$ as in (b) exists, it is unique.
We can now prove that for any line $L$ with $b_{i}\in L$, either $|L|=2$ or $L=\operatorname{cl}(\{b_{i},b_{j}\})$ for some $j\in[r]-\{i\}$. This is because if $|L|\neq 2$, then $L\subsetneq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$ for some $\{j,k\}\subseteq[r]-\{i\}$, and $|L\cap\operatorname{cl}(\{b_{j},b_{k}\})|\leq 1$, so $L$ has a point in one of $\operatorname{cl}(\{b_{i},b_{j}\})$ or $\operatorname{cl}(\{b_{i},b_{k}\})$ in addition to $b_{i}$, and so $L$ is that line.
We now prove (b) by induction on $r$. This was shown above for $r\in\{3,4\}$. For the induction step, assume that $r>4$ and that (b) holds for ranks less than $r$. Thus, $M|H_{1}$ has a basis $\{b_{1},b_{2},\ldots,b_{r-1}\}$ that satisfies (b). Now $M|H_{2}$ and $M|(H_{1}\cap H_{2})$ also have such bases. By the uniqueness of the basis of $M|(H_{1}\cap H_{2})$ satisfying (b) and noting that the large lines of $M|(H_{1}\cap H_{2})$ are large lines of $M|H_{1}$ and $M|H_{2}$ as well, it follows that, up to relabeling, the basis of $M|H_{2}$ that satisfies (b) is $\{b_{2},b_{3},\ldots,b_{r-1},b_{r}\}$ for some element $b_{r}$. We claim that the basis of $M$ that satisfies (b) is $B=\{b_{1},b_{2},\ldots,b_{r}\}$. To prove this, it suffices to show that the line $\operatorname{cl}(\{b_{1},b_{r}\})$ is large since any other line spanned by two elements of $B$ is in at least one of $M|H_{1}$ and $M|H_{2}$ and so is large. Consider the hyperplane $H=\operatorname{cl}(B-\{b_{2}\})$. We know that the lines $\operatorname{cl}(\{b_{i},b_{j}\})$, for $\{i,j\}\subsetneq[r]-\{2\}$ and $\{i,j\}\neq\{1,r\}$, are large since each is a large line of at least one of $M|H_{1}$ and $M|H_{2}$; thus, $|H|\geq r-1+g\bigl{(}\binom{r-1}{2}-1\bigl{)}$, so $H$ is a large hyperplane. By the induction hypothesis, $M|H$ has a basis that satisfies (b), and, as we showed above, that basis accounts for all large lines of $M|H$. Thus, that basis must be $B-\{b_{2}\}$, and so the line $\operatorname{cl}(\{b_{1},b_{r}\})$ is large, as needed. Thus, (b) holds.
We now turn to (c). As shown above, for the basis $B=\{b_{1},b_{2},\ldots,b_{r}\}$ in (b) and any $3$-point line $L$ of $M$, there is a subset $\{i,j,k\}$ of $[r]$ for which $L\subseteq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$; also, $L\cap B=\emptyset$. The bound $\binom{r}{3}g^{2}$ follows since $\binom{r}{3}$ is the number of large planes $\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$, and if $L\subseteq\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$, then $L$ is determined by its intersections with any two of the large lines in $\operatorname{cl}(\{b_{i},b_{j},b_{k}\})$, each of which has $g$ elements that are not in $B$.
Now assume that $M$ has $\binom{r}{3}g^{2}$ three-point lines. Properties (1) and (2) of Theorem 6.1 follow from (b). Since $M$ has $\binom{r}{3}g^{2}$ three-point lines, each choice of a subset $\{i,j,k\}\subseteq[r]$ and elements $e\in\operatorname{cl}(\{b_{i},b_{j}\})-\{b_{i},b_{j}\}$ and $f\in\operatorname{cl}(\{b_{i},b_{k}\})-\{b_{i},b_{k}\}$ must give a $3$-point line $\operatorname{cl}(\{e,f\})$. Thus, property (3) of Theorem 6.1 also holds, so, as claimed, $M$ is a Dowling geometry. ∎
The theorem and proof above easily yield the following alternative characterization of Dowling geometries by optimizing valuative invariants.
Theorem 6.3.
Fix integers $r\geq 3$ and $g\geq 2$. Let $\mathcal{C}$ be the set of all simple rank-$r$ matroids on $r+g\binom{r}{2}$ elements in which lines have $2$, 33, or $g+2$ points and, for each flat $F$ of rank $h$ with $2<h<r$, either $|F|=h+g\binom{h}{2}$ or $|F|\leq h+g\binom{h-1}{2}$. Call rank-$i$ flats of size $i+g\binom{i}{2}$ large, and call a flag large if, for each $i\in[r]\cup\{0\}$, it has a large rank-$i$ flat. Dowling geometries (including those over quasigroups if $r=3$) are the matroids in $\mathcal{C}$ that maximize the number of large flags and, after that, maximize the number of $3$-point lines.
Proof.
Dowling geometries in $\mathcal{C}$ have r!(g+2)/2r!\,(g+2)/2 large flags since a flag includes one of the $r$ large hyperplanes, and one of the r1r-1 large rank-(r2)(r-2) flats that it contains, and so on through one of the three large lines in a large plane, and then one of the $g+2$ points in that large line. Furthermore, since, by Theorem 6.2, matroids in $\mathcal{C}$ have at most $r$ large hyperplanes, each of which contains at most r1r-1 large rank-(r2)(r-2) flats, etc., if a matroid in $\mathcal{C}$ has r!(g+2)/2r!\,(g+2)/2 large flags, then it must have $r$ large hyperplanes. From that, part (c) of Theorem 6.2 gives the result. ∎
The next result completes the proof that Dowling geometries are extremal.
Theorem 6.4.
Fix integers $r\geq 3$ and $g\geq 2$. All Dowling geometries of rank $r$ based on groups (or quasigroups if $r=3$) of order $g$ yield the same point in the polytope $\Omega_{r,r+g\binom{r}{2}}$ of unlabeled matroids. Thus, Dowling geometries are extremal.
Proof.
Let the group (or, if $r=3$, quasigroup) $G$ have order $g$. To prove the result, we show that the point in the polytope $\Omega_{r,r+g\binom{r}{2}}$ given by $Q_{r}(G)$ is determined by $r$ and $g$.
Recall (L) and (S) from earlier in this section. Call a connected component of a cyclic flat $F\in\mathcal{Z}(Q_{r}(G))$ large or small according to whether it is a large or small flat. At most one connected component of a cyclic flat is large. Let $[F]$, the type of $F$, be $(T,U)$ where $T$ is the set of supports of the small components of $F$, and if $F$ has a large component, then $U$ is its support, otherwise $U=\emptyset$. The type $[C]$ of a chain $C\in\mathcal{C}(\mathcal{Z}(Q_{r}(G)))$ is the set of the types of its members. The type $[F]=(T,U)$ gives $r(F)$ and $|F|$:
$r(F)=|U|+\sum_{X\in T}(|X|-1)\quad\text{ and }\quad|F|=|U|+g\,\binom{|U|}{2}+\sum_{X\in T}\binom{|X|}{2}$.
Thus, from the type $[C]$ of $C\in\mathcal{C}(\mathcal{Z}(Q_{r}(G)))$, we get the isomorphism type of the nested matroid that $C$ contributes to Equation (2.1) and its symmetrized counterpart.
By (L) and (S), for each type $(T,U)$, the number of cyclic flats $F$ of $Q_{r}(G)$ of type $(T,U)$ depends only on $g$ and $(T,U)$: it is $g^{t}$ where $t=\sum_{X\in T}(|X|-1)$ since $F$ is determined by $(T,U)$ along with, for each $X\in T$, say with least element $h$, the choice of one of $g$ elements $a_{(h,k)}\in\operatorname{cl}(\{b_{h},b_{k}\})-\{b_{h},b_{k}\}$ for each $k\in X-\{h\}$.
For $F,F^{\prime}\in\mathcal{Z}(Q_{r}(G))$ with $[F]=(T,U)$ and $F^{\prime}]=(T^{\prime},U^{\prime})$, if $F\subseteq F^{\prime}$, then (i) $U\subseteq U^{\prime}$ and (ii) for each $X\in T$, either $X\subseteq U^{\prime}$ or $X\subseteq X^{\prime}$ for some $X^{\prime}\in T^{\prime}$. We claim that, given $F\in\mathcal{Z}(Q_{r}(G))$ with $[F]=(T,U)$ and $(T^{\prime},U^{\prime})$ for which properties (i) and (ii) hold, the number of $F^{\prime}\in\mathcal{Z}(Q_{r}(G))$ with $F^{\prime}]=(T^{\prime},U^{\prime})$ and $F\subseteq F^{\prime}$ depends just on $(T,U)$, $(T^{\prime},U^{\prime})$, and $g$; specifically, it is $g^{t}$ where $t$ is the sum of the terms that we identify next. An element in $U^{\prime}-U$ is either in no set in $T$ or, if it is in set $X$ of $T$, then $X\subseteq U^{\prime}-U$; in both cases, by (L), no choice of group elements is made, so such elements contribute nothing to $t$. Some sets $X^{\prime}$ of $T^{\prime}$ may be disjoint from all sets in $T$; such a set contributes a small component to $F^{\prime}$, and $g^{|X^{\prime}|-1}$ small flats have support $X^{\prime}$ by (S), so such a set contributes $|X^{\prime}|-1$ to the sum that defines $t$. Sets that are common to $T$ and $T^{\prime}$ contribute nothing. All other sets in $T^{\prime}$ are formed by merging pairs of sets (i.e., replacing two sets by their union) and adjoining an element that is in neither $U$ nor a set in $T$ to a set; both merging and adjoining can be applied multiple times; it follows from (S) that each merger of two sets or adjoining of one element to a set requires choosing one group element and so contributes $1$ to the sum that defines $t$.
[…]
7. Spikes
Spikes have played many important roles in matroid theory in recent decades (see [24]). The following generalization of spikes was introduced in [7]: for $r,s,t\in\mathbb{N}$ with $r\geq 3$, $s\geq r-1$, and $t\geq 2$, an $(r,s,t+1)$-spike with tip $a$ is a rank-$r$ simple matroid whose ground set is the union of $s$ lines $L_{1},L_{2},\ldots,L_{s}$, each having $t+1$ elements and containing $a$, for which, for any $k\in[r-1]$, the union of any $k$ of these lines has rank $k+1$. An $r$-spike is an $(r,r,3)$-spike. For example, $\operatorname{PG}(2,2)$ is a $3$-spike and (atypically) any element can be the tip; for a given point $a$ of $\operatorname{PG}(2,2)$, relaxing any one or more of the four circuit-hyperplanes that do not contain $a$ gives more examples of $3$-spikes. Also, there is only one (r,r1,t+1)(r,r-1,t+1)-spike up to isomorphism, namely, the parallel connection at $a$ of r1r-1 lines, each having $t+1$ elements.
In [7], $(r,s,t+1)$-spikes with $r\geq 5$ were shown to be distinguishable from all other matroids (but not necessarily from each other) by their Tutte polynomials, and it was shown which $(r,s,t+1)$-spikes have the same Tutte polynomial; those results imply that certain $(r,s,t+1)$-spikes are Tutte unique. We identify some $(r,s,t+1)$-spikes that are extremal. (We treat only spikes with tips. Peter Nelson (private communications) has treated $r$-spikes without tips.)
In any $(r,s,t+1)$-spike $M$ with $(t+1)$-point lines $L_{1},L_{2},\ldots,L_{s}$ that contain $a$, any $3$-subset of any $L_{i}$ is a circuit, as is any $4$-set $C$ for which, for some $\{i,j\}\subseteq[s]$, we have $|C\cap(L_{i}-\{a\})|=2=|C\cap(L_{j}-\{a\})|$. All other circuits of $M$ have size $r$ or $r+1$. An $r$-set $X$ for which $|X\cap(L_{i}-\{a\})|=1$ for $r$ integers $i\in[s]$ is either a basis or a circuit; thus, there can be two types of $r$-circuits if $r\in\{3,4\}$. The hyperplane that is spanned by an $r$-circuit $X$ of the type just identified intersects each set $L_{i}-\{a\}$ in at most one element, and if $s=r$, then $X$ is a circuit-hyperplane.
For any $s$, $t$, and $r>4$, there is a unique $(r,s,t+1)$-spike, up to isomorphism, that has no $r$-circuits; this is the free $(r,s,t+1)$-spike. Similarly, in the free $(3,s,t+1)$-spike, the only $3$-circuits are the $3$-subsets of the lines $L_{i}$, and in the free $(4,s,t+1)$-spike, the only $4$-circuits are those that span a plane $L_{i}\cup L_{j}$. Equivalently, the free $(r,s,t+1)$-spike is the truncation to rank $r$ of the parallel connection of $s$ lines at $a$, each having $t+1$ elements. The free $(r,s,t+1)$-spike with $r\geq 4$ is Tutte unique (see [7]).
Among $3$-spikes, $\operatorname{PG}(2,2)$ maximizes the number of circuit-hyperplanes. For $r>3$, let $C$ be an $r$-circuit in $\operatorname{PG}(r-1,2)$, let $a$ be an element of $\operatorname{PG}(r-1,2)$ that is not in the hyperplane $\operatorname{cl}(C)$, and let $X$ be the union of the lines $L_{e}=\operatorname{cl}(\{a,e\})$ for $e\in C$. The restriction $M=\operatorname{PG}(r-1,2)|X$ is an $r$-spike, and $C$ is a circuit-hyperplane of $M$. It is easy to show that the symmetric difference of $C$ with $k$ of the sets $L_{e}-\{a\}$, for $e\in C$, is a circuit-hyperplane if and only if $k$ is even, otherwise the symmetric difference is a basis of $M$. Thus, $M$ has $2^{r-1}$ circuit-hyperplanes. Also, for $r>3$, the maximum number of circuit-hyperplanes in an $r$-spike is $2^{r-1}$, and using the ideas about symmetric differences above, it is not hard to construct an isomorphism between $M$ and any $r$-spike with $2^{r-1}$ circuit-hyperplanes. This justifies calling the $r$-spike with $2^{r-1}$ circuit-hyperplanes, for $r>3$, the binary $r$-spike. Of course, $\operatorname{PG}(2,2)$ is the binary $3$-spike. The binary $r$-spike with $r\geq 5$ is Tutte unique (see [7]).
The next result shows how to identify $(r,s,t+1)$-spikes, where $r\geq 5$, by optimizing valuative invariants, and we show that the free $(r,s,t+1)$-spike, the (r,r1,t+1)(r,r-1,t+1)-spike, and the binary $r$-spike are extremal.
Theorem 7.1.
Fix integers $r\geq 5$, $t\geq 2$, and $s\geq r-1$. Let $M$ be a rank-$r$ matroid on $st+1$ elements. Assume that for each $i\in[r-1]$, each rank-$i$ flat has at most (i1)t+1(i-1)t+1 elements. For $i\in[r-1]$, call rank-$i$ flats with (i1)t+1(i-1)t+1 elements large. Assume that $M$ minimizes the valuative invariant $N\mapsto F_{2,3}(N;t+1,w)$ for all $w\neq 2t+1$, so each plane that contains a large line is large. Then $M$ has at most $s$ large lines, and if $M$ has $s$ large lines, then $M$ is an $(r,s,t+1)$-spike. Also, the free $(r,s,t+1)$-spike, the binary $r$-spike, and the (r,r1,t+1)(r,r-1,t+1)-spike are extremal.
Proof.
By the sizes of large lines and planes, coplanar large lines intersect in a point. Let $L$ be any large line of $M$. Any plane $P$ with $L\subsetneq P$ is large, so |PL|=t|P-L|=t. Now |E(M)L|=(s1)t|E(M)-L|=(s-1)t and $\pi=\{P-L\,:\,P\text{ is a plane, }L\subsetneq P\}$ is a partition of E(M)LE(M)-L, so $\pi$ has s1s-1 blocks; say $\pi=\{B_{i}\,:\,i\in[s-1]\}$. The union of two planes $L\cup B_{i}$ and $L\cup B_{j}$ has rank $4$ and size $3t+1$, so $L\cup B_{i}\cup B_{j}$ is a flat. Let $L^{\prime}$ be a line of $M$ for which $r(L\cup L^{\prime})=4$. Thus, $\operatorname{cl}(L\cup L^{\prime})=L\cup B_{i}\cup B_{j}$ for some $\{i,j\}\subseteq[s-1]$, and $|L^{\prime}\cap B_{i}|=1=|L^{\prime}\cap B_{j}|$, so $|L^{\prime}|=2$. Thus, any large line of $M$ other than $L$ is $B_{i}\cup\{a\}$ for some $a\in L$ and $i\in[s-1]$, so there are at most $s$ large lines.
Since $L$ is any large line, the argument above shows that any two large lines of $M$ are coplanar. All large lines $L^{\prime}\neq L$ contain the same point of $L$ since two such large lines $L^{\prime}$ and $L^{\prime\prime}$ intersect in some point, which is therefore in the planes $L\cup L^{\prime}$ and $L\cup L^{\prime\prime}$ and so in their intersection, $L$. Thus, if $M$ has $s$ large lines, then $M$ is an $(r,s,t+1)$-spike.
The free $(r,s,t+1)$-spike is extremal since, among $(r,s,t+1)$-spikes, it is the only one that minimizes the number of $r$-circuits. The binary $r$-spike is extremal since, among $r$-spikes, it is the only one that maximizes the number of circuit-hyperplanes. Finally, the (r,r1,t+1)(r,r-1,t+1)-spike is extremal since it is the only spike with its parameters. ∎
We next give another characterization of (r,r1,t+1)(r,r-1,t+1)-spikes so that in the next section we can show that direct sums of such spikes, where $t$ is fixed, are extremal. The notion of an (r,r1,t+1)(r,r-1,t+1)-spike makes sense even when $r=2$, in which case a $(2,1,t+1)$-spike is a $(t+1)$-point line, $U_{2,t+1}$. What we gain by this mild extension is a slightly stronger result about direct sums in the next section. While Theorem 7.1 is limited to ranks $5$ and higher, Theorem 7.2 applies for ranks $2$ and higher.
Theorem 7.2.
Fix integers $r,t\in\mathbb{N}-\{1\}$. Let $M$ be a rank-$r$ matroid on (r1)t+1(r-1)t+1 elements. Assume that for each $i\in[r]$, each rank-$i$ flat has at most (i1)t+1(i-1)t+1 elements. For $i\in[r]$, call rank-$i$ flats with (i1)t+1(i-1)t+1 elements large, and call a flag of flats of $M$ large if it is counted by $F_{0,r}(N;0,1,t+1,2t+1,\ldots,(r-1)t+1)$. Then $M$ has at most (r1)!(t+1)(r-1)!(t+1) large flags, and $M$ has (r1)!(t+1)(r-1)!(t+1) large flags if and only if $M$ is an (r,r1,t+1)(r,r-1,t+1)-spike. Also, if $r>2$, then $M$ has (r1)!(t+1)(r-1)!(t+1) large flags if and only if $M$ has r1r-1 large hyperplanes, and that is the maximum number of large hyperplanes under the assumptions on $M$.
Proof.
The statement is easy to see if $r=2$, so we focus on $r\geq 3$. We first treat the bound on the number of large hyperplanes. A rank-$3$ matroid on $2t+1$ points has at most two large lines, so the bound holds for $r=3$. Now assume that $r>3$ and that the bound holds for smaller ranks. Let $H_{1},H_{2},\ldots,H_{k}$ be the large hyperplanes of $M$. For $\{i,j\}\subseteq[k]$, we have $H_{j}-H_{i}\subseteq E(M)-H_{i}$ and $|E(M)-H_{i}|=t\leq|H_{j}-(H_{i}\cap H_{j})|$, so $E(M)-H_{i}=H_{j}-H_{i}$; likewise, $E(M)-H_{j}=H_{i}-H_{j}$. Thus, $H_{i}\cup H_{j}=E(M)$ and $H_{i}\cap H_{j}$ is a large rank-(r2)(r-2) flat. Also, all intersections $H_{i}\cap H_{j}$, for $\{i,j\}\subseteq[k]$, are different since $H_{i}$ and $H_{j}$ are the only hyperplanes that contain $H_{i}\cap H_{j}$. The flats $H_{i}\cap H_{k}$, for $i\in[k-1]$, are large hyperplanes of $M|H_{k}$, so $k-1\leq r-2$ by the induction hypothesis, so $k\leq r-1$, as needed.
A routine induction now shows that the intersection of any $i$ large hyperplanes is a large rank-(ri)(r-i) flat. Thus, all such intersections are different. Also, if k=r1k=r-1, then $\bigcap_{i\in[r-1]}H_{i}$ is a rank-$1$ flat, say $\{a\}$, and $M$ has $\binom{r-1}{r-2}=r-1$ large lines, all of which contain $a$, so $M$ is an (r,r1,t+1)(r,r-1,t+1)-spike, and $M$ has (r1)!(t+1)(r-1)!(t+1) large flags.
Arguments like those given in earlier sections show that, when $r>2$, having r1r-1 large hyperplanes is equivalent to having (r1)!(t+1)(r-1)!(t+1) large flags. ∎
8. Direct sums of some extremal matroids
Ferroni and Fink [20] conjectured that direct sums of extremal matroids are extremal. While it is far from settling the conjecture, in Theorem 8.2 below we show that direct sums of extremal matroids are extremal if they share certain features that can be described by optimizing valuative invariants. The relevant common features occur in, for instance, (a) cycle matroids of complete graphs, (b) projective geometries of rank at least four and the same order, (c) affine geometries of rank at least four and the same order, (d) Dowling geometries based on a group that is the only group of its order, and (e) (r,r1,t+1)(r,r-1,t+1)-spikes where $t$ is fixed.
We start with another special case of the conjecture; we treat direct sums of uniform matroids. Ferroni and Fink [20, Example 4.4] observed that uniform matroids are extremal since they maximize the number of bases; alternatively, Theorem 3.1 applies since uniform matroids are perfect matroid designs. The proof of the next result uses the lexicographic order on $(n,r)$-sequences: for two $(n,r)$-sequences $\underline{x}=x_{1}\ldots x_{n}$ and $\underline{y}=y_{1}\ldots y_{n}$, we have $\underline{x}<\underline{y}$ if, for the least $i$ with $x_{i}\neq y_{i}$, we have $x_{i}<y_{i}$, so $x_{i}=0$ and $y_{i}=1$.
Theorem 8.1.
Direct sums of uniform matroids are extremal.
Proof.
Let $M$ be $U_{r_{1},n_{1}}\oplus U_{r_{2},n_{2}}\oplus\cdots\oplus U_{r_{k},n_{k}}$ where $r_{1}\leq r_{2}\leq\cdots\leq r_{k}$ and, if $r_{i}=r_{i+1}$, then $n_{i}\geq n_{i+1}$. Set $r=r(M)$ and $n=|E(M)|$. If $M$ has $s$ loops and $t$ coloops, then we can focus on matroids that have $s$ loops and $t$ coloops by maximizing the valuative invariant $N\mapsto f_{0}(N;s,0)+f_{r}(N;n,t)$. Thus, without loss of generality, we can assume that $s=t=0$, so $0<r_{i}<n_{i}$ for all $i\in[k]$. The sum of the coefficients of the terms $x^{i}y^{j}$, for $i+j<k$, in the Tutte polynomial of a matroid is a valuative invariant, and by minimizing it, we can focus on matroids that have at least $k$ connected components. The number of cyclic flats in a matroid is a valuative invariant, and the matroids that minimize it, within the class of matroids that have at least $k$ connected components and neither loops nor coloops, have exactly $k$ connected components and their $2^{k}$ cyclic flats are the unions of none, some, or all of the connected components; therefore the restriction to such a connected component $X$ has just two cyclic flats, $\emptyset$ and $X$, so such restrictions are uniform. Let $\underline{x}=x_{1}\ldots x_{n}$ be the $(n,r)$-sequence
[…]
Note that $\underline{x}$ encodes the rank and size of each connected component of $M$, and it is the lexicographically least rank sequence $\underline{r}$ for which $g_{\underline{r}}(M)$, the coefficient of $\underline{r}]$ in $\mathcal{G}(M)$, is nonzero. For each $(n,r)$-sequence $\underline{y}$ with $\underline{y}<\underline{x}$ in lexicographic order, minimize the valuative invariant $N\mapsto g_{\underline{y}}(N)$. Then maximize the valuative invariant $N\mapsto g_{\underline{x}}(N)$. Let $N$ be a matroid that emerges from this sequence of optimization problems, so $N$ is an $n$-element rank-$r$ direct sum of $k$ uniform matroids for which $\underline{x}$ is the lexicographically least rank sequence $\underline{r}$ for which $g_{\underline{r}}(N)\neq 0$. Thus, the components of $N$ have the same ranks and size as those of $M$, so $N$ is isomorphic to $M$. Thus, $M$ is extremal. ∎
We next give a result on direct sums of extremal matroids that applies to direct sums of many of the matroids that we treated earlier in this paper. Looking ahead to Corollary 8.3 may make the notation below easier to digest. Characterizations of the type used in this result are illustrated by Theorem 5.3 for $M(K_{r+1})$. In many applications, $\mathcal{M}$ contains, up to isomorphism, one matroid of each rank (e.g,. $\mathcal{M}$ can be the set of cycle matroids of complete graphs of rank two and greater). For the general case, an example to keep in mind is Dowling geometries, where many matroids maximize the number of large flags, but only Dowling geometries also maximize the number of $3$-point lines.
Theorem 8.2.
Let $\alpha_{0}=0,\alpha_{1}=1,\alpha_{2},\alpha_{3},\ldots$ be integers for which
  • (1)
    $\alpha_{2}\geq 3$ and $\alpha_{i}+\alpha_{j}\leq\alpha_{i+j-1}+1$ for all $i,j\in\mathbb{N}$.
Fix a set $A\subseteq[\alpha_{2}]-\{1\}$ with $2,\alpha_{2}\in A$. Fix an integer $d\geq 2$. Let $\mathcal{M}$ be a set of simple matroids of rank at least $d$ that satisfies the following conditions:
  • (2)
    for all $M\in\mathcal{M}$ and $i\leq r(M)$, some rank-$i$ flat of $M$ has size $\alpha_{i}$,
  • (3)
    for each $M\in\mathcal{M}$ and flat $F$ of $M$, if $2<r(F)<r(M)$, then either $|F|=\alpha_{r(F)}$ or $|F|\leq\alpha_{r(F)-1}+1$, while if $r(F)=2$, then $|F|\in A$, and
  • (4)
    for $r\geq d$, a rank-$r$ simple matroid $M$ is in $\mathcal{M}$ if and only if, among rank-$r$ simple matroids that satisfy conditions (2)-(3), the valuative invariant that maps $N$ to $F_{0,r}(N;\alpha_{0},\alpha_{1},\alpha_{2},\ldots,\alpha_{r})$ is maximized by $M$.
If, for some sequence of valuative invariants $f_{1},f_{2},\ldots,f_{k}$ that are additive on direct sums (i.e., $f_{i}(M\oplus N)=f_{i}(M)+f_{i}(N)$), the matroids $M_{1},M_{2},\ldots,M_{t}$ are, up to isomorphism, the only matroids in $\mathcal{M}$ of their ranks that maximize $f_{1}$, and then $f_{2}$, and so on, through $f_{k}$, then $M_{1}\oplus M_{2}\oplus\cdots\oplus M_{t}$ is extremal.
Proof.
Call rank-$i$ flats of size $\alpha_{i}$ in a matroid $M\in\mathcal{M}$ large, and call a flag of $M$ large if it has flats of each rank from 0 to $r(M)$, all of which are large. Thus, if $M\in\mathcal{M}$, then $F_{0,r}(M;\alpha_{0},\alpha_{1},\alpha_{2},\ldots,\alpha_{r})$ is the number of large flags of $M$. It follows from conditions (1)-(2) that each matroid in $\mathcal{M}$ is connected.
Fix any matroids $M_{1},M_{2},\ldots,M_{t}$ in $\mathcal{M}$ on disjoint ground sets. We can assume that $r(M_{i})\geq r(M_{i+1})$ for $i\in[t-1]$. Set $M=M_{1}\oplus M_{2}\oplus\cdots\oplus M_{t}$. We start with some observations about $M$.
By condition (1), for $i\geq 2$, no rank-$i$ flat in $M$ that has elements in two or more of $M_{1},M_{2},\ldots,M_{t}$ has size more than $\alpha_{i-1}+1$. Thus, $M$ satisfies the bounds in condition (3) and any rank-$i$ flat of $M$ of size $\alpha_{i}$ is a flat of some $M_{j}$. Therefore, we can focus on direct sums that satisfy the bounds in condition (3).
For $s\in[t+1]$, let $p_{s}$ be the sum of the first s1s-1 terms among $r(M_{1}),r(M_{2}),\ldots,r(M_{t})$, so $p_{1}=0$ and $p_{t+1}=r(M)$. If $i\in[r(M)-1]$, then $i=p_{s}+u$ for unique $s\in[t]$ and $u\in[r(M_{s})-1]\cup\{0\}$, and the largest rank-$i$ flats in $M$ have size
$\beta_{i}=\alpha_{r(M_{1})}+\alpha_{r(M_{2})}+\cdots+\alpha_{r(M_{s-1})}+\alpha_{u}.$
Set $\beta_{r(M)}=|E(M)|$. We extend the terminology of large flats and large flags to $M$ in the obvious way, using the sequence $\beta_{0},\beta_{1},\ldots,\beta_{r(M)}$. Note that $\beta_{i}\leq\alpha_{i}$ for all $i$, with equality if and only if $i\leq r(M_{1})$. To get the large flags of $M$, start with a large flag of a matroid among $M_{1},M_{2},\ldots,M_{t}$ of largest rank, then take the union of the ground set of that matroid with, in turn, the nonempty sets in a large flag of a matroid of largest rank among $M_{1},M_{2},\ldots,M_{t}$ that has not yet been selected, and repeat until we reach $E(M)$. Condition (1) gives $\beta_{k+1}-\beta_{k}=1$ for $t$ integers $k\in[r(M)]$, namely, $k=p_{s}$ with $s\in[t]$. If, for some $s\in[t]$, a sequence $\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)}$ has $\gamma_{p_{s}}=\beta_{p_{s}}$ and $\gamma_{p_{s}+1}-\gamma_{p_{s}}>1$, then $\gamma_{p_{s}+1}>\beta_{p_{s}+1}$ and so $F_{0,r(M)}(M;\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)})=0$ since $\beta_{p_{s}+1}$ is the largest size of a rank-$(p_{s}+1)$ flat of $M$.
Consider matroids of rank $r(M)$ and size $|E(M)|=\beta_{r(M)}$. We claim that optimizing the valuative invariants below, in the order listed, isolates the direct sums $N_{1}\oplus N_{2}\oplus\cdots\oplus N_{t}$ where $N_{i}\in\mathcal{M}$ and, if the components are indexed so that $r(N_{i})\geq r(N_{i+1})$ for all $i\in[t-1]$, then $r(M_{i})=r(N_{i})$ for all $i\in[t]$.
  • (a)
    To focus on simple matroids, maximize $N\mapsto f_{1}(N;1)$.
  • (b)
    As justified above, minimize $N\mapsto f_{2}(N;j)$ for each $j\not\in A$; also, for all $i,j$ with $3\leq i\leq r(M)$ and $\alpha_{i-1}+1<j<\alpha_{i}$, minimize $N\mapsto f_{i}(N;j)$.
  • (c)
    To focus on matroids with at least $t$ connected components, minimize the sum of the coefficients of the terms $x^{i}y^{j}$ with $i+j<t$ in the Tutte polynomial $T(N;x,y)$.
  • (d)
    For each sequence $\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)}$ where, for some $s\in[t]$, we have $\gamma_{p_{s}}=\beta_{p_{s}}$ and $\gamma_{p_{s}+1}-\gamma_{p_{s}}>1$, minimize $N\mapsto F_{0,r(M)}(N;\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)})$.
  • (e)
    Maximize $N\mapsto F_{0,r(M)}(N;\beta_{0},\beta_{1},\ldots,\beta_{r(M)})$.
Since $\beta_{k+1}-\beta_{k}=1$ if and only if $k=p_{s}$ where $s\in[t]$, simple matroids that satisfy (c) and (e) have $t$ connected components. Let $N=N_{1}\oplus N_{2}\oplus\cdots\oplus N_{t}$ satisfy the sequence of conditions (a)-(e), so each $N_{i}$ is connected. Let $F_{0}\subsetneq F_{1}\subsetneq\cdots\subsetneq F_{r(M)}$ be a flag that $F_{0,r(M)}(N;\beta_{0},\beta_{1},\ldots,\beta_{r(M)})$ counts. Since $\beta_{r(M_{1})}=\alpha_{r(M_{1})}$, we have $F_{r(M_{1})}\subseteq E(N_{j})$ for some $j\in[t]$; we may assume that $j=1$. If $E(N_{1})\neq F_{r(M_{1})}$, then $|X-F_{r(M_{1})}|>1$ for some flat $X$ of $N_{1}$ that covers $F_{r(M_{1})}$ (otherwise more than $t$ sets $F_{i+1}-F_{i}$ would be singletons, so more than $t$ terms $\beta_{i+1}-\beta_{i}$ would be $1$); thus, $F_{0,r(M)}(N;\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)})>0$ for some sequence $\gamma_{0},\gamma_{1},\ldots,\gamma_{r(M)}$ where $\gamma_{i}=\beta_{i}$ for $i\leq r(M_{1})$ and $\gamma_{r(M_{1})+1}-\gamma_{r(M_{1})}=|X-F_{r(M_{1})}|>1$, but the minimization in item (d) prevents that. Thus, $E(N_{1})=F_{r(M_{1})}$. The same ideas show that, up to reindexing, $r(N_{i})=r(M_{i})$ and $E(N_{i})=F_{p_{i+1}}-F_{p_{i}}$ for all $i\in[t]$. We maximize $F_{0,r(M)}(N;\beta_{0},\beta_{1},\ldots,\beta_{r(M)})$ when each $F_{0,r(N_{i})}(N_{i};\alpha_{0},\alpha_{1},\ldots,\alpha_{r(N_{i})})$ is maximized, so each $N_{i}$ is in $\mathcal{M}$, as needed.
We now use the hypothesis that there are valuative invariants $f_{1},f_{2},\ldots,f_{k}$ for which, among all matroids in $\mathcal{M}$, the matroids $M_{1},M_{2},\ldots,M_{t}$ are, up to isomorphism, the unique matroids of their ranks that maximize $f_{1}$, and then $f_{2}$, and so on, through $f_{k}$. By the assumption that each $f_{i}$ is additive on direct sums, $f_{i}(N)$ is maximized precisely when each $f_{i}(N_{j})$ is maximized. By assumption, that is when $N_{j}$ is isomorphic to $M_{j}$. Thus, maximizing $f_{1}$, and then $f_{2}$, and so on through $f_{k}$ leaves one matroid, up to isomorphism, namely, $M$, as claimed. ∎
The assumption that $f_{1},f_{2},\ldots,f_{k}$ are additive can be replaced by any condition that ensures that each $f_{i}$ is maximized on a matroid precisely when it is maximized on the restrictions to its connected components.
Corollary 8.3.
Direct sums of cycle matroids of complete graphs of rank at least two are extremal. For a fixed prime power $q$, direct sums of projective geometries of order $q$ and rank at least four are extremal, and the same holds for affine geometries. If there is only one projective or affine plane of order $q$, up to isomorphism, then planes and lines of order $q$ can be included in the previous statement. For a group $G$ for which $G$ is the only group of order $|G|$, direct sums of Dowling geometries of rank at least four based on $G$ are extremal. For a fixed integer $t\geq 2$, direct sums of (r,r1,t)(r,r-1,t)-spikes of rank at least two are extremal.
Proof.
For all cases except Dowling geometries, the set $\mathcal{M}$ in Theorem 8.2 has, up to isomorphism, a unique matroid of each rank. The first statement follows directly from Theorems 5.3 and 8.2 with $\alpha_{i}=\binom{i+1}{2}$, $A=\{2,3\}$, and $d=2$. The assertion about projective geometries follows Theorems 3.1 and 8.2 with $\alpha_{i}=\frac{q^{i}-1}{q-1}$, $A=\{2,q+1\}$, and $d=4$; likewise for affine geometries, with $\alpha_{i}=q^{i-1}$, for $i\geq 1$, $A=\{2,q\}$, and $d=4$; when, as noted above, those results extend to planes and lines, use $d=2$. For Dowling geometries, the result follows from Theorems 6.3, 6.4, and 8.2 with $\alpha_{i}=i+|G|\,\binom{i}{2}$, $A=\{2,3,|G|+2\}$, and $d=4$ by using a single additional valuative invariant $f_{1}$, namely, the number of $3$-point lines. For (r,r1,t)(r,r-1,t)-spikes, the result follows from Theorems 7.2 and 8.2 with $\alpha_{i}=1+(i-1)t$ for $i\geq 1$, $A=\{2,t+1\}$, and $d=2$. ∎
This result specifies rank at least four for Dowling geometries since only for the first few orders $g$ is there a unique quasigroup of order $g$. That stands in contrast to groups, where there is a unique group of order $g$ if and only if $\gcd(g,\phi(g))=1$, where $\phi(g)$ is Euler’s $\phi$-function; see [28].
Acknowledgments
The author is grateful to Luis Ferroni, Alex Fink, and Peter Nelson for helpful and encouraging e-mail exchanges while this material was being developed.